### v1

# Robust variational neural posterior estimation for simulation-based inference

Abstract
Recent advances in neural density estimation have enabled powerful simulation-based inference (SBI) methods that can flexibly approximate Bayesian inference for intractable stochastic models. Although these methods have demonstrated reliable posterior estimation when the simulator accurately represents the underlying data generative process (GDP), recent work has shown that they perform poorly in the presence of model misspecification. This poses a significant problem for their use on real-world problems, due to simulators always misrepresenting the true DGP to a certain degree. In this paper, we introduce robust variational neural posterior estimation (RVNP), a method which addresses the problem of misspecification in amortised SBI by bridging the simulation-to-reality gap using variational inference and error modelling. We test RVNP on multiple benchmark tasks, including using real data from astronomy, and show that it can recover robust posterior inference in a data-driven manner without adopting tunable hyperparameters or priors governing the misspecification.
1 Introduction
Simulator models are ubiquitous in many areas of the natural sciences and engineering, enabling researchers to approximate complex real-world data-generating processes (DGP) using physically grounded forward models. However, these simulators are often computationally expensive, non-differentiable, and lack closed-form likelihoods, making traditional inference methods inapplicable. Implicitly, the simulator defines an intractable likelihood $p(\bm{x}_{\rm sim}|\,\bm{\theta})$ over $\mathcal{X}_{\rm sim}\subseteq\mathbb{R}^{n}$, relating the simulated observations and the parameters of interest $\bm{\theta}\in\Theta\subseteq\mathbb{R}^{m}$. As a result of intractability and the computational expense of running simulations, solving the inverse problem of inferring simulator parameters from observed data $\bm{x}_{\rm obs}$ poses a significant challenge. Simulation-based inference (SBI, Cranmer et al. 2020) or likelihood-free inference provides methods to approximately infer the posterior distribution of the simulator parameters conditioned on observed data.
A range of SBI methods have emerged to solve the likelihood-free inference problem, beginning with traditional approaches such as approximate Bayesian computation (ABC; Rubin 1984; Beaumont et al. 2002) and Bayesian synthetic likelihood (BSL, Price et al. 2018). Recent work has introduced methods based on neural density estimation, such as neural posterior estimation (NPE, Papamakarios & Murray 2016; Lueckmann et al. 2017; Greenberg et al. 2019), neural likelihood estimation, (NLE, Lueckmann et al. 2019; Papamakarios et al. 2019), neural ratio estimation (NRE, Izbicki et al. 2014; Cranmer et al. 2016; Hermans et al. 2020; Durkan et al. 2020), and diffusion-based methods (Glöckler et al., 2024). SBI methods can be categorised into amortised and non-amortised inference methods. In the context of neural SBI, non-amortised methods such as sequential neural posterior estimation (SNPE; Greenberg et al. 2019), sequential neural likelihood estimation (SNLE; Papamakarios et al. 2019), and sequential neural ratio estimation (SNRE; Hermans et al. 2020) target a single posterior conditioned on fixed data, adapting their inference procedure with each simulation round. After an up-front simulation budget, amortised methods aim to learn a global inference model over a given prior, making them well-suited for scenarios where repeated or scalable inference is required. In this paper, we focus on amortised SBI methods for inferring posterior distributions for any observation within the support of the simulator model.
SBI methods have been widely used in fields such as astronomy (Mishra-Sharma & Cranmer, 2022), particle physics (The Atlas Collaboration, 2025), cosmology (Lemos et al. 2023; Zeghal et al. 2024), and neuroscience (Oesterle et al. 2020; Hashemi et al. 2024), to name a few. However, recent work has shown that they can yield overconfident posterior approximations (Hermans et al., 2022) and suffer significantly when the true DGP does not lie within the family of distributions defined by the statistical model (Cannon et al. 2022; Schmitt et al. 2024), known as model misspecification. Model misspecification may be caused by a variety of factors, such as contamination in the data or unaccounted-for physical processes in the modelling that can lead to overconfident posteriors (Hermans et al., 2022). This discrepancy between the simulated data and the real observations is known as the simulation-to-reality gap (Miglino et al., 1995) or simulation gap.
Methods for mitigating against misspecification in neural SBI have done so mainly by addressing the simulation-to-reality gap. This is based on the assumption that misspecification appears as a divergence-based discrepancy between the true DGP $p^{*}(\bm{x}_{\rm obs})$ and the distribution described by the simulator model $p(\bm{x}_{\rm sim})=\mathbb{E}_{p(\bm{\theta})}[p(\bm{x}_{\rm sim}|\,\bm{\theta})]$, under the prior distribution $p(\bm{\theta})$. Robust SBI methods usually address the simulation-to-reality gap through error modelling and adjustment parameters (Ward et al. 2022; Frazier & Drovandi 2021; Kelly et al. 2024), domain adaptation approaches (Huang et al. 2023; Swierc et al. 2024; Elsemüller et al. 2025; Mishra et al. 2025), or generalized Bayesian inference (Dellaporta et al., 2022). The success of most of these methods relies on the observed points appearing as out-of-distribution (OOD) with respect to the simulated observations. However, recent work has underscored the importance of within-distribution (ID) points in a misspecified SBI (Schmitt et al. 2024;Frazier et al. 2024;Elsemüller et al. 2025) as the errors in the model may still produce summary statistics which lie ID relative to the simulations. Wehenkel et al. (2025) shows that using a reliable calibration set can aid towards robust amortised SBI under such modelling errors. Often, a reliable calibration set will not exist, making such problems highly difficult to solve. Recently, unsupervised domain adaptation (UDA) methods have been implemented in robust amortised SBI using Maximum Mean Discrepancy and domain-adversarial neural networks (Elsemüller et al., 2025), and using consistency loss regularisation (Mishra et al., 2025). As amortised SBI looks to construct general posteriors for a range of observations, it is natural to consider the misspecification problem for situations involving many observations where all points appear OOD, or when a significant number of points appear OOD.
Despite their success in robust SBI, existing methods for robust SBI encounter issues in the context of robust amortised SBI. In particular, the error modelling and correction parameter approaches scale poorly to amortised Bayesian inference due to their dependence on an MCMC sampling step. On the other hand, they benefit from their Bayesian formulation, particularly through the connection between hyperparameter choice and Bayesian prior adoption (Ward et al. 2022; Frazier & Drovandi 2021; Kelly et al. 2024). Domain adaptation methods scale more favourably to amortised SBI, but come at the cost of a non-Bayesian interpretation of the domain adaptation hyperparameters, a lack of interoperability of the domain adaptation (Elsemüller et al., 2025), and a lack of clarity between the trade-off in the domain adaptation and the inference algorithm (Chen et al., 2021). Furthermore, it is not always desirable to use domain-adapted neural embedding statistics if expert knowledge on the summary embedding space is available, such as known sufficient statistics on a low-dimensional observation space in physically motivated units. Data-driven methods that have a reliable Bayesian interpretation and do not rely on hyperparameters are desirable in this context.
Figure 1: Summary statistics for the pendulum task, where many of the misspecified observations (red) will appear within high-probability regions of the marginal density $p(\bm{x}_{\rm sim})$ (underlying hue). Multiple observations will provide information on the simulation-to-reality gap for the pendulum task. Furthermore, it highlights the issue of fitting for the misspecification using a single observation that, if it appears within the distribution, will contain no information about the misspecification. The two right-most images show that by increasing the number of observations, we recover a more reliable inference.
1.1 Our contributions
We propose robust variational neural posterior estimation (RVNP) and its tuned variant (RVNP-T) that addresses the misspecification problem in amortized SBI by pre-training the simulator likelihood $p_{\bm{\Psi}}(\bm{x}_{\rm sim}|\,\bm{\theta})$ for parameters $\bm{\Psi}$, adopting a flexible error model $p_{\bm{\xi}(\bm{\theta})}(\bm{x}_{\rm obs}|\,\bm{x}_{\rm sim})$, and using an importance weighted autoencoder (Burda et al., 2015) scheme to maximize the evidence of the true data under the variational posterior $p_{\bm{\phi}}(\bm{\theta}|\,\bm{x}_{\rm obs})$ for the parameters $\bm{\xi}$ and $\bm{\phi}$. The variational posterior is the main objective of the RVNP algorithm. The posterior $p_{\bm{\phi}}(\bm{\theta}|\,\bm{x}_{\rm obs})$ is optionally tuned on the adapted synthetic DGP to return the RVNP-T posterior distribution. The error modelling we choose to adopt can be seen as a neural adaptation of the mean adjustment and covariance inflation from Frazier & Drovandi (2021).
The main claim of our paper is that robust variational neural posterior estimation (RVNP) and its tuned variant can recover robust amortised posterior inference under misspecification by bridging the simulation-to-reality gap using error modelling. The error model parameters are adapted in a data-driven way when we have many observations. We summarise our contributions as follows:
  1. 1.
    We introduce RVNP, an amortised SBI method that uses a pre-trained simulator likelihood, an error model, and an importance-weighted autoencoder (Burda et al., 2015) scheme to return robust amortised posterior inference under misspecification without adopting tunable parameters or priors over the misspecification. We also introduce RVNP-T, which tunes the final posterior using the simulator and the noise induced by the error model.
  2. 2.
    We investigate the effect of the number of observed data points on error modelling in robust SBI for the first time.
  3. 3.
    To our knowledge, this is the first example of using amortised variational autoencoders to address the misspecification problem in SBI.
In amortised SBI, it is usually necessary to have a pre-defined up-front simulator sample size, which may change if the model is learnt in rounds. We assume that we have a fixed simulation budget and that the neural statistic embedding is either pre-trained or adopted by expert knowledge. The simulation-to-reality gap should inform our error model as we observed more data from the true DGP (Figure 1). We impose a strong inductive bias on the domain adaptation so that the error model can only inflate the synthetic DGP to account for the simulation gap, allowing us to have a form of model criticism. Cranmer et al. (2020) suggests augmenting the simulator with an error model to account for the simulation-to-reality gap. However, choosing an error model can be difficult and arbitrary and requires multiple independent trainings to test the different error models (Ward et al., 2022). We instead consider using an error model that is inferred by maximising the evidence lower bound jointly with inferring the posterior distribution of the parameters. Our method is well-suited to situations when the misspecified points appear, on average, OOD relative to the simulated points. Current unsupervised robust SBI methods rely on a clear simulation-to-reality gap driven by misspecification. In situations where the entire inference set lies within or very near the original simulated distribution, a strong calibration set is most likely necessary.
Overview of Paper. In Section 2, we provide an overview of the background necessary for the paper. Section 3 describes the methods that will be applied to experiments in Section 4. In Section 5, we discuss related works. We conclude the paper in Section 6 with a discussion and conclusion.
2 Background
2.1 SBI formalism
Let $\bm{\theta}\in\Theta\subseteq\mathbb{R}^{m}$ be the target parameters of interest which are to be inferred after adopting a prior distribution $p(\bm{\theta})$.
Let $\bm{x}_{\rm sim}\in\mathcal{X}_{\rm sim}\subseteq\mathbb{R}^{n}$ denote a simulated observation. The simulator is a family of distributions parametrized by $\bm{\theta}$ which can be represented by an unknown density $p(\bm{x}_{\rm sim}|\,\bm{\theta})$ over $\mathcal{X}_{\rm sim}$, relating the simulated observations and the parameters of interest $\bm{\theta}\in\mathbb{R}^{m}$.
We denote $\bm{x}_{\rm obs}\in\mathcal{X}_{\rm obs}\subseteq\mathbb{R}^{n}$ as a true observation and $\bm{\theta}^{*}\in\mathbb{R}^{m}$ as the ground truth of the parameter $\bm{\theta}$ for an experiment. We let $p^{*}(\bm{x}_{\rm obs})$ denote the true, unknown, DGP.
We define the error model as $p_{\bm{\xi}(\bm{\theta})}(\bm{x}_{\rm obs}|\,\bm{x}_{\rm sim})$, a family of distributions parametrized by $\bm{\xi}(\bm{\theta})\in\mathbb{R}^{k}$ and $\bm{x}_{\rm sim}$, which relates the simulations to observations. The density explicitly depends on $\bm{x}_{\rm sim}$ and implicitly on $\bm{\theta}$ due to the error model parameters $\bm{\xi}$ being a function of $\bm{\theta}$.
We assume that we have $D=\{\bm{\theta}^{(i)},\bm{x}_{\rm sim}^{(i)}\}_{i=1}^{N_{\rm sim}}$, a fixed number of points generated from the synthetic DGP $p(\bm{x}_{\rm sim}|\bm{\theta})p(\bm{\theta})$, and a set of observations, $O=\{\bm{x}_{\rm obs}^{(i)}\}_{i=1}^{N_{\rm obs}}$ each associated to a different, unknown true $\bm{\theta}^{*}$ value.
We let $\iota_{\omega}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{l}$, $\bm{x}_{\rm sim}\mapsto\bm{z}_{\rm sim}$ denote a statistical embedding parametrized by $\omega$. This embedding can represent fixed user-defined summary statistics, a pre-defined embedding, or a neural statistic estimator (NSE) where the parameters $\bm{\omega}$ are to be learnt. Lower-dimensional embeddings are important when dealing with high-dimensional data, but can come at the cost of information loss when the embedding is not a sufficient statistic of $\bm{x}_{\rm sim}$ for $\bm{\theta}$ (Blum et al., 2013).
(a) CS posterior
(b) SIR posterior
(c) Pendulum posterior
(d) Spectra posterior
Figure 2: Samples from posterior distribution conditional on a single observed point when RVNP was trained on $N_{\rm obs}=1000$ different observations. The green corresponds to RVNP, the blue corresponds to NPE, and the red point (dashed line) corresponds to the true $\bm{\theta}^{*}$. The grey corresponds to the training samples. We see that RVNP is significantly more robust than NPE, particularly in the complex pendulum and spectra task.
2.2 Amortised neural posterior and neural likelihood estimation
The goal of amortised neural posterior estimation is to approximate the unknown posterior distribution $p^{*}(\bm{\theta}|\bm{x_{\rm obs}})$ for all $\bm{x}_{\rm obs}\in\mathcal{X}_{\rm obs}$. After choosing a conditional density estimation architecture $p_{\bm{\phi}}(\bm{\theta}|\,\bm{x}_{\rm obs})$ parametrized by $\bm{\phi}$, and an architecture for the neural statistic embedding, $\iota_{\omega}$, NPE (Papamakarios & Murray 2016; Lueckmann et al. 2017; Greenberg et al. 2019) fits for the parameters $\bm{\omega}$ and $\bm{\phi}$ by minimizing the the expected forward Kullback-Leibler (KL) divergence between analytic and approximate posterior
$\mathcal{L}_{\rm NPE}({\bm{\phi},\bm{\omega}})$ $=\mathbb{E}_{p^{*}(x_{\rm obs})}\big{[}\mathbb{KL}[p(\bm{\theta}|\,\iota_{\bm{\omega}}(\bm{x}_{\rm obs}))||p_{\bm{\phi}}(\bm{\theta}|\,\iota_{\bm{\omega}}(\bm{x}_{\rm obs}))]\big{]}$ (1)
$=\mathbb{E}_{p^{*}(x_{\rm obs})}\big{[}\mathbb{E}_{p(\bm{\theta}|\,\iota_{\bm{\omega}}(\bm{x}_{\rm obs}))}[-\log p_{\bm{\phi}}(\bm{\theta}|\,\iota_{\bm{\omega}}(\bm{x}_{\rm obs}))]\big{]},$
where the expectation is over the unknown true data-generating distribution $p^{*}(\bm{x}_{\rm obs})$ and the second line follows because the true unknown posterior and its entropy do not depend on the trainable parameters. As noted in Schmitt et al. (2024), this amortised posterior objective function is not feasible as we may not have enough real data to approximate the expectation with respect to $p^{*}(\bm{x}_{\rm obs})$, and the true posterior is unknown and intractable. Instead, the unknown $p^{*}(\bm{x}_{\rm obs})$ is replaced with the marginal likelihood $p(\bm{x_{\rm obs}})=\int p(\bm{x}_{\rm obs}|\,\bm{\theta})p(\bm{\theta})d\bm{\theta}$. Under the model assumption we have $\bm{x}_{\rm obs}$ can be replaced with $\bm{x}_{\rm sim}$ and the objective becomes
$\mathcal{L}(\bm{\phi},\bm{\omega})_{\rm NPE}:=-\mathbb{E}_{\bm{x}_{\rm sim},\bm{\theta}}[\log p_{\bm{\phi}}(\bm{\theta}|\,\iota_{\bm{\omega}}(\bm{x}_{\rm sim}))].$ (2)
This is minimised with respect to the parameters $\bm{\omega}$ and $\bm{\phi}$. The success of this objective depends on the assumption that sampling from the evidence is equivalent to sampling from $p^{*}(\bm{x}_{\rm obs})$.
On the other hand, neural likelihood estimation (NLE) learns a distribution maximising the conditional log-probability of the simulated data
$\mathcal{L}(\bm{\Psi})_{\rm NLE}:=\mathbb{E}_{p(\bm{x}_{\rm sim}|\bm{\theta})p(\bm{\theta})}[\log p_{\bm{\Psi}}(\bm{x}_{\rm sim}|\bm{\theta})]$ (3)
with respect to the flow parameters $\bm{\Psi}$ (where we dropped reference to $\iota_{\bm{\omega}}$), which is equivalent to minimising the KL divergence between the flow and the target distribution. The assumption of the learnt distribution being a likelihood identifies $\bm{x}_{\rm obs}$ with $\bm{x}_{\rm sim}$. However, one can also view the NLE objective as a surrogate training objective to approximate the sampling distribution given by the simulator, without any initial assumptions about the connection between true data and simulator output. In this sense, the NLE objective becomes a neural conditional prior proxy, and a likelihood can be introduced to relate the forward model to the true $\bm{x}_{\rm obs}$.
2.3 Misspecification in SBI
The simulator, $p(\bm{x}_{\rm sim}|\,\bm{\theta})$, is said to be misspecified if the true data-generating process does not fall within the family of distributions defined by the simulator on the support of the prior of $\bm{\theta}$. That is, $q^{*}\notin\{p(x_{\rm sim}|\,\bm{\theta})\,;\bm{\theta}\in\rm{supp}(p(\bm{\theta}))\}$ (Cannon et al., 2022). For single observations, the definition of misspecification was extended to using summary statistics (Kelly et al., 2024) by defining $\bm{b}(\bm{\theta})=\mathbb{E}_{p(\bm{x}_{\rm sim}\mid\bm{\theta})}[\iota_{\omega}(\bm{x}_{\rm sim})]$ and $\bm{b}_{j}=\mathbb{E}_{q^{*}(\bm{x}_{\rm obs})}[\iota_{\omega}(\bm{x}_{\rm obs}^{(j)})]$, for each $\bm{x}_{\rm obs}^{(j)}\in O$ where $O=\{\bm{x}_{\rm obs}^{(i)}\}_{i=1}^{N_{\rm obs}}$ is the set of observations each associated to a different, unknown true $\bm{\theta^{(j)}}$ value. Then the the simulator is misspecified if there is no $\{\bm{\theta}\}_{i=1}^{N_{\rm obs}}$ in the support of the prior for which $\bm{b}(\bm{\theta}^{(j)})=\bm{b}_{j}$ for each $j$.
In this paper, we adopt the alternate, but similar, definition of misspecification provided in Wehenkel et al. (2025), which is more aligned with amortised SBI. They define a simulator to be misspecified if $\exists\mathcal{S}\subseteq\Theta\times\mathcal{X}_{\rm obs}:\forall(\bm{\theta},\bm{x}_{\rm obs})\in\mathcal{S}$
$p(\bm{\theta})=p^{*}(\bm{\theta})\text{ and }p^{*}({\bm{\theta}|\,\bm{x}_{\rm obs}})\neq p({\bm{\theta}|\,\bm{x}_{\rm sim}=\bm{x}_{\rm obs}}).$ (4)
This definition aligns with the amortised SBI task as it identifies misspecification as a set-wise phenomenon. Our goal is to recover robust and reliable posterior inference for all $\bm{x}_{\rm obs}\in\mathcal{X}_{\rm obs}$. This definition does, however, ignore situations when the prior distribution is misspecified. Our method can handle prior misspecification more favourably through traditional Bayesian model iterations, provided the support of the proposed prior is a subset of the support of the original prior.
2.4 Importance weighted autoencoders
Importance-weighted autoencoders (IWAE) were inspired by the vanilla variational autoencoder (VAE, Kingma & Welling 2013), which introduces a latent space $\mathcal{Z}\subseteq\mathbb{R}^{m}$ and two distributions, $p_{\bm{\xi}}(\bm{x}|\,\bm{z})$ and $p_{\bm{\phi}}(\bm{z}|\,\bm{x})$, known as the decoder and the encoder respectively. The vanilla VAE objective aims to maximise the evidence of the data via the evidence lower bound
$\log p(\mathbf{x})\;\geq\;\mathbb{E}_{p_{\bm{\phi}}(\mathbf{z}|\mathbf{x})}\left[\log p_{\bm{\xi}}(\mathbf{x}|\mathbf{z})\right]-\mathrm{KL}\!\left(p_{\bm{\phi}}(\mathbf{z}|\mathbf{x})\,\|\,p(\mathbf{z})\right)$ (5)
with respect to $\bm{\xi}$ and $\bm{\phi}$, where $p(\bm{z})$ is a prior over the latent parameter. However, using this bound has been shown to induce mode-seeking behaviour, leading to overly simplified representations and poor inference. Burda et al. (2015) introduced IWAE, which is based on a strictly tighter evidence lower bound derived from importance sampling. The log-evidence for a single point is given as
$\log p_{\bm{\theta}}(\textbf{x})$ $\approx\log\left[\frac{1}{k}\sum_{l=1}^{k}\frac{p_{\bm{\xi}}\left(\textbf{x},\textbf{z}^{(l)}\right)}{p_{\bm{\phi}}\left(\textbf{z}^{(l)}|\textbf{x}\right)}\right]={\mathcal{L}}^{\text{IWAE}}_{k}\left(\bm{\xi},\bm{\phi};\textbf{x}\right)$ (6)
$\text{with}\quad\textbf{z}^{(l)}\sim p_{\bm{\phi}}\left(\textbf{z}|\textbf{x}\right),$ (7)
which is a mass-covering objective that targets the evidence.
3 Method: RVNP
Motivated by a desire to have a method that can recover robust amortized posterior inference with a fixed embedding space such that the simulation-to-reality gap is bridged with an interpretable, flexible error model, we extend the approaches of Ward et al. (2022) and Kelly et al. (2024) to amortized SBI, and use an importance weighted autoencoder (Burda et al., 2015) amortized variational inference scheme to define robust variational neural posterior estimation (RVNP) and its tuned variant (RVNP-T). In what follows, we drop the explicit embedding notation $\iota_{\omega}$ unless necessary. Our goal is to use a normalizing flow parametrized by $\bm{\phi}$, $p_{\bm{\phi}}(\bm{\theta}|\,\bm{x}_{\rm obs})$, that can approximate the true posterior distribution. Algorithm 1 provides an overview of the RVNP and RVNP-T algorithms. We chose a fixed simulation budget of $N_{\rm sim}=100,000$ for each task, with $10\%$ retained for validation.
RVNP builds on variational methods for solving the inverse problem under a learnt likelihood in SBI (Glöckler et al., 2022), and relies on a conditional neural density estimator to approximate the likelihood from a fixed budget of simulations from the simulator model. This overcomes the computational expense of using Hamiltonian Monte Carlo Neal (2011) or other Markov Chain methods to jointly infer the parameters of interest and the parameters of the error model. Furthermore, variational methods allow us to use more expressive models for the error model distribution. We assume a pre-training or fixing of the neural statistic embedding $\iota_{\omega}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{l}$ occurred.
3.1 Generative model
Throughout, we assume that the neural embedding or summary statistics are pre-trained and are not jointly learnt. The first step of RVNP is to train the normalising flow $p_{\bm{\Psi}}(\bm{x}|\,\bm{\theta})$ to approximate the likelihood from the generated samples from the simulator using the NLE objective (Equation 3). Once the normalizing flow has been trained, we assume that $p(\bm{x}_{\rm sim}|\,\bm{\theta})\approx p_{\bm{\Psi}}(\bm{x}_{\rm sim}|\,\bm{\theta})$ and include the surrogate in the forward model.
We assume that the true DGP can be modelled as
  1. 1.
    $p(\bm{\xi})$ a (pseudo-)prior over the error model parameters.
  2. 2.
    $\bm{\theta}\sim p(\bm{\theta})$, where $p(\bm{\theta})$ is known and tractable.
  3. 3.
    $\bm{x}_{\rm sim}\sim p_{\bm{\Psi}}(\bm{x}_{\rm sim}|\,\bm{\theta})$.
  4. 4.
    $\bm{x}_{\rm obs}\sim p_{\bm{\xi}(\bm{\theta})}(\bm{x}_{\rm obs}|\,\bm{x}_{\rm sim})$, where $p_{\bm{\xi}(\bm{\theta})}(\bm{x}_{\rm obs}|\,\bm{x}_{\rm sim})=\mathcal{N}(\bm{x}_{\rm obs};\,\bm{x}_{\rm sim},\bm{\xi}(\bm{\theta}))$ is an adopted error model conditional on $\bm{\xi}$, where the covariance matrix is the output of a neural network $\rm{NN}(\bm{\theta})$ parametrized by $\bm{\alpha}$.
Under this generative model, the posterior distribution is proportional to
$p(\{\bm{\theta}^{(i)}\}_{i=1}^{N_{\rm obs}},\bm{\xi}\mid O)\propto\prod_{i=1}^{N_{\rm obs}}\int p_{\bm{\xi}(\bm{\theta}^{(i)})}(\bm{x}^{(i)}_{\rm obs}\mid\bm{x}^{(i)}_{\rm sim})\,p_{\bm{\Psi}}(\bm{x}^{(i)}_{\rm sim}\mid\bm{\theta}^{(i)})\,p(\bm{\theta}^{(i)})\,p(\bm{\xi}(\bm{\theta}^{(i)}))\,d\bm{x}^{(i)}_{\rm sim}$. (8)
3.2 Variational posterior
From the posterior distribution of our forward model (Equation 8), we can express the log-evidence of the data as
$\log p(O)=\sum_{i=1}^{N_{\rm obs}}\log\mathbb{E}_{p_{\bm{\Psi}}(\bm{x}^{(i)}_{\rm sim}|\,\bm{\theta}^{(i)})}[p_{\bm{\xi}(\bm{\theta}^{(i)})}(\bm{x}^{(i)}_{\rm obs}|\,\bm{x}^{(i)}_{\rm sim})p(\bm{\theta}^{(i)})p(\bm{\xi}(\bm{\theta}^{(i)}))].$ (9)
We can use the IWAE lower bound on the log-evidence (Equation 6) to derive the variational loss function for RVNP as
$\mathcal{L}(\bm{\phi},\bm{\alpha})_{V}\approx-\sum_{i=1}^{N_{\rm obs}}\log\left[\frac{1}{K}\sum_{l=1}^{K}\frac{\mathbb{E}_{p_{\bm{\Psi}}(\mathbf{x}_{\rm sim}\mid\bm{\theta}^{(l)})}\!\left[\,p_{\bm{\xi}(\bm{\theta}^{(i)}\,;\,\bm{\alpha})}\!\left(\mathbf{x}_{\rm obs}^{(i)}\mid\mathbf{x}_{\rm sim}\right)\right]\,p(\bm{\theta}^{(l)})p(\bm{\xi}(\bm{\theta}^{(i)}\,;\,\bm{\alpha}))}{p_{\bm{\phi}}(\bm{\theta}^{(l)}\mid\mathbf{x}_{\rm obs}^{(i)})}\right]$ (10)
where $\bm{\theta}^{(l)}\sim p_{\bm{\phi}}(\bm{\theta}\mid\mathbf{x}_{\rm obs}^{(i)})$, $\bm{\alpha}$ are the weights of the neural network $\bm{\xi}(\bm{\theta}^{(i)}\,;\,\bm{\alpha})$, and for each $\bm{\theta}^{(l)}$ we approximate $\mathbb{E}_{p_{\bm{\Psi}}(\mathbf{x}_{\rm sim}\mid\bm{\theta}^{(l)})}\!\left[\,p_{\bm{\xi}(\bm{\theta}^{(i)}\,;\,\bm{\alpha})}\!\left(\mathbf{x}_{\rm obs}^{(i)}\mid\mathbf{x}_{\rm sim}\right)\right]$ using a Monte Carlo (MC) estimate. We use the logsumexp function to ensure the MC estimates are stable.
The second step of RVNP is to minimize $\mathcal{L}(\bm{\phi},\bm{\alpha})_{V}$ for $\bm{\phi},\bm{\alpha}$. Assuming that the likelihood function has been learnt exactly, this objective is theoretically motivated by maximising the evidence of the data. This returns error model parameters and posterior parameters that maximise the evidence lower bound.
3.3 Posterior tuning
RVNP-T, the tuned variant of RVNP, includes an extra tuning step that fixes the neural network parameters of the error model $\bm{\alpha}$, and uses the original simulated dataset $D$ to optimise the adjusted NPE objective
$\mathcal{L}(\bm{\phi})_{\rm{NPE}(\alpha)}:=-\mathbb{E}_{p(\bm{x}_{\rm sim},\bm{\theta})}\mathbb{E}_{p_{\bm{\xi}(\bm{\theta})}(\bm{x}_{\rm obs}|\,\bm{x}_{\rm sim})}[\log p_{\bm{\phi}}(\bm{\theta}|\,\bm{x}_{\rm obs})].$ (11)
This final objective can be identified with the noisy neural posterior estimation (NNPE, Ward et al. 2022) objective and the error augmentation method suggested by Cranmer et al. (2020). However, our error model does not have to be globally fixed and has been inferred using variational inference.
3.4 Error modelling
In the experiments, we adopt two error models. The default error model that we adopt in RVNP is given by the Gaussian covariance matrix
$\bm{\xi}(\bm{\theta})=\rm{Diag}(\rm{NN}(\bm{\theta}\,;\bm{\alpha}))+\bm{\Lambda},$ (12)
a neural network that outputs the diagonal components of the covariance matrix and the non-diagonal components $\bm{\Lambda}$ are globally learnt. We also include a global error model.
$\bm{\xi}(\bm{\theta})=\bm{\Sigma},$ (13)
defined through a full rank Gaussian covariance matrix that is parametrised in terms of a Cholesky decomposition. This error model is constant across the parameter space and does not explicitly depend on $\bm{\theta}$. Our method generalises very easily to include any inductive bias that we believe explains the simulation-to-reality gap.
Our approach can be understood in terms of the simulator defining a population prior $p(\bm{\theta},\bm{x}_{\rm sim})=p_{\bm{\Psi}}(\bm{x}_{\rm sim}|\,\bm{\theta})p(\bm{\theta})$, while the error model can be identified with the likelihood of $\bm{x}_{\rm sim}$ for the observed data. In this framework, the simulator model describes our prior understanding of the true DGP. In this paper, we allow only for Gaussian error-type models that allow the forward model to inflate the noise on the output of the simulator. This allows us to account for the misspecification using a Gaussian forget mechanism.
Including an error model is an assumption about the true DGP, and the exact corruption may be unknown. However, we follow the Box principle of all models are wrong, some are useful (Box, 1976) and attempt to forward model the DGP using the simulator and an error model term that can account for corruption in the data. Error modelling in robust SBI is not new; Ward et al. (2022) introduced a spike and slab error model, which assumes that a summary statistic is either well-specified or misspecified. The parameter adjustment method of Kelly et al. (2024) can be thought of an error model $p_{\bm{\xi}}(\bm{x}_{\rm obs}|\,\bm{x}_{\rm sim})=\delta(\bm{x}_{\rm obs}-\bm{x}_{\rm sim}-\bm{\xi})$, integrating with respect to $\bm{x}_{\rm sim}$ first.
Adjustment parameters will move the observed point to a region of high probability with respect to the simulated samples. On a point-by-point basis, there are infinitely many solutions to this problem, and there is no guarantee of where an OOD point should be mapped back to in the original sample without a high volume of data. Furthermore, the spike-and-slab error model invokes the assumption that the true corruption model is unknown, and that the misspecification forces us to forget certain summary statistics if they lie OOD. This assumption relies on the misspecification occurring along individual axes, which is unrealistic in many situations and is subject to the hyperparameter choices of the spike-and-slab error model. In general, misspecification has no guarantee of being represented exactly along the given axes, even if the components represent summary statistics. To consider why, let $p_{\gamma}(\bm{x})$ be the probability density function parametrised by $\gamma$ and assume that $T(\bm{x})$ represents a sufficient statistic of $\bm{\theta}$. By the Fisher-Neyman factorization theorem (Hogg et al. 2005, pp 376-377), there exist non-negative functions $k_{\gamma}$ and $h$ such that $p_{\gamma}(\bm{x})=h(\bm{x})k_{\gamma}(\bm{x})$. If $g$ is a bijection, then
$f_{\gamma}(x)=h(x)\,k_{\gamma}(g^{-1}(g(T(x)))),$ (14)
$f_{\gamma}(x)=h(x)\,(k_{\gamma}\circ g^{-1})(g\circ T(x)),$ (15)
and, therefore, $g\circ T(x)$ is also a sufficient statistic. In this paper, we choose to model a full covariance matrix to account for misspecification that may not occur exactly along each axis. While suitable pre-processing may help address this issue, we choose to model the full covariance and infer this.
3.5 Prior over the correction model parameters
In RVNP, a prior over the error model parameters can be introduced to regularise the influence of the error model and make the inference more robust in the single-posterior case. Adopting an error model makes the generative model overparameterized, and the inference process will be heavily dependent on the prior assumptions about misspecification in the single-point inference. In our experiments, we do not use any prior distribution over the error model parameters. This allows us to focus on the robustness imparted by multiple observations.
4 Experiments and results
We evaluate our method and test the main claim of our paper: that RVNP and its tuning variant can recover robust amortised posterior inference. In the experiments section, we test four variants of our algorithm
  • RVNP: the standard RVNP algorithm with no final tuning step, using the error model given in Equation 12.
  • RVNP-Global: the standard RVNP algorithm with no final tuning step, using the global error model given in Equation 13.
  • RVNP-T: the tuned version of RVNP.
  • RVNP-Global-T: the tuned version of RVNP-Global.
In each experiment, we only test the tuned algorithm for whichever algorithm performed better, RVNP or RVNP-Global. The architecture and training procedure for each task are described in the Appendix.
Benchmarking Algorithms. In these tasks, we benchmark RVNP against noisy neural posterior estimation (NNPE) from Ward et al. (2022) due to the similar performance of NNPE with RNPE therein and NNPE’s amortisation ability. We do not benchmark against the MMD or consistency loss methods (Elsemüller et al. 2025; Mishra et al. 2025) due to their dynamic adaptation of the neural statistics. We also benchmark against vanilla NPE.
Summary of Results. We find that RVNP and RVNP-T can recover robust posterior inference in an amortised manner across a range of different tasks, including cases where a significant number of points have ID data. Furthermore, we test RVNP and RVNP-T using low-resolution spectra cross-matched with high-resolution spectroscopic data, and we validate our posterior recovery against classically derived results. For each task, except the Spectra task, we test our method for $N_{\rm obs}\in\{1,10,100,1000,10000\}$ points and evaluate the effectiveness of our methods. We note that in the absence of misspecification, RVNP collapses towards NPE and is well-specified and calibrated.
RVNP and RVNP-T can recover robust posterior inference without adopting priors over misspecification or introducing ad hoc hyperparameters that must be tuned. We display examples of posterior inference in the $N_{\rm obs}=1000$ case for RVNP against NPE for each of the tasks for a single example in Figure 2. We argue that the success of RVNP, the interpretability of the error correction model, and the lack of hyperparameters to be tuned (other than the optimiser hyperparameters) point to a significant contribution to robust amortised SBI. In all tasks, our algorithm is overconfident for $N_{\rm obs}=1$. This is not surprising, as without any other information, our model collapses to the NPE solution. With stronger priors over the misspecification, we could make this step more robust, but this defeats the purpose of allowing the data to drive the error model.
4.1 Metrics for assessing misspecification
We consider three main metrics to assess the robustness of the inference. Let $\bm{\theta}^{*}$ be the true value of the parameter. Assuming that we have a labelled test set $T=\{\bm{\theta}^{*(i)},\bm{x_{\rm obs}}^{(i)}\}_{i=1}^{N_{\rm test}}$, we evaluate the log posterior probability (LPP) of the true parameter over the dataset to assess the performance of RVNP. LPP has been extensively used in SBI literature (Papamakarios & Murray 2016; Hermans et al. 2020; Ward et al. 2022;Kelly et al. 2024; Wehenkel et al. 2025).
Given a credible level $\gamma$, let $\rm{HDR}_{p(\bm{\theta}|\,\bm{x}_{\rm obs})}(1-\gamma)$ represent the $1-\gamma$ highest posterior density region of the posterior $p(\bm{\theta}|\,\bm{x}_{\rm obs})$. The expected posterior coverage (EPC) at a given confidence level over a test set is given by
$\rm{EPC}(\gamma):=\mathbb{E}_{\bm{\theta}^{*},\bm{x}_{\rm obs}\sim T}[\mathbf{1}\{\theta^{*}\in\rm{HDR}_{p(\bm{\theta}|\,\bm{x}_{\rm obs})}(1-\gamma)\}],$ (16)
where $\mathbf{1}$ is the indicator function. EPC is a commonly used metric to assess robustness and calibration of posterior distributions, particularly when looking at single observation situations. When comparing posterior calibration across a range of amortised datasets, we compute the average expected posterior coverage AEPC
$\alpha:=\int_{0}^{1}[\rm{EPC}(\gamma)-\gamma]d\gamma,$ (17)
which represents the average calibration across the test set. Finally, we also compute the normalised (root) mean squared error (NMSE):
$\text{NMSE}=\frac{1}{N_{\rm obs}}\sum_{j=1}^{N_{\rm obs}}\frac{\sqrt{\frac{1}{S}\sum_{s=1}^{N_{\rm samples}}\left(\theta^{*}_{j}-\theta^{(s)}_{j}\right)^{2}}}{(\max(\theta_{\rm{prior}})-\min(\theta_{\rm{prior}}))},$ (18)
which evaluates the normalised accuracy of the posterior prediction to the truth relative to the prior width.
4.2 Task A: CS
We reproduce the cancer and stromal cell development benchmark task from Ward et al. (2022). The simulator models the development of cancer and stromal cells in 2D space based on the locality of a cell relative to unobserved parents. This is emulated conditional on three Poisson rate parameters $\bm{\theta}=(\lambda_{c},\lambda_{p},\lambda_{d})$. The total number of cells $N^{c}$, number of unobserved parents $N^{p}$, and the number of daughter cells for each parent $N_{i}^{d}$ are sampled as $N^{c}\sim\rm{Poisson}(\lambda_{c})$, $N^{p}\sim\rm{Poisson}(\lambda_{p})$, and $N^{d}_{i}\sim\rm{Poisson}(\lambda_{d})$ for $i=1,...,N^{p}$. Cell locations $\{c_{i}\}_{i=1}^{N_{c}}$ and disease origin points $\{p_{i}\}_{i=1}^{N_{p}}$ were sampled uniformly across the 2D domain using homogeneous spatial point processes. For each origin point $p_{i}$, $r_{i}$ is the Euclidean distance to its $N_{d}^{i}$-th nearest cell. Cells falling within or on the boundary of this radius from $p_{i}$ are designated as cancerous. Distance-based summary measures were estimated by randomly sampling 50 stromal cells. The summary statistics are as follows: N Cancer and N Stromal, the number of cancer and stromal cells, respectively; and (Mean Min Dist) and (Max Min Dist), the mean and maximum distance from the stromal cells to their nearest cancer cell, respectively. The Numba just-in-time implementation of this task was taken directly from the data products of Ward et al. (2022).
Misspecification. The misspecification in the observed data is introduced by removing cells in the core regions of tumours, which mimics necrosis.
Results. We display the results for the CS task in Figure 3 and provide an example posterior in the $N_{\rm obs}=1000$ case in Figure 2. Finally, we display the simulated observations together with observed points in Figure 7. In this task, the misspecification is minimal, and all versions of our algorithm can recover robust posterior inference beginning at the $N_{\rm obs}=10$ point. We find that all variants of RVNP perform similarly to NNPE in this task. This is not surprising, as the misspecification can be described by inflating the covariance along two specific axes. The RVNP algorithms slightly outperform NNPE in NMSE.
Figure 3: Results for the CS task. We conclude that RVNP and its variants can recover robust posterior inference in amortised simulation-based inference. The hue in the middle plots indicates the error bar on the NPE and NNPE algorithms. For $\alpha$ nearest to 0 is best, with positive values representing underconfidence and negative values representing overconfidence. For the log-probability, higher values are better. For NMSE, lower values are better.
4.3 Task B: SIR
We include the misspecified susceptible-infected-recovered (SIR) task from Ward et al. (2022), which takes the stochastic model of epidemic spread modelled conditional on $\bm{\theta}=(\beta,\gamma)$, the time-varying infection rate and the recovery rate, respectively. The SIR model emulates ideal disease transmission dynamics from the susceptible ($s$), infected ($i$), and recovered ($r$) parameters as
$\frac{ds}{dt}=-\beta si,\quad\frac{di}{dt}=\beta si-\gamma i,\quad\frac{dr}{dt}=\gamma i.$ (19)
Ward et al. (2022) employs a stochastic extension by using time-dependent transmission dynamics through a variable infection rate $\tilde{\beta}_{t}$, accounting for external factors such as policy interventions or pathogen mutations. This stochastic process is characterized using the basic reproduction number $R_{0t}=\frac{\tilde{\beta}_{t}}{\gamma}$, which follows the mean-reverting stochastic differential equation:
$dR_{0t}=\eta\left(\frac{\beta}{\gamma}-R_{0t}\right)dt+\sigma R_{0}dW_{t},$ (20)
where $\eta$ controls the mean reversion strength of $R_{0t}$ toward the equilibrium value $\frac{\beta}{\gamma}$, $\sigma$ represents the volatility parameter, and $W_{t}$ denotes standard Brownian motion. $\eta=0.05$ and $\sigma=0.05$ are fixed and the goal is to infer the parameters $\beta$ and $\gamma$. The Julia code to sample from this process was taken directly from the data products of Ward et al. (2022). The summary statistics produced in this task are the mean, median and maximum number of infections, the day of the maximum number of infections, and the day at which half of the total number of infections was reached, and the mean autocorrelation of infections with lag 1.
Figure 4: Results for the SIR task. We conclude that RVNP and its variants can recover robust posterior inference in amortised simulation-based inference. The hue in the middle plots indicates the error bar on the NPE and NNPE algorithms. For $\alpha$ nearest to 0 is best, with positive values representing underconfidence and negative values representing overconfidence. For the log-probability, higher values are better. For NMSE, lower values are better.
Misspecification To introduce misspecification in the observations, a small reporting delay is adopted where weekend infection counts are reduced by $5\%$ and are added to the Monday count.
Results. We display the results for the SIR task in Figure 4 and provide an example posterior in the $N_{\rm obs}=1000$ case in Figure 2. Finally, we display the simulated observations together with observed points in Figure 8. In this task, the misspecification is extreme but occurs only along the final axis, and all versions of our algorithm can recover robust posterior inference beginning at the $N_{\rm obs}=10$ point. As the misspecification occurs along a given axis, we expect NNPE to perform well in this task. We find that each of our algorithms witnesses slight performance gains in each of the metrics as the number of observations increases.
4.4 Task C: Pendulum Task
We describe a stochastic pendulum simulator that, given $\bm{\theta}:=[\omega_{0},A]$, samples the horizontal position of a frictionless pendulum at 200 time points evenly sampled every 0.05 seconds $\bm{x}_{\rm sim}=(f(t_{0}),...,f(t_{200}))$ where $f(t)=A\cos(\omega_{0}t+\phi)$ for $\phi\sim U(-\pi,\pi)$. In this task, $\omega_{0}$ and $A$ denote the fundamental frequency and amplitude, respectively, of the frictionless pendulum. $\phi$ is a stochastic phase shift. This task was inspired by the task from Wehenkel et al. (2025). However, it differs significantly and was adjusted to test our claim that increasing the number of posterior observations will better constrain the parameters.
Misspecification We synthesise a time calibration error in the instrumentation that causes the instrument to take 200 measurements every 0.075 seconds instead of the simulated 0.05 seconds. There is a significant probability that the misspecified point will appear ID due to the mild misspecification.
Neural Statistic Estimation In this example, each of the data points is a single observation. The neural statistic estimator is an embedding of the full pendulum time series into a lower-dimensional representation of the data. Following Chen et al. (2021), we use the Shannon-Jensen InfoMax objective (Hjelm et al., 2019) to target sufficient neural statistics, $\iota_{\bm{\omega}}$, to encode the high-dimensional data. This objective function maximises the mutual information between $\bm{x}_{\rm sim}$ and $\bm{\theta}$ using a discriminator network.
Figure 5: Results for the Pendulum task. We conclude that RVNP and its variants can recover robust posterior inference in amortised simulation-based inference. The hue in the middle plots indicates the error bar on the NPE and NNPE algorithms. For $\alpha$ nearest to 0 is best, with positive values representing underconfidence and negative values representing overconfidence. For the log-probability, higher values are better. For NMSE, lower values are better.
Results. We display the results for the Pendulum task in Figure 5 and provide an example posterior in the $N_{\rm obs}=1000$ case in Figure 2. Finally, we display the simulated observations together with observed points in Figure 9. In this task, the misspecification is geometrically significantly more complex. Moreover, due to the nature of the misspecification, most of the points will appear ID. Due to the complexity of the misspecification, NNPE struggles significantly. However, RVNP and RVNP-Global recover robust posterior inference for this task. In particular, RVNP is very well calibrated after the $N_{\rm obs}=100$ point and recovers robust posterior inference across a broad range in parameter space. Due to the complexity of the error model, the tuned variant, RVNP-T, does not see the same degree of performance increase. However, it does perform better than both NPE and NNPE.
4.5 Task D: Spectra Task - Real Gaia BP/RP Data
The third data release of the European Space Agency’s Gaia telescope (Gaia Collaboration et al., 2016) contain over 220 million flux-calibrated, low-resolution, optical stellar spectra. These spectra are measured by two instruments, the “Blue Photometer” (BP, 330-680 nm coverage in wavelength) and the “Red Photometer” (RP, 640-1050 nm). The processed and calibrated (De Angeli, F. et al. 2023; Montegriffo, P. et al. 2023) BP/RP (XP) spectra from Gaia DR3 are low-resolution, contaminated spectra which have multiple difficult systematics to overcome (Huang et al., 2024). However, the XP spectra are expected to contain significant information about different stellar parameters (Witten et al., 2022) and robust posterior inference using stellar evolution simulators conditional on the Gaia XP spectra would provide an efficient method for understanding the Milky Way.
Simulator We use the MIST Choi et al. (2016) stellar evolution models to generate stellar parameters compatible with high Galactic latitudes and map each of the effective temperature, log-surface gravity, and metallicity to a medium-resolution synthetic Castelli & Kurucz (2004) model spectral energy distribution. This defines the simulator relating $\bm{\theta}=(T_{\rm eff},\log g,\rm{[Fe/H]})$ to $\bm{x}_{\rm sim}$. We cut the spectra in their native resolution between 330-1050 nm to define a $301$ dimensional vector which overlaps with the Gaia XP spectra wavelength range but at a significantly different resolution.
Embedding We pre-train an NSE estimator using the same method as described in the pendulum synthetic task.
Misspecification: We can view the problem of inferring stellar parameters using real Gaia XP spectra and synthetic stellar evolution models as a misspecification problem. The Gaia XP spectra are expected to have two significant differences from the synthetic simulation model. Firstly, the XP spectra are lower resolution than the synthetic spectra and, therefore, contain less information than high-resolution spectra from the simulator. Secondly, there are inherent systematic errors in the processing of the XP spectra that cause a simulation gap even if we know the spectroscopic parameters using traditional high-resolution methods. For the real dataset, we target high Galactic latitudes due to the minimal impact of photometric extinction in these regions and a relatively homogenous population of main-sequence stars (O’Callaghan et al., 2024). We select all stars with absolute Galactic latitude $|b|>80^{\circ}$ that have valid LAMOST (Wang, 2022) spectroscopically determined stellar parameters. These spectroscopically determined stellar parameters will act as ground truth for our experiment, but we should note that they have their own errors arising from the spectroscopic determination. Furthermore, we select all Gaia recommended quality cuts and choose stars with confident distance estimates between $300$ and $700$ pc. This leaves us with a dataset of size $N_{\rm obs}=1053$.
Misspecification Summary. For those not familiar with stellar astronomy, the observed dataset is generated from real data with ground truth values that were selected so that measurement error and external factors have a minimal impact on the data, helping us better isolate the model misspecification.
Results. We display the results for the Spectra task in Figure 5 and provide an example posterior in the $N_{\rm obs}=1000$ case in Figure 2. Finally, we display the simulated observations together with observed points in Figure 10. In this task, the misspecification is geometrically very complex due to both the image of the neural statistic and the complexity of stellar evolution. Many of the points appear ID relative to the simulated points in this task. We naively applied the neural statistic, so that the neural statistic knows nothing about the structure of the Gaia XP spectra. We find that the neural statistic struggles to identify the metallicity parameter, most likely due to the neural statistic fitting for high-resolution features in the synthetic spectra.
Due to the complexity of the misspecification, NNPE struggles significantly. However, RVNP and RVNP-Global recover robust posterior inference for this task. In particular, RVNP-Global is very well calibrated after the $N_{\rm obs}=10$ point and recovers robust posterior inference across a broad range in parameter space. In this task, we displayed the tuned variant of RVNP-T and found that it struggled significantly to recover robust posterior inference. It is worth noting that we chose a poor prior for this problem, and we hypothesise that this is why RVNP-Global performs better than RVNP-Global, as the neural network covariance matrix accesses many unseen parts of parameter space.
Figure 6: Results for the Spectra task. We conclude that RVNP and its variants can recover robust posterior inference in amortised simulation-based inference. The hue in the middle plots indicates the error bar on the NPE and NNPE algorithms. For $\alpha$ nearest to 0 is best, with positive values representing underconfidence and negative values representing overconfidence. For the log-probability, higher values are better. For NMSE, lower values are better.
5 Related work
Misspecification in SBI Model misspecification is understood in likelihood-based methods (Davison 2003, pages 147–148); however, a systematic theory of likelihood-free methods is lacking. Misspecification in SBI has been studied in the context of ABC (Frazier et al. 2020; Bharti et al. 2022; Fujisawa et al. 2021), BSL (Frazier et al., 2024), generalized Bayesian inference (Dellaporta et al., 2022), and neural conditional density estimation (Ward et al. 2022; Kelly et al. 2024; Huang et al. 2023; Elsemüller et al. 2025; Mishra et al. 2025; Schmitt et al. 2024; Wehenkel et al. 2025). In this paper, we focused on neural-based methods, where empirically it has been shown that SBI struggles under model misspecification (Cannon et al. 2022; Schmitt et al. 2024). Robust posterior recovery under model misspecification is essential for the success of SBI, and different methods have emerged to mitigate against it. Kelly et al. (2025) identifies three main strategies currently used to account for model misspecification in SBI: robust summary statistics, generalised Bayesian inference, and error modelling and adjustment parameters. Most of the early solutions in robust SBI were intended for a single data set (Ward et al. 2022; Kelly et al. 2024; Huang et al. 2023).
Robust amortised SBI Recently, robust amortised neural SBI has been addressed using optimal transport for domain shifts when a calibration set exists (Wehenkel et al., 2025), using unsupervised domain adaptation (Elsemüller et al., 2025), and consistency losses regularisation (Mishra et al., 2025). Moreover, Glöckler et al. (2023) proposes regularisation techniques to increase the robustness of the learnt posterior against adversarial attacks. Of these methods, only the consistency loss method targets both the likelihood and the posterior. The noisy neural posterior estimation from Ward et al. (2022) can also be viewed as an amortised SBI method, where a pre-defined error model is used during training to corrupt the simulations.
Variational methods in SBI Variational methods have been used in multiple capacities to date. Wiqvist et al. (2021) introduced Sequential Neural Posterior and Likelihood Approximation, which proposes using variational inference (VI) to speed up the inference of likelihood-based methods, similar to the likelihood-based Bayesian approach to VI. Glöckler et al. (2022) introduces a framework that uses VI for simulation-based inference by using a pre-trained likelihood (or likelihood-ratio) and learn the posterior using VI, then refining the posterior using sampling importance resampling (Rubin, 1987). Nautiyal et al. (2024) introduces a generative modelling approach based on the variational inference framework and learns an encoder-decoder model in terms of latent variables. They introduce a latent variable that can account for complex structures and dependencies in the simulator model. Simons et al. (2022) propose a simulation-based inference algorithm that iteratively updates particles to more match the posterior in a variational likelihood-free gradient descent manner.
6 Discussion
In this paper, we introduced RVNP(-T), a robust amortised Bayesian inference method for simulation-based inference that jointly infers the simulation-to-reality gap and the amortised posterior using an importance-weighted autoencoder framework. This is the first case of using variational inference for robust simulation-based inference. Moreover, it is the first approach that does not rely on tuning hyperparameters of the loss function to recover robust posterior inference. We argue that this is an important step toward reliable posterior inference in amortised SBI. Previous work’s reliance on tuning parameters or misspecification priors to control the robustness of the posterior implies that there is an unknown parameter that controls the posterior inference in a way that may be nontrivial for real inference tasks. Furthermore, the error model is interpretable, overcoming the issues with unsupervised domain adaptation.
6.1 Neural statistic estimators
In this paper, we adopt a neural statistic estimator (NSE) before training the neural likelihood surrogate. There is an architectural limitation from the necessity of learning the likelihood proxy in RVNP because learning the NSE simultaneously with the likelihood is not a well-posed problem (Brehmer & Cranmer, 2020). However, we argue that in amortised SBI, this is not a significant problem. Usually, amortised SBI requires a fixed simulation budget up front, and in neural posterior estimation (NPE), the neural conditional density estimator is trained on those simulations. When training an NSE for an NPE task, usually the NSE is trained simultaneously with the posterior proxy. However, this objective function has no theoretical guarantees of the sufficiency of the statistic (Chen et al., 2021). In this paper, we chose to pre-train the NSE using the InfoMax objective (Hjelm et al., 2019) because of its theoretical guarantees of sufficiency.
We also considered hand-crafted statistics in this paper, as the robustness guarantees occur during the final variational inference step. This allows RVNP to include problems where an understanding of the error model is a significant part of the inference task. Such as in situations where the chosen summary statistics are in units known to the scientist, and they wish to fit for discrepancies between the model and the data.
6.2 Connection to domain adaptation
Unsupervised domain adaptation couples posterior learning to adapting the neural statistics. In this paper, we saw that for some tasks, learning a simple covariance matrix parameterised by a neural network on a subset of data does not necessarily generalise well to other areas of parameter space, even with strong inductive bias. Complex domain adaptation in a nonlinear fashion will give rise to infinite solutions that can account for the simulation-to-reality gap, and even in simple inductive bias error models throughout this paper, it highlights that the domain adaptation may cause serious issues if many of the points lie ID. A separate body of work should discuss error modelling using calibration to compare with the results of Wehenkel et al. (2025).
6.3 Prior misspecification
A type of misspecification not explicitly addressed in this paper is if the prior $p(\bm{\theta})$ is itself misspecified. RVNP should have the same limitations as likelihood-based prior misspecification, except with two crucial differences. Any prior chosen must be defined within the support of the prior used to train the likelihood. Furthermore, changing the prior distribution may make the likelihood proxy learn different parts of the parameter space suboptimally. In the Spectra task, we defined an incorrect prior relative to the observed data, which occurred unintentionally due to the selection effects of the real data. It is worth considering the pendulum task for this problem. The OOD points appear OOD under this misspecification because the effect of the misspecification makes the points appear they have a higher fundamental frequency. However, the information about the misspecification is only available to us because we assume that the prior is well-specified. We would lose the robustness of the inference if we assumed a wider prior on the fundamental frequency, causing the synthetic DGP to cover the observed data points.
6.4 Higher dimensional problems
Our method inherits the limitations of both neural likelihood estimation and amortised variational inference. It is highly desirable to use NSE in higher-dimensional problems to reduce the dimension of the data to help the computational capabilities of the normalising flow on the likelihood estimation. In general, RVNP should scale well to higher-dimensional problems, but may struggle to compete with domain adaptation methods for more complex, multimodal data. Future work will look at the limitations of using pre-training the NSE and learning the likelihood proxy on the embedded summary statistics.
Broader Impact Statement
Author Contributions
Acknowledgments
Appendix A Appendix
A.1 Misspecified Observations Overlaying Training Simulations
Here, we display the simulated samples and the misspecified points together for each task.
Figure 7: The simulated samples and the misspecified points for the CS Task.
Figure 8: The simulated samples and the misspecified points for the SIR Task.
Figure 9: The simulated samples and the misspecified points for the Pendulum Task.
Figure 10: The simulated samples and the misspecified points for the Spectra Task.
A.2 RVNP Training Pipeline
We provide an overview of the RVNP training pipeline in the following algorithms.
Algorithm 1 RVNP Full Pipeline
1:Input: Flags train_flow,train_variational,tune_posteriortrain\_flow,train\_variational,tune\_posterior
2:Input: All datasets and models
3:if train_flowtrain\_flow then
4:  TrainNormalizingFlow()
5:end if
6:if train_variationaltrain\_variational then
7:  TrainVariationalPosteriorAndErrorModel()
8:end if
9:if tune_posteriortune\_posterior then
10:  PosteriorTuning()
11:end if
Algorithm 2 Train Normalising Flow Likelihood
1:Input: Simulator dataset $D=\{(\mathbf{x}_{\rm sim},\bm{\theta})\}$, normalizing flow $p_{\Psi}$
2:Input: Number of epochs $E_{\rm flow}$, optimizer $\text{Optimizer}_{\Psi}$
3:for epoch =1=1 to $E_{\rm flow}$ do
4:  Sample minibatch $B_{\rm sim}\subset D$
5:  Compute NLE loss:
$\mathcal{L}_{\Psi}=-\mathbb{E}_{B_{\rm sim}}[\log p_{\Psi}(\mathbf{x}_{\rm sim}\mid\bm{\theta})]$
6:  Update $\Psi\leftarrow\text{Optimizer}_{\Psi}(\nabla_{\Psi}\mathcal{L}_{\Psi})$
7:end for
Algorithm 3 Train Variational Posterior and Error Model (RVNP)
1:Input: Observed dataset $\mathcal{O}$, normalizing flow $p_{\Psi}$, posterior $p_{\phi}$, error model network $\bm{\xi}_{\alpha}$
2:Input: Number of epochs $E_{\rm var}$, MC samples $K$, optimizers $\text{Optimizer}_{\phi},\text{Optimizer}_{\alpha}$
3:for epoch =1=1 to $E_{\rm var}$ do
4:  for each $\mathbf{x}_{\rm obs}^{(i)}\in\mathcal{O}$ do
5:   Sample $\bm{\theta}^{(l)}\sim p_{\phi}(\bm{\theta}\mid\mathbf{x}_{\rm obs}^{(i)}),\;l=1..K$
6:   For each $\bm{\theta}^{(l)}$, sample $\mathbf{x}_{\rm sim}^{(l,m)}\sim p_{\Psi}(\mathbf{x}_{\rm sim}\mid\bm{\theta}^{(l)})$
7:   Compute IWAE variational loss:
$\mathcal{L}_{V}^{(i)}=-\log\frac{1}{K}\sum_{l=1}^{K}\frac{\mathbb{E}_{\mathbf{x}_{\rm sim}^{(l,m)}}\big{[}p_{\bm{\xi}_{\alpha}}(\mathbf{x}_{\rm obs}^{(i)}\mid\mathbf{x}_{\rm sim}^{(l,m)})\big{]}\,p(\bm{\theta}^{(l)})\,p(\bm{\xi}_{\alpha}(\bm{\theta}^{(l)}))}{p_{\phi}(\bm{\theta}^{(l)}\mid\mathbf{x}_{\rm obs}^{(i)})}$
8:  end for
9:  Update $\phi,\alpha$ via $\text{Optimizer}_{\phi},\text{Optimizer}_{\alpha}$
10:end for
Algorithm 4 Posterior Tuning (RVNP-T)
1:Input: Simulator dataset $D$, posterior $p_{\phi}$, error model $\bm{\xi}_{\alpha}$ (fixed)
2:Input: Number of epochs $E_{\rm tune}$, optimizer $\text{Optimizer}_{\phi}$
3:for epoch =1=1 to $E_{\rm tune}$ do
4:  Sample minibatch $B_{\rm sim}\subset D$
5:  Compute NPE-style loss with fixed error model:
$\mathcal{L}_{\rm NPE}=-\mathbb{E}_{(\mathbf{x}_{\rm sim},\bm{\theta})\in B_{\rm sim}}\mathbb{E}_{p_{\bm{\xi}_{\alpha}}(\mathbf{x}_{\rm obs}\mid\mathbf{x}_{\rm sim})}[\log p_{\phi}(\bm{\theta}\mid\mathbf{x}_{\rm sim})]$
6:  Update $\phi\leftarrow\text{Optimizer}_{\phi}(\nabla_{\phi}\mathcal{L}_{\rm NPE})$
7:end for
A.3 Training Procedure
In every task, RVNP is defined using a variational posterior model based on a rational quadratic spline (RQS) flow with $B=10$ and $15$ knots. The depth of the flow is set to $5$ layers, while the neural network conditioner has a hidden block dimension of $52$. The flow dimension is the dimension of $\theta_{\text{dim}}$, and the conditioning dimension corresponds to that of $x_{\text{sim}}$. The simulator flow has the same architecture as the input and output dimensions swapped.
The importance weighted autoencoder objective was trained with a batch size of $1024$ over $500$ iterations, using the Adam optimizer with learning rate $10^{-3}$, momentum term $\beta_{1}=0.9$, $\epsilon=10^{-8}$, weight decay $10^{-5}$, gradient clipping at $10.0$, and a warmup schedule of $1000$ steps. Early stopping is applied with a patience of $100$ iterations, and $10\%$ of the data is reserved for validation. In the forward modelling of the posterior and the simulator, $K_{\text{obs samples}}=30$ is the number of samples used in the importance weighting. In each task, the simulator was trained using the same optimiser parameters but using the maximum likelihood loss.
Neural Statistic Estimator To train the neural statistic estimator on the InfoMax objective, we adopt a neural statistic and a discriminator model. We use the same optimiser parameters and batch size for the main training routine. All models are implemented in Equinox and trained using JAX. The encoder outputs a deterministic latent representation $z$ without variational sampling, and the discriminator maximises the mutual information between $z$ and $\bm{\theta}$. The spectra encoder uses one-dimensional convolutional feature extraction and global attention modelling using a Conformer block. The hidden dimension for both the embeddings and the discriminator is 100. We describe the algorithm in 5.
Algorithm 5 Spectra Encoder Forward Pass
0: Input sequence $x\in\mathbb{R}^{C\times L}$
1:xGELU(Conv1($)x\leftarrow\text${GELU}($\text${Conv1}(x))
[…]
7:xGELU(fc_hidden($)x\leftarrow\text${GELU}($\text${fc\_hidden}(x))
[…]
9:return  $z$
The pendulum encoder follows a similar structure but uses a single convolutional layer followed by a Conformer block. (Algorithm 6).
Algorithm 6 Pendulum Encoder Forward Pass
0: Input sequence $x\in\mathbb{R}^{C\times L}$
1:xGELU(Conv1($)x\leftarrow\text${GELU}($\text${Conv1}(x))
[…]
8:return  $z$
The discriminator is a simple multilayer perceptron (MLP) that takes the concatenation of the latent embedding $z$ and conditioning variable $\theta$ as input, and outputs a scalar logit (Algorithm 7).
Algorithm 7 Discriminator Forward Pass
0: Latent embedding $z\in\mathbb{R}^{d_{z}}$, condition vector $\theta\in\mathbb{R}^{d_{\theta}}$
[…]
2:xReLU(fc1($)x\leftarrow\text${ReLU}($\text${fc1}(x))
3:xReLU(fc2($)x\leftarrow\text${ReLU}($\text${fc2}(x))
[…]
5:return  logitlogit
InfoMax Loss Function
To train the encoders, we maximise the mutual information (MI) between the latent embeddings $z$ and the conditioning variables $\theta$. Algorithm 8 summarises the loss function.
Algorithm 8 InfoMax (Shannon) Loss Computation
0: Input batch $x\in\mathbb{R}^{B\times L}$, real batch $x_{\text{real}}$, condition vectors $\theta\in\mathbb{R}^{B\times d_{\theta}}$, encoder $E(\cdot)$, discriminator $D(\cdot)$, number of shuffles $S$
1: Sample randomness keys for encoder and discriminator
[…]
4: Compute joint discriminator outputs: $l_{\text{joint}}\leftarrow D(z,\theta)$
5: Initialise marginal loss accumulator
6:for $s=1\dots S$ do
7:  Generate random permutation $\pi_{s}$ of $\{1,\dots,B\}$
8:  $\theta_{\text{shuffled}}\leftarrow\theta[\pi_{s}]$
9:  $l_{\text{marginal}}^{(s)}\leftarrow D(z,\theta_{\text{shuffled}})$
10:  Accumulate: $m^{(s)}\leftarrow-\mathrm{softplus}(l_{\text{marginal}}^{(s)})$
11:end for
12: Compute joint term: $J\leftarrow-\mathrm{softplus}(-l_{\text{joint}})$
13: Compute marginal term: $M\leftarrow\frac{1}{S}\sum_{s=1}^{S}m^{(s)}$
14: Estimate MI lower bound: $\widehat{I}(z;\theta)\leftarrow\mathbb{E}[J]+\mathbb{E}[M]$
15: Shannon loss: $\mathcal{L}_{\text{Shannon}}\leftarrow-\widehat{I}(z;\theta)$
16:return  $\mathcal{L}_{\text{Shannon}}$

### v2

@@
-Robust variational neural posterior estimation for simulation-based inference
+Misspecification-robust amortised simulation-based inference using variational methods
 Abstract
 Recent advances in neural density estimation have enabled powerful simulation-based inference (SBI) methods that can flexibly approximate Bayesian inference for intractable stochastic models.
-Although these methods have demonstrated reliable posterior estimation when the simulator accurately represents the underlying data generative process (GDP), recent work has shown that they perform poorly in the presence of model misspecification.
-This poses a significant problem for their use on real-world problems, due to simulators always misrepresenting the true DGP to a certain degree.
+Although these methods have demonstrated reliable posterior estimation when the simulator accurately represents the underlying data generative process (DGP), recent work has shown that they perform poorly in the presence of model misspecification.
+This poses a significant issue for their use in real-world problems, due to simulators always misrepresenting the true DGP to a certain degree.
 In this paper, we introduce robust variational neural posterior estimation (RVNP), a method which addresses the problem of misspecification in amortised SBI by bridging the simulation-to-reality gap using variational inference and error modelling.
-We test RVNP on multiple benchmark tasks, including using real data from astronomy, and show that it can recover robust posterior inference in a data-driven manner without adopting tunable hyperparameters or priors governing the misspecification.
+We test RVNP on multiple benchmark tasks, including using real data from astronomy, and show that it can recover robust posterior inference in a data-driven manner without adopting hyperparameters or priors governing the misspecification influence.
 1 Introduction
-Simulator models are ubiquitous in many areas of the natural sciences and engineering, enabling researchers to approximate complex real-world data-generating processes (DGP) using physically grounded forward models.
+Simulator models are ubiquitous in many areas of the natural sciences and engineering, enabling researchers to approximate complex real-world data generative processes (DGP) using physically grounded forward models.
 However, these simulators are often computationally expensive, non-differentiable, and lack closed-form likelihoods, making traditional inference methods inapplicable.
@@
 As a result of intractability and the computational expense of running simulations, solving the inverse problem of inferring simulator parameters from observed data $\bm{x}_{\rm obs}$ poses a significant challenge.
-Simulation-based inference (SBI, Cranmer et al.
-2020) or likelihood-free inference provides methods to approximately infer the posterior distribution of the simulator parameters conditioned on observed data.
-A range of SBI methods have emerged to solve the likelihood-free inference problem, beginning with traditional approaches such as approximate Bayesian computation (ABC; Rubin 1984; Beaumont et al.
-2002) and Bayesian synthetic likelihood (BSL, Price et al.
-2018).
-Recent work has introduced methods based on neural density estimation, such as neural posterior estimation (NPE, Papamakarios & Murray 2016; Lueckmann et al.
-2017; Greenberg et al.
-2019), neural likelihood estimation, (NLE, Lueckmann et al.
-2019; Papamakarios et al.
-2019), neural ratio estimation (NRE, Izbicki et al.
-2014; Cranmer et al.
-2016; Hermans et al.
-2020; Durkan et al.
-2020), and diffusion-based methods (Glöckler et al., 2024).
+Simulation-based inference (SBI, Cranmer_2020) or likelihood-free inference provides methods to approximately infer the posterior distribution of the simulator parameters conditioned on observed data.
+A range of SBI methods have emerged to solve the likelihood-free inference problem, beginning with traditional approaches such as approximate Bayesian computation (ABC; 10.1214/aos/1176346785; abc_genetics) and Bayesian synthetic likelihood (BSL, Price02012018).
+Recent work has introduced methods based on neural density estimation, such as neural posterior estimation (NPE, 2016arXiv160506376P; NIPS2017_addfa9b7; greenberg2019automaticposteriortransformationlikelihoodfree), neural likelihood estimation (NLE, pmlr-v96-lueckmann19a; papamakarios2019sequentialneurallikelihoodfast), neural ratio estimation (NRE, pmlr-v33-izbicki14; cranmer2016approximatinglikelihoodratioscalibrated; hermans2020likelihoodfreemcmcamortisedapproximate; pmlr-v119-durkan20a), and modelling the joint distribution of data and simulation parameters (DBLP:conf/icml/GlocklerDWWM24).
 SBI methods can be categorised into amortised and non-amortised inference methods.
-In the context of neural SBI, non-amortised methods such as sequential neural posterior estimation (SNPE; Greenberg et al.
-2019), sequential neural likelihood estimation (SNLE; Papamakarios et al.
-2019), and sequential neural ratio estimation (SNRE; Hermans et al.
-2020) target a single posterior conditioned on fixed data, adapting their inference procedure with each simulation round.
+In the context of neural SBI, non-amortised methods such as sequential neural posterior estimation (SNPE; 2016arXiv160506376P; NIPS2017_addfa9b7; greenberg2019automaticposteriortransformationlikelihoodfree), sequential neural likelihood estimation (SNLE; papamakarios2019sequentialneurallikelihoodfast), and sequential neural ratio estimation (SNRE; hermans2020likelihoodfreemcmcamortisedapproximate) target a single posterior conditioned on fixed data, adapting their inference procedure with each simulation round.
 After an up-front simulation budget, amortised methods aim to learn a global inference model over a given prior, making them well-suited for scenarios where repeated or scalable inference is required.
-In this paper, we focus on amortised SBI methods for inferring posterior distributions for any observation within the support of the simulator model.
-SBI methods have been widely used in fields such as astronomy (Mishra-Sharma & Cranmer, 2022), particle physics (The Atlas Collaboration, 2025), cosmology (Lemos et al.
-2023; Zeghal et al.
-2024), and neuroscience (Oesterle et al.
-2020; Hashemi et al.
-2024), to name a few.
-However, recent work has shown that they can yield overconfident posterior approximations (Hermans et al., 2022) and suffer significantly when the true DGP does not lie within the family of distributions defined by the statistical model (Cannon et al.
-2022; Schmitt et al.
-2024), known as model misspecification.
-Model misspecification may be caused by a variety of factors, such as contamination in the data or unaccounted-for physical processes in the modelling that can lead to overconfident posteriors (Hermans et al., 2022).
-This discrepancy between the simulated data and the real observations is known as the simulation-to-reality gap (Miglino et al., 1995) or simulation gap.
+In this paper, we focus on SBI methods for inferring posterior distributions that are amortised over a dataset of observations.
+SBI methods have been widely used in fields such as astronomy (PhysRevD.105.063017), particle physics (atlascollaboration2024implementationneuralsimulationbasedinference), cosmology (Lemos_2023; zeghal2024simulationbasedinferencebenchmarklsst), and neuroscience (10.7554/eLife.54997; 2024MLS&T...5c5019H), to name a few.
+However, recent work has shown that they can yield overconfident posterior approximations (2021arXiv211006581H) and suffer significantly when the true DGP does not lie within the family of distributions defined by the statistical model (cannon2022investigating; schmitt2024detecting), known as model misspecification.
+Model misspecification may be caused by a variety of factors, such as contamination in the data or unaccounted-for physical processes in the modelling that can lead to overconfident posteriors (2021arXiv211006581H).
+This discrepancy between the simulated data and the real observations is known as the simulation-to-reality gap (miglino1995evolving) or simulation gap.
 Methods for mitigating against misspecification in neural SBI have done so mainly by addressing the simulation-to-reality gap.
-This is based on the assumption that misspecification appears as a divergence-based discrepancy between the true DGP $p^{*}(\bm{x}_{\rm obs})$ and the distribution described by the simulator model $p(\bm{x}_{\rm sim})=\mathbb{E}_{p(\bm{\theta})}[p(\bm{x}_{\rm sim}|\,\bm{\theta})]$, under the prior distribution $p(\bm{\theta})$.
-Robust SBI methods usually address the simulation-to-reality gap through error modelling and adjustment parameters (Ward et al.
-2022; Frazier & Drovandi 2021; Kelly et al.
-2024), domain adaptation approaches (Huang et al.
-2023; Swierc et al.
-2024; Elsemüller et al.
-2025; Mishra et al.
-2025), or generalized Bayesian inference (Dellaporta et al., 2022).
+This is based on the assumption that misspecification appears as a divergence-based discrepancy between the true DGP $p^{*}(\bm{x}_{\rm obs})$ and the marginal distribution described by the simulator model $p(\bm{x}_{\rm sim})=\mathbb{E}_{p(\bm{\theta})}[p(\bm{x}_{\rm sim}|\,\bm{\theta})]$, under the prior distribution $p(\bm{\theta})$.
+Robust SBI methods often address the simulation-to-reality gap through error modelling and adjustment parameters (ward2022robust; frazier2021robust; kelly2024misspecificationrobust), domain adaptation approaches (huang2023learning; swierc2024domain; elsemüller2025doesunsuperviseddomainadaptation; mishra2025robust), or generalised Bayesian inference (dellaporta2022robust).
 The success of most of these methods relies on the observed points appearing as out-of-distribution (OOD) with respect to the simulated observations.
-However, recent work has underscored the importance of within-distribution (ID) points in a misspecified SBI (Schmitt et al.
-2024;Frazier et al.
-2024;Elsemüller et al.
-2025) as the errors in the model may still produce summary statistics which lie ID relative to the simulations.
-Wehenkel et al.
-(2025) shows that using a reliable calibration set can aid towards robust amortised SBI under such modelling errors.
+However, recent work has underscored the importance of within-distribution (ID) points in a misspecified SBI (schmitt2024detecting; Frazier26092024; elsemüller2025doesunsuperviseddomainadaptation), as the errors in the model may still produce summary statistics which lie ID relative to the simulations.
+wehenkel2025addressing showed that using a reliable calibration set can aid towards robust amortised SBI under such modelling errors.
 Often, a reliable calibration set will not exist, making such problems highly difficult to solve.
-Recently, unsupervised domain adaptation (UDA) methods have been implemented in robust amortised SBI using Maximum Mean Discrepancy and domain-adversarial neural networks (Elsemüller et al., 2025), and using consistency loss regularisation (Mishra et al., 2025).
+Recently, unsupervised domain adaptation (UDA) methods have been implemented in robust amortised SBI using maximum mean discrepancy (MMD), domain-adversarial neural networks (elsemüller2025doesunsuperviseddomainadaptation), and consistency loss regularisation (mishra2025robust).
 As amortised SBI looks to construct general posteriors for a range of observations, it is natural to consider the misspecification problem for situations involving many observations where all points appear OOD, or when a significant number of points appear OOD.
-Despite their success in robust SBI, existing methods for robust SBI encounter issues in the context of robust amortised SBI.
-In particular, the error modelling and correction parameter approaches scale poorly to amortised Bayesian inference due to their dependence on an MCMC sampling step.
-On the other hand, they benefit from their Bayesian formulation, particularly through the connection between hyperparameter choice and Bayesian prior adoption (Ward et al.
-2022; Frazier & Drovandi 2021; Kelly et al.
-2024).
-Domain adaptation methods scale more favourably to amortised SBI, but come at the cost of a non-Bayesian interpretation of the domain adaptation hyperparameters, a lack of interoperability of the domain adaptation (Elsemüller et al., 2025), and a lack of clarity between the trade-off in the domain adaptation and the inference algorithm (Chen et al., 2021).
+Despite their success in robust SBI, existing methods encounter issues in the context of robust amortised SBI.
+In particular, the error modelling and correction parameter approaches scale poorly to amortised Bayesian inference due to their dependence on a Markov Chain Monte Carlo (MCMC) sampling step.
+On the other hand, they benefit from their Bayesian formulation, particularly through the connection between hyperparameter choice and Bayesian prior adoption (ward2022robust; frazier2021robust; kelly2024misspecificationrobust).
+Domain adaptation methods scale more favourably to amortised SBI, but come at the cost of a non-Bayesian interpretation of the domain adaptation hyperparameters, a lack of interoperability of the domain adaptation (elsemüller2025doesunsuperviseddomainadaptation), and a lack of clarity between the trade-off in the domain adaptation and the inference algorithm (chen2021neuralapproximatesufficientstatistics).
 Furthermore, it is not always desirable to use domain-adapted neural embedding statistics if expert knowledge on the summary embedding space is available, such as known sufficient statistics on a low-dimensional observation space in physically motivated units.
@@
 1.1 Our contributions
-We propose robust variational neural posterior estimation (RVNP) and its tuned variant (RVNP-T) that addresses the misspecification problem in amortized SBI by pre-training the simulator likelihood $p_{\bm{\Psi}}(\bm{x}_{\rm sim}|\,\bm{\theta})$ for parameters $\bm{\Psi}$, adopting a flexible error model $p_{\bm{\xi}(\bm{\theta})}(\bm{x}_{\rm obs}|\,\bm{x}_{\rm sim})$, and using an importance weighted autoencoder (Burda et al., 2015) scheme to maximize the evidence of the true data under the variational posterior $p_{\bm{\phi}}(\bm{\theta}|\,\bm{x}_{\rm obs})$ for the parameters $\bm{\xi}$ and $\bm{\phi}$.
-The variational posterior is the main objective of the RVNP algorithm.
-The posterior $p_{\bm{\phi}}(\bm{\theta}|\,\bm{x}_{\rm obs})$ is optionally tuned on the adapted synthetic DGP to return the RVNP-T posterior distribution.
-The error modelling we choose to adopt can be seen as a neural adaptation of the mean adjustment and covariance inflation from Frazier & Drovandi (2021).
-The main claim of our paper is that robust variational neural posterior estimation (RVNP) and its tuned variant can recover robust amortised posterior inference under misspecification by bridging the simulation-to-reality gap using error modelling.
-The error model parameters are adapted in a data-driven way when we have many observations.
+Problem setting: Given a dataset $\mathcal{O}=\{\bm{x}_{\rm obs}^{(i)}\}_{i=1}^{N_{\rm obs}}$ where each observation corresponds to a different unknown parameter $\bm{\theta}^{*(i)}$, we seek to infer all posterior distributions $\{p(\bm{\theta}\,|\,\bm{x}_{\rm obs}^{(i)})\}_{i=1}^{N_{\rm obs}}$, when the simulator model is misspecified.
+Amortised SBI: In this paper, we refer to amortised SBI as amortising the Bayesian inference over all data points in $\mathcal{O}$.
+This definition is in line with amortised variational inference (margossian2023amortised).
+Our solution: We propose robust variational neural posterior estimation (RVNP) and its tuned variant (RVNP-T), methods that address the misspecification problem in amortised SBI.
+RVNP pre-trains a simulator likelihood $p_{\bm{\Psi}}(\bm{x}_{\rm sim}|\,\bm{\theta})$, adopts a flexible error model $p_{\bm{\xi}(\bm{\theta})}(\bm{x}_{\rm obs}|\,\bm{x}_{\rm sim})$, and uses an importance weighted autoencoder (burda2015importance) inference scheme to maximize the evidence of the true data under the variational posterior $q_{\bm{\phi}}(\bm{\theta}|\,\bm{x}_{\rm obs})$ for the parameters $\bm{\xi}$ and $\bm{\phi}$.
+The main claim of our paper is that RVNP and its tuned variant can recover robust amortised posterior inference under misspecification by bridging the simulation-to-reality gap using error modelling.
+The error model parameters are adapted in a data-driven way when we have many observations and the method requires no extra manually tuned misspecification-related hyperparameters or priors that govern the influence of the misspecification.
 We summarise our contributions as follows:
   1. 1.
-    We introduce RVNP, an amortised SBI method that uses a pre-trained simulator likelihood, an error model, and an importance-weighted autoencoder (Burda et al., 2015) scheme to return robust amortised posterior inference under misspecification without adopting tunable parameters or priors over the misspecification.
+    We introduce RVNP, an amortised SBI method that uses a pre-trained simulator likelihood, an error model, and an importance-weighted autoencoder (burda2015importance) scheme to return robust amortised posterior inference under misspecification without adopting user-tuned misspecification-related hyperparameters or priors.
 We also introduce RVNP-T, which tunes the final posterior using the simulator and the noise induced by the error model.
@@
   3. 3.
-    To our knowledge, this is the first example of using amortised variational autoencoders to address the misspecification problem in SBI.
-In amortised SBI, it is usually necessary to have a pre-defined up-front simulator sample size, which may change if the model is learnt in rounds.
-We assume that we have a fixed simulation budget and that the neural statistic embedding is either pre-trained or adopted by expert knowledge.
-The simulation-to-reality gap should inform our error model as we observed more data from the true DGP (Figure 1).
-We impose a strong inductive bias on the domain adaptation so that the error model can only inflate the synthetic DGP to account for the simulation gap, allowing us to have a form of model criticism.
-Cranmer et al.
-(2020) suggests augmenting the simulator with an error model to account for the simulation-to-reality gap.
-However, choosing an error model can be difficult and arbitrary and requires multiple independent trainings to test the different error models (Ward et al., 2022).
-We instead consider using an error model that is inferred by maximising the evidence lower bound jointly with inferring the posterior distribution of the parameters.
-Our method is well-suited to situations when the misspecified points appear, on average, OOD relative to the simulated points.
-Current unsupervised robust SBI methods rely on a clear simulation-to-reality gap driven by misspecification.
-In situations where the entire inference set lies within or very near the original simulated distribution, a strong calibration set is most likely necessary.
+    To our knowledge, this is the first example of using amortised variational inference to address the misspecification problem in SBI.
 Overview of Paper.
@@
 This embedding can represent fixed user-defined summary statistics, a pre-defined embedding, or a neural statistic estimator (NSE) where the parameters $\bm{\omega}$ are to be learnt.
-Lower-dimensional embeddings are important when dealing with high-dimensional data, but can come at the cost of information loss when the embedding is not a sufficient statistic of $\bm{x}_{\rm sim}$ for $\bm{\theta}$ (Blum et al., 2013).
+Lower-dimensional embeddings are important when dealing with high-dimensional data, but can come at the cost of information loss when the embedding is not a sufficient statistic of $\bm{x}_{\rm sim}$ for $\bm{\theta}$ (blum2013comparative).
+In this paper, we consider examples where we have expert-informed summary statistics, and neural embedding summary statistics.
 (a) CS posterior
@@
 The goal of amortised neural posterior estimation is to approximate the unknown posterior distribution $p^{*}(\bm{\theta}|\bm{x_{\rm obs}})$ for all $\bm{x}_{\rm obs}\in\mathcal{X}_{\rm obs}$.
-After choosing a conditional density estimation architecture $p_{\bm{\phi}}(\bm{\theta}|\,\bm{x}_{\rm obs})$ parametrized by $\bm{\phi}$, and an architecture for the neural statistic embedding, $\iota_{\omega}$, NPE (Papamakarios & Murray 2016; Lueckmann et al.
-2017; Greenberg et al.
-2019) fits for the parameters $\bm{\omega}$ and $\bm{\phi}$ by minimizing the the expected forward Kullback-Leibler (KL) divergence between analytic and approximate posterior
-$\mathcal{L}_{\rm NPE}({\bm{\phi},\bm{\omega}})$ $=\mathbb{E}_{p^{*}(x_{\rm obs})}\big{[}\mathbb{KL}[p(\bm{\theta}|\,\iota_{\bm{\omega}}(\bm{x}_{\rm obs}))||p_{\bm{\phi}}(\bm{\theta}|\,\iota_{\bm{\omega}}(\bm{x}_{\rm obs}))]\big{]}$ (1)
-$=\mathbb{E}_{p^{*}(x_{\rm obs})}\big{[}\mathbb{E}_{p(\bm{\theta}|\,\iota_{\bm{\omega}}(\bm{x}_{\rm obs}))}[-\log p_{\bm{\phi}}(\bm{\theta}|\,\iota_{\bm{\omega}}(\bm{x}_{\rm obs}))]\big{]},$
-where the expectation is over the unknown true data-generating distribution $p^{*}(\bm{x}_{\rm obs})$ and the second line follows because the true unknown posterior and its entropy do not depend on the trainable parameters.
-As noted in Schmitt et al.
-(2024), this amortised posterior objective function is not feasible as we may not have enough real data to approximate the expectation with respect to $p^{*}(\bm{x}_{\rm obs})$, and the true posterior is unknown and intractable.
-Instead, the unknown $p^{*}(\bm{x}_{\rm obs})$ is replaced with the marginal likelihood $p(\bm{x_{\rm obs}})=\int p(\bm{x}_{\rm obs}|\,\bm{\theta})p(\bm{\theta})d\bm{\theta}$.
-Under the model assumption we have $\bm{x}_{\rm obs}$ can be replaced with $\bm{x}_{\rm sim}$ and the objective becomes
-$\mathcal{L}(\bm{\phi},\bm{\omega})_{\rm NPE}:=-\mathbb{E}_{\bm{x}_{\rm sim},\bm{\theta}}[\log p_{\bm{\phi}}(\bm{\theta}|\,\iota_{\bm{\omega}}(\bm{x}_{\rm sim}))].$ (2)
+After choosing a conditional density estimation architecture $q_{\bm{\phi}}(\bm{\theta}|\,\bm{x}_{\rm obs})$ parametrized by $\bm{\phi}$, and an architecture for the neural statistic embedding, $\iota_{\omega}$, NPE (2016arXiv160506376P; NIPS2017_addfa9b7; greenberg2019automaticposteriortransformationlikelihoodfree) fits for the parameters $\bm{\omega}$ and $\bm{\phi}$ by minimizing the the expected forward Kullback-Leibler (KL) divergence between analytic and approximate posterior
+$\mathcal{L}_{\rm NPE}({\bm{\phi},\bm{\omega}})$ $=\mathbb{E}_{p^{*}(x_{\rm obs})}\big[\mathbb{KL}[p(\bm{\theta}|\,\iota_{\bm{\omega}}(\bm{x}_{\rm obs}))||q_{\bm{\phi}}(\bm{\theta}|\,\iota_{\bm{\omega}}(\bm{x}_{\rm obs}))]\big]$ (1)
+$=\mathbb{E}_{p^{*}(x_{\rm obs})}\big[\mathbb{E}_{p(\bm{\theta}|\,\iota_{\bm{\omega}}(\bm{x}_{\rm obs}))}[-\log q_{\bm{\phi}}(\bm{\theta}|\,\iota_{\bm{\omega}}(\bm{x}_{\rm obs}))]\big],$
+where the expectation is over the unknown true data-generating distribution $p^{*}(\bm{x}_{\rm obs})$.
+The second line follows from the first because the true unknown posterior and its entropy do not depend on the trainable parameters.
+This amortised posterior objective function is not realistic in most situations (schmitt2024detecting), as we rarely have enough real data to approximate the expectation with respect to $p^{*}(\bm{x}_{\rm obs})$, and the true posterior is unknown.
+Instead, the unknown $p^{*}(\bm{x}_{\rm obs})$ is replaced by the marginal likelihood $p(\bm{x_{\rm obs}})=\int p(\bm{x}_{\rm obs}|\,\bm{\theta})p(\bm{\theta})d\bm{\theta}$.
+Under this assumption that there is no misspecification, $\bm{x}_{\rm obs}$ can be replaced with $\bm{x}_{\rm sim}$ and the objective becomes
+$\mathcal{L}(\bm{\phi},\bm{\omega})_{\rm NPE}:=-\mathbb{E}_{\bm{x}_{\rm sim},\bm{\theta}}[\log q_{\bm{\phi}}(\bm{\theta}|\,\iota_{\bm{\omega}}(\bm{x}_{\rm sim}))],$ (2)
+where the expectation is over simulator input and output pairs.
 This is minimised with respect to the parameters $\bm{\omega}$ and $\bm{\phi}$.
 The success of this objective depends on the assumption that sampling from the evidence is equivalent to sampling from $p^{*}(\bm{x}_{\rm obs})$.
-On the other hand, neural likelihood estimation (NLE) learns a distribution maximising the conditional log-probability of the simulated data
+Neural likelihood estimation (NLE) learns a distribution maximising the conditional log-probability of the simulated data
 $\mathcal{L}(\bm{\Psi})_{\rm NLE}:=\mathbb{E}_{p(\bm{x}_{\rm sim}|\bm{\theta})p(\bm{\theta})}[\log p_{\bm{\Psi}}(\bm{x}_{\rm sim}|\bm{\theta})]$ (3)
 with respect to the flow parameters $\bm{\Psi}$ (where we dropped reference to $\iota_{\bm{\omega}}$), which is equivalent to minimising the KL divergence between the flow and the target distribution.
-The assumption of the learnt distribution being a likelihood identifies $\bm{x}_{\rm obs}$ with $\bm{x}_{\rm sim}$.
-However, one can also view the NLE objective as a surrogate training objective to approximate the sampling distribution given by the simulator, without any initial assumptions about the connection between true data and simulator output.
-In this sense, the NLE objective becomes a neural conditional prior proxy, and a likelihood can be introduced to relate the forward model to the true $\bm{x}_{\rm obs}$.
+Using neural embeddings for the summary statistics must be implemented in a two-stage manner because it cannot be trained in an end-to-end fashion along with the likelihood proxy (brehmer2020flows).
+Sampling from the posterior amounts to adopting either a Markov Chain Monte Carlo (MCMC) sampling method or a variational inference algorithm, which will increase the number of architecture choices and hyperparameter tuning, but can recover accurate posterior inference (glockler2022variational).
 2.3 Misspecification in SBI
-The simulator, $p(\bm{x}_{\rm sim}|\,\bm{\theta})$, is said to be misspecified if the true data-generating process does not fall within the family of distributions defined by the simulator on the support of the prior of $\bm{\theta}$.
-That is, $q^{*}\notin\{p(x_{\rm sim}|\,\bm{\theta})\,;\bm{\theta}\in\rm{supp}(p(\bm{\theta}))\}$ (Cannon et al., 2022).
-For single observations, the definition of misspecification was extended to using summary statistics (Kelly et al., 2024) by defining $\bm{b}(\bm{\theta})=\mathbb{E}_{p(\bm{x}_{\rm sim}\mid\bm{\theta})}[\iota_{\omega}(\bm{x}_{\rm sim})]$ and $\bm{b}_{j}=\mathbb{E}_{q^{*}(\bm{x}_{\rm obs})}[\iota_{\omega}(\bm{x}_{\rm obs}^{(j)})]$, for each $\bm{x}_{\rm obs}^{(j)}\in O$ where $O=\{\bm{x}_{\rm obs}^{(i)}\}_{i=1}^{N_{\rm obs}}$ is the set of observations each associated to a different, unknown true $\bm{\theta^{(j)}}$ value.
-Then the the simulator is misspecified if there is no $\{\bm{\theta}\}_{i=1}^{N_{\rm obs}}$ in the support of the prior for which $\bm{b}(\bm{\theta}^{(j)})=\bm{b}_{j}$ for each $j$.
-In this paper, we adopt the alternate, but similar, definition of misspecification provided in Wehenkel et al.
-(2025), which is more aligned with amortised SBI.
+The simulator, $p(\bm{x}_{\rm sim}|\,\bm{\theta})$, is said to be misspecified if the true data generative process does not fall within the family of distributions defined by the simulator on the support of the prior of $\bm{\theta}$.
+That is, $q^{*}\notin\{p(x_{\rm sim}|\,\bm{\theta})\,;\bm{\theta}\in\bm{\Theta})\}$ (cannon2022investigating).
+For single observations, the definition of misspecification can be extended to using summary statistics (kelly2024misspecificationrobust) by defining $\bm{b}(\bm{\theta})=\mathbb{E}_{p(\bm{x}_{\rm sim}\mid\bm{\theta})}[\iota_{\omega}(\bm{x}_{\rm sim})]$ and $\bm{b}_{j}=\mathbb{E}_{q^{*}(\bm{x}_{\rm obs})}[\iota_{\omega}(\bm{x}_{\rm obs}^{(j)})]$, for each $\bm{x}_{\rm obs}^{(j)}\in O$ where $O=\{\bm{x}_{\rm obs}^{(i)}\}_{i=1}^{N_{\rm obs}}$ is the set of observations each associated to a different, unknown true $\bm{\theta^{(j)}}$ value.
+The simulator is said to be misspecified if there is no $\{\bm{\theta}\}_{i=1}^{N_{\rm obs}}$ in the support of the prior for which $\bm{b}(\bm{\theta}^{(j)})=\bm{b}_{j}$ for each $j$.
+In this paper, we adopt the alternate, but similar, definition of misspecification provided in wehenkel2025addressing, which is more aligned with amortised SBI.
 They define a simulator to be misspecified if $\exists\mathcal{S}\subseteq\Theta\times\mathcal{X}_{\rm obs}:\forall(\bm{\theta},\bm{x}_{\rm obs})\in\mathcal{S}$
@@
 This definition does, however, ignore situations when the prior distribution is misspecified.
-Our method can handle prior misspecification more favourably through traditional Bayesian model iterations, provided the support of the proposed prior is a subset of the support of the original prior.
+Prior misspecification is addressed at the end of the paper in the discussion.
 2.4 Importance weighted autoencoders
-Importance-weighted autoencoders (IWAE) were inspired by the vanilla variational autoencoder (VAE, Kingma & Welling 2013), which introduces a latent space $\mathcal{Z}\subseteq\mathbb{R}^{m}$ and two distributions, $p_{\bm{\xi}}(\bm{x}|\,\bm{z})$ and $p_{\bm{\phi}}(\bm{z}|\,\bm{x})$, known as the decoder and the encoder respectively.
-The vanilla VAE objective aims to maximise the evidence of the data via the evidence lower bound
-$\log p(\mathbf{x})\;\geq\;\mathbb{E}_{p_{\bm{\phi}}(\mathbf{z}|\mathbf{x})}\left[\log p_{\bm{\xi}}(\mathbf{x}|\mathbf{z})\right]-\mathrm{KL}\!\left(p_{\bm{\phi}}(\mathbf{z}|\mathbf{x})\,\|\,p(\mathbf{z})\right)$ (5)
-with respect to $\bm{\xi}$ and $\bm{\phi}$, where $p(\bm{z})$ is a prior over the latent parameter.
+Importance-weighted autoencoders (IWAE) were inspired by the vanilla variational autoencoder (VAE, kingma2013auto), which introduces a latent space ${\Theta}\subseteq\mathbb{R}^{m}$ and two distributions, $p_{\bm{\xi}}(\bm{x}|\,\bm{\theta})$ and $q_{\bm{\phi}}(\bm{\theta}|\,\bm{x})$, known as the decoder and the encoder respectively.
+The vanilla VAE aims to maximise the evidence of the data using the evidence lower bound
+$\log p(\mathbf{x})\;\geq\;\mathbb{E}_{q_{\bm{\phi}}(\mathbf{\theta}|\mathbf{x})}\left[\log p_{\bm{\xi}}(\mathbf{x}|\mathbf{\theta})\right]-\mathrm{KL}\!\left(q_{\bm{\phi}}(\mathbf{\theta}|\mathbf{x})\,\|\,p(\mathbf{\theta})\right)$ (5)
+with respect to $\bm{\xi}$ and $\bm{\phi}$, where $p(\bm{\theta})$ is a prior over the latent parameter.
 However, using this bound has been shown to induce mode-seeking behaviour, leading to overly simplified representations and poor inference.
-Burda et al.
-(2015) introduced IWAE, which is based on a strictly tighter evidence lower bound derived from importance sampling.
+burda2015importance introduced IWAE, which is based on a strictly tighter evidence lower bound derived from importance sampling.
 The log-evidence for a single point is given as
-$\log p_{\bm{\theta}}(\textbf{x})$ $\approx\log\left[\frac{1}{k}\sum_{l=1}^{k}\frac{p_{\bm{\xi}}\left(\textbf{x},\textbf{z}^{(l)}\right)}{p_{\bm{\phi}}\left(\textbf{z}^{(l)}|\textbf{x}\right)}\right]={\mathcal{L}}^{\text{IWAE}}_{k}\left(\bm{\xi},\bm{\phi};\textbf{x}\right)$ (6)
-$\text{with}\quad\textbf{z}^{(l)}\sim p_{\bm{\phi}}\left(\textbf{z}|\textbf{x}\right),$ (7)
+$\log p_{\bm{\theta}}(\bm{x})$ $\approx\log\left[\frac{1}{k}\sum_{l=1}^{k}\frac{p_{\bm{\xi}}\left(\textbf{x},\bm{\theta}^{(l)}\right)}{q_{\bm{\phi}}\left(\bm{\theta}^{(l)}|\textbf{x}\right)}\right]={\mathcal{L}}^{\text{IWAE}}_{k}\left(\bm{\xi},\bm{\phi};\textbf{x}\right)$ (6)
+$\text{with}\quad\bm{\theta}^{(l)}\sim q_{\bm{\phi}}\left(\bm{\theta}|\textbf{x}\right),$ (7)
 which is a mass-covering objective that targets the evidence.
+We use the The doubly reparameterized gradient (DReG) estimator (tucker2018doubly) to avoid low signal-to-noise in the gradient estimation.
+cremer2017reinterpretingimportanceweightedautoencoders proposed using the implicit distribution given by importance sampling from the true posterior using the variational posterior as a proposal distribution.
+Given samples $\bm{\theta}_{2:N_{s}}:=\bm{\theta}_{2},...,\bm{\theta}_{N_{s}}$, they define the implicit distribution arising from importance sampling as
+$\tilde{q}_{IW}(\mathbf{\theta}|\,\mathbf{x},\bm{\theta}_{2:N_{s}}):=\frac{p(\bm{x},\bm{\theta})}{\frac{1}{N_{s}}\Big(\frac{p(\bm{x},\bm{\theta})}{q_{\mathbf{\phi}}(\mathbf{\theta}|\,\mathbf{x})}+\sum_{k=1}^{N_{s}}\frac{p(\bm{x},\bm{\theta}_{j})}{q_{\mathbf{\phi}}(\mathbf{\theta}_{j}|\,\mathbf{x})}\Big)},$ (8)
+which has the property that
+$q_{\text{SIR}}(\bm{\theta}|\,\bm{x}):=\mathbb{E}_{\mathbf{\theta}_{2:N_{s}}}[\tilde{q}_{IW}(\mathbf{\theta}|\,\mathbf{x},\bm{\theta}_{2:N_{s}})]\rightarrow p(\mathbf{\theta}|\,\mathbf{x}),\,\rm{when}\,N_{s}\rightarrow\infty$ (9)
+the expectation of the implicit distribution approaches the true posterior.
+Sampling from this distribution is detailed explicitly in cremer2017reinterpretingimportanceweightedautoencoders, and is done by sampling-importance-resampling (SIR).
 3 Method: RVNP
-Motivated by a desire to have a method that can recover robust amortized posterior inference with a fixed embedding space such that the simulation-to-reality gap is bridged with an interpretable, flexible error model, we extend the approaches of Ward et al.
-(2022) and Kelly et al.
-(2024) to amortized SBI, and use an importance weighted autoencoder (Burda et al., 2015) amortized variational inference scheme to define robust variational neural posterior estimation (RVNP) and its tuned variant (RVNP-T).
+Motivated by a desire to have a method that can recover robust amortised posterior inference with a fixed embedding space such that the simulation-to-reality gap is bridged with an interpretable, flexible, and data-driven error model, we extend neural likelihood estimation and the error modelling approaches for model misspecification (ward2022robust; kelly2024misspecificationrobust) to amortised SBI and propose robust variational neural posterior estimation (RVNP).
+RVNP builds on variational methods for solving the inverse problem under a learnt likelihood in SBI (glockler2022variational) by adopting an importance weighted autoencoder amortised variational inference scheme (burda2015importance) to define robust variational neural posterior estimation (RVNP) and its tuned variant (RVNP-T).
 In what follows, we drop the explicit embedding notation $\iota_{\omega}$ unless necessary.
-Our goal is to use a normalizing flow parametrized by $\bm{\phi}$, $p_{\bm{\phi}}(\bm{\theta}|\,\bm{x}_{\rm obs})$, that can approximate the true posterior distribution.
+Our goal is to use a normalising flow parametrized by $\bm{\phi}$, $q_{\bm{\phi}}(\bm{\theta}|\,\bm{x}_{\rm obs})$, that can approximate the true posterior distribution in an amortised manner.
 Algorithm 1 provides an overview of the RVNP and RVNP-T algorithms.
-We chose a fixed simulation budget of $N_{\rm sim}=100,000$ for each task, with $10\%$ retained for validation.
-RVNP builds on variational methods for solving the inverse problem under a learnt likelihood in SBI (Glöckler et al., 2022), and relies on a conditional neural density estimator to approximate the likelihood from a fixed budget of simulations from the simulator model.
-This overcomes the computational expense of using Hamiltonian Monte Carlo Neal (2011) or other Markov Chain methods to jointly infer the parameters of interest and the parameters of the error model.
-Furthermore, variational methods allow us to use more expressive models for the error model distribution.
-We assume a pre-training or fixing of the neural statistic embedding $\iota_{\omega}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{l}$ occurred.
 3.1 Generative model
-Throughout, we assume that the neural embedding or summary statistics are pre-trained and are not jointly learnt.
-The first step of RVNP is to train the normalising flow $p_{\bm{\Psi}}(\bm{x}|\,\bm{\theta})$ to approximate the likelihood from the generated samples from the simulator using the NLE objective (Equation 3).
-Once the normalizing flow has been trained, we assume that $p(\bm{x}_{\rm sim}|\,\bm{\theta})\approx p_{\bm{\Psi}}(\bm{x}_{\rm sim}|\,\bm{\theta})$ and include the surrogate in the forward model.
+Throughout, we assume a pre-training or fixing of the neural statistic embedding $\iota_{\omega}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{l}$ occurred.
+The first step of RVNP is to train the normalising flow $p_{\bm{\Psi}}(\bm{x}|\,\bm{\theta})$ to approximate the likelihood using the generated samples from the simulator by minimising the NLE objective (Equation 3).
+Once the normalising flow has been trained, we assume that $p(\bm{x}_{\rm sim}|\,\bm{\theta})\approx p_{\bm{\Psi}}(\bm{x}_{\rm sim}|\,\bm{\theta})$ and include the surrogate in the forward model.
 We assume that the true DGP can be modelled as
@@
 We can use the IWAE lower bound on the log-evidence (Equation 6) to derive the variational loss function for RVNP as
-$\mathcal{L}(\bm{\phi},\bm{\alpha})_{V}\approx-\sum_{i=1}^{N_{\rm obs}}\log\left[\frac{1}{K}\sum_{l=1}^{K}\frac{\mathbb{E}_{p_{\bm{\Psi}}(\mathbf{x}_{\rm sim}\mid\bm{\theta}^{(l)})}\!\left[\,p_{\bm{\xi}(\bm{\theta}^{(i)}\,;\,\bm{\alpha})}\!\left(\mathbf{x}_{\rm obs}^{(i)}\mid\mathbf{x}_{\rm sim}\right)\right]\,p(\bm{\theta}^{(l)})p(\bm{\xi}(\bm{\theta}^{(i)}\,;\,\bm{\alpha}))}{p_{\bm{\phi}}(\bm{\theta}^{(l)}\mid\mathbf{x}_{\rm obs}^{(i)})}\right]$ (10)
-where $\bm{\theta}^{(l)}\sim p_{\bm{\phi}}(\bm{\theta}\mid\mathbf{x}_{\rm obs}^{(i)})$, $\bm{\alpha}$ are the weights of the neural network $\bm{\xi}(\bm{\theta}^{(i)}\,;\,\bm{\alpha})$, and for each $\bm{\theta}^{(l)}$ we approximate $\mathbb{E}_{p_{\bm{\Psi}}(\mathbf{x}_{\rm sim}\mid\bm{\theta}^{(l)})}\!\left[\,p_{\bm{\xi}(\bm{\theta}^{(i)}\,;\,\bm{\alpha})}\!\left(\mathbf{x}_{\rm obs}^{(i)}\mid\mathbf{x}_{\rm sim}\right)\right]$ using a Monte Carlo (MC) estimate.
+$\mathcal{L}(\bm{\phi},\bm{\alpha})_{V}\approx-\sum_{i=1}^{N_{\rm obs}}\log\left[\frac{1}{K}\sum_{l=1}^{K}\frac{\mathbb{E}_{p_{\bm{\Psi}}(\mathbf{x}_{\rm sim}\mid\bm{\theta}^{(l)})}\!\left[\,p_{\bm{\xi}(\bm{\theta}^{(i)}\,;\,\bm{\alpha})}\!\left(\mathbf{x}_{\rm obs}^{(i)}\mid\mathbf{x}_{\rm sim}\right)\right]\,p(\bm{\theta}^{(l)})p(\bm{\xi}(\bm{\theta}^{(i)}\,;\,\bm{\alpha}))}{q_{\bm{\phi}}(\bm{\theta}^{(l)}\mid\mathbf{x}_{\rm obs}^{(i)})}\right]$ (12)
+where $\bm{\theta}^{(l)}\sim q_{\bm{\phi}}(\bm{\theta}\mid\mathbf{x}_{\rm obs}^{(i)})$, $\bm{\alpha}$ are the weights of the neural network $\bm{\xi}(\bm{\theta}^{(i)}\,;\,\bm{\alpha})$, and for each $\bm{\theta}^{(l)}$ we approximate $\mathbb{E}_{p_{\bm{\Psi}}(\mathbf{x}_{\rm sim}\mid\bm{\theta}^{(l)})}\!\left[\,p_{\bm{\xi}(\bm{\theta}^{(i)}\,;\,\bm{\alpha})}\!\left(\mathbf{x}_{\rm obs}^{(i)}\mid\mathbf{x}_{\rm sim}\right)\right]$ using a Monte Carlo (MC) estimate.
 We use the logsumexp function to ensure the MC estimates are stable.
@@
 RVNP-T, the tuned variant of RVNP, includes an extra tuning step that fixes the neural network parameters of the error model $\bm{\alpha}$, and uses the original simulated dataset $D$ to optimise the adjusted NPE objective
-$\mathcal{L}(\bm{\phi})_{\rm{NPE}(\alpha)}:=-\mathbb{E}_{p(\bm{x}_{\rm sim},\bm{\theta})}\mathbb{E}_{p_{\bm{\xi}(\bm{\theta})}(\bm{x}_{\rm obs}|\,\bm{x}_{\rm sim})}[\log p_{\bm{\phi}}(\bm{\theta}|\,\bm{x}_{\rm obs})].$ (11)
-This final objective can be identified with the noisy neural posterior estimation (NNPE, Ward et al.
-2022) objective and the error augmentation method suggested by Cranmer et al.
-(2020).
+$\mathcal{L}(\bm{\phi})_{\rm{NPE}(\alpha)}:=-\mathbb{E}_{p(\bm{x}_{\rm sim},\bm{\theta})}\mathbb{E}_{p_{\bm{\xi}(\bm{\theta})}(\bm{x}_{\rm obs}|\,\bm{x}_{\rm sim})}[\log q_{\bm{\phi}}(\bm{\theta}|\,\bm{x}_{\rm obs})].$ (13)
+This final objective can be identified with the noisy neural posterior estimation (NNPE, ward2022robust) objective and the error augmentation method suggested by Cranmer_2020.
 However, our error model does not have to be globally fixed and has been inferred using variational inference.
 3.4 Error modelling
-In the experiments, we adopt two error models. The default error model that we adopt in RVNP is given by the Gaussian covariance matrix
+Theoretically, any model can be introduced within the RVNP framework.
+However, we opted to only include error models that can inflate the error on the simulator model.
+The default error model that we adopt in RVNP is given by the Gaussian covariance matrix
 $\bm{\xi}(\bm{\theta})=\rm{Diag}(\rm{NN}(\bm{\theta}\,;\bm{\alpha}))+\bm{\Lambda},$ (14)
-a neural network that outputs the diagonal components of the covariance matrix and the non-diagonal components $\bm{\Lambda}$ are globally learnt. We also include a global error model.
+a neural network that outputs the diagonal components of the covariance matrix and the non-diagonal components $\bm{\Lambda}$ are globally inferred.
+The second model we include is a global Gaussian model
 $\bm{\xi}(\bm{\theta})=\bm{\Sigma},$ (15)
@@
 Our method generalises very easily to include any inductive bias that we believe explains the simulation-to-reality gap.
-Our approach can be understood in terms of the simulator defining a population prior $p(\bm{\theta},\bm{x}_{\rm sim})=p_{\bm{\Psi}}(\bm{x}_{\rm sim}|\,\bm{\theta})p(\bm{\theta})$, while the error model can be identified with the likelihood of $\bm{x}_{\rm sim}$ for the observed data.
-In this framework, the simulator model describes our prior understanding of the true DGP.
-In this paper, we allow only for Gaussian error-type models that allow the forward model to inflate the noise on the output of the simulator.
-This allows us to account for the misspecification using a Gaussian forget mechanism.
-Including an error model is an assumption about the true DGP, and the exact corruption may be unknown.
-However, we follow the Box principle of all models are wrong, some are useful (Box, 1976) and attempt to forward model the DGP using the simulator and an error model term that can account for corruption in the data.
-Error modelling in robust SBI is not new; Ward et al.
-(2022) introduced a spike and slab error model, which assumes that a summary statistic is either well-specified or misspecified.
-The parameter adjustment method of Kelly et al.
-(2024) can be thought of an error model $p_{\bm{\xi}}(\bm{x}_{\rm obs}|\,\bm{x}_{\rm sim})=\delta(\bm{x}_{\rm obs}-\bm{x}_{\rm sim}-\bm{\xi})$, integrating with respect to $\bm{x}_{\rm sim}$ first.
-Adjustment parameters will move the observed point to a region of high probability with respect to the simulated samples.
-On a point-by-point basis, there are infinitely many solutions to this problem, and there is no guarantee of where an OOD point should be mapped back to in the original sample without a high volume of data.
-Furthermore, the spike-and-slab error model invokes the assumption that the true corruption model is unknown, and that the misspecification forces us to forget certain summary statistics if they lie OOD.
-This assumption relies on the misspecification occurring along individual axes, which is unrealistic in many situations and is subject to the hyperparameter choices of the spike-and-slab error model.
-In general, misspecification has no guarantee of being represented exactly along the given axes, even if the components represent summary statistics.
-To consider why, let $p_{\gamma}(\bm{x})$ be the probability density function parametrised by $\gamma$ and assume that $T(\bm{x})$ represents a sufficient statistic of $\bm{\theta}$.
-By the Fisher-Neyman factorization theorem (Hogg et al.
-2005, pp 376-377), there exist non-negative functions $k_{\gamma}$ and $h$ such that $p_{\gamma}(\bm{x})=h(\bm{x})k_{\gamma}(\bm{x})$.
-If $g$ is a bijection, then
-$f_{\gamma}(x)=h(x)\,k_{\gamma}(g^{-1}(g(T(x)))),$ (14)
-$f_{\gamma}(x)=h(x)\,(k_{\gamma}\circ g^{-1})(g\circ T(x)),$ (15)
-and, therefore, $g\circ T(x)$ is also a sufficient statistic.
-In this paper, we choose to model a full covariance matrix to account for misspecification that may not occur exactly along each axis.
-While suitable pre-processing may help address this issue, we choose to model the full covariance and infer this.
+Our approach can be understood in terms of the simulator defining a population prior $p(\bm{\theta},\bm{x}_{\rm sim})=p_{\bm{\Psi}}(\bm{x}_{\rm sim}|\,\bm{\theta})p(\bm{\theta})$ and the error model representing the likelihood of $\bm{x}_{\rm sim}$ for the observed data.
 3.5 Prior over the correction model parameters
-In RVNP, a prior over the error model parameters can be introduced to regularise the influence of the error model and make the inference more robust in the single-posterior case.
-Adopting an error model makes the generative model overparameterized, and the inference process will be heavily dependent on the prior assumptions about misspecification in the single-point inference.
-In our experiments, we do not use any prior distribution over the error model parameters.
-This allows us to focus on the robustness imparted by multiple observations.
+In RVNP, a (pseudo-)prior over the error model parameters $p(\bm{\xi}(\bm{\theta}))$ can be introduced to regularise the influence of the error model and make the inference more robust in the single-posterior case.
+In our experiments, we only use a uniform prior distribution over the error model parameters, as we focus on the robustness imparted by multiple observations.
+Algorithm 1 Robust Variational Neural Posterior Estimation (RVNP)
+[…]
+2:repeat
+3:  Sample minibatch $B\subset D$
+4:  $\mathcal{L}_{\rm NLE}(\bm{\Psi})\leftarrow-\mathbb{E}_{B}[\log p_{\bm{\Psi}}(\bm{x}_{\rm sim}\mid\bm{\theta})]$
+5:  $\bm{\Psi}\leftarrow\bm{\Psi}-\eta\nabla_{\bm{\Psi}}\mathcal{L}_{\rm NLE}$
+6:until convergence
+7:return $p_{\bm{\Psi}^{*}}(\bm{x}_{\rm sim}\mid\bm{\theta})$
+8:
+9:LearnPosteriorAndErrorModel($p_{\bm{\Psi}^{*}}$, $\mathcal{O}=\{\bm{x}_{\rm obs}\}$)
+10:repeat
+11:  for each $\bm{x}_{\rm obs}\in$ minibatch $\mathcal{O}$ do
+12:   $\{\bm{\theta}^{(l)}\}_{l=1}^{K}\sim q_{\bm{\phi}}(\bm{\theta}\mid\bm{x}_{\rm obs})$
+13:   for $l=1$ to $K$ do
+14:    $\{\bm{x}_{\rm sim}^{(l,m)}\}_{m=1}^{M}\sim p_{\bm{\Psi}^{*}}(\bm{x}_{\rm sim}\mid\bm{\theta}^{(l)})$
+15:    $w^{(l)}\leftarrow\frac{1}{M}\sum_{m=1}^{M}\frac{p_{\bm{\xi}_{\alpha}(\bm{\theta}^{(l)})}(\bm{x}_{\rm obs}\mid\bm{x}_{\rm sim}^{(l,m)})p(\bm{\theta}^{(l)})p(\bm{\xi}_{\alpha})}{q_{\bm{\phi}}(\bm{\theta}^{(l)}\mid\bm{x}_{\rm obs})}$
+16:   end for
+17:   $\mathcal{L}_{V}\leftarrow-\log\left(\frac{1}{K}\sum_{l=1}^{K}w^{(l)}\right)$
+18:   $(\bm{\phi},\bm{\alpha})\leftarrow(\bm{\phi},\bm{\alpha})-\eta\nabla_{(\bm{\phi},\bm{\alpha})}\mathcal{L}_{V}$
+19:  end for
+20:until convergence
+21:return $q_{\bm{\phi}^{*}}(\bm{\theta}\mid\bm{x}_{\rm obs})$, $p_{\bm{\xi}_{\alpha}^{*}(\bm{\theta})}(\bm{x}_{\rm obs}\mid\bm{x}_{\rm sim})$
+22:
+23:TunePosterior($q_{\bm{\phi}^{*}}$, $p_{\bm{\xi}_{\alpha}^{*}}$, $D$) // Optional: RVNP-T
+24:repeat
+25:  Sample $(\bm{x}_{\rm sim},\bm{\theta})\sim D$, $\bm{x}_{\rm obs}\sim p_{\bm{\xi}_{\alpha}^{*}(\bm{\theta})}(\bm{x}_{\rm obs}\mid\bm{x}_{\rm sim})$
+26:  $\mathcal{L}_{\rm tune}(\bm{\phi})\leftarrow-\log q_{\bm{\phi}}(\bm{\theta}\mid\bm{x}_{\rm obs})$
+27:  $\bm{\phi}\leftarrow\bm{\phi}-\eta\nabla_{\bm{\phi}}\mathcal{L}_{\rm tune}$
+28:until convergence
+29:return $q_{\bm{\phi}^{**}}(\bm{\theta}\mid\bm{x}_{\rm obs})$
+30:
+31:Infer($\bm{x}_{\rm obs}$, $N_{\rm samples}$) // With optional SIR
+32:for $i=1$ to $N_{\rm samples}$ do
+33:  $\bm{\theta}^{(i)}\sim q_{\bm{\phi}^{*}}(\bm{\theta}\mid\bm{x}_{\rm obs})$
+34:end for
+35: Optionally apply sample-importance-resampling to $\{\bm{\theta}^{(i)}\}$
+36:return $\{\bm{\theta}^{(i)}\}_{i=1}^{N_{\rm samples}}$
 4 Experiments and results
-We evaluate our method and test the main claim of our paper: that RVNP and its tuning variant can recover robust amortised posterior inference.
-In the experiments section, we test four variants of our algorithm
+We test the main claim of our paper: that RVNP and its tuned variant can recover robust amortised posterior inference.
+In each task, we chose a fixed simulation budget of $N_{\rm sim}=100,000$ and retained $10\%$ for validation.
+For the three synthetic tasks, we test our methods for $N_{\rm obs}\in\{1,10,100,1000,10000\}$ points.
+For the task using real astronomical data, we test our methods for $N_{\rm obs}\in\{1,10,100,1000\}$ due to limitations on the available data.
+Furthermore, we implemented sample-importance-resampling with $5000$ particles and $100$ samples per particle.
+In the experiments section, we test three variants of our algorithm
   • RVNP: the standard RVNP algorithm with no final tuning step, using the error model given in Equation 14.
-  • RVNP-Global: the standard RVNP algorithm with no final tuning step, using the global error model given in Equation 13.
+  • RVNP-G: the standard RVNP algorithm with no final tuning step, using the global error model given in Equation 15.
   • RVNP-T: the tuned version of RVNP.
-  • RVNP-Global-T: the tuned version of RVNP-Global.
-In each experiment, we only test the tuned algorithm for whichever algorithm performed better, RVNP or RVNP-Global. The architecture and training procedure for each task are described in the Appendix.
+For the RVNP and the RVNP-G algorithm, we evaluate the metrics with and without sample-importance-resampling.
+For both the variational posterior and the simulator proxy, we used neural spline flows (NeuralSplineFlows), and the hyperparameters were kept consistent among the tasks.
+An overview of the architecture and training procedure for each task is described in the appendices.
 Benchmarking Algorithms.
-In these tasks, we benchmark RVNP against noisy neural posterior estimation (NNPE) from Ward et al.
-(2022) due to the similar performance of NNPE with RNPE therein and NNPE’s amortisation ability.
-We do not benchmark against the MMD or consistency loss methods (Elsemüller et al.
-2025; Mishra et al.
-2025) due to their dynamic adaptation of the neural statistics.
-We also benchmark against vanilla NPE.
+In these tasks, we benchmark RVNP against noisy neural posterior estimation (NNPE) from ward2022robust because NNPE is a robust amortised Bayesian inference algorithm.
+Furthermore, NNPE exhibits similar performance relative to their robust neural posterior estimation algorithm, which is not suited to amortised inference.
+We do not benchmark against the unsupervised domain adaptation or consistency loss methods for amortised Bayesian inference (elsemüller2025doesunsuperviseddomainadaptation; mishra2025robust) due to their dynamic adaptation of the embedding space.
+We also benchmark against standard NPE.
 Summary of Results.
-We find that RVNP and RVNP-T can recover robust posterior inference in an amortised manner across a range of different tasks, including cases where a significant number of points have ID data.
-Furthermore, we test RVNP and RVNP-T using low-resolution spectra cross-matched with high-resolution spectroscopic data, and we validate our posterior recovery against classically derived results.
-For each task, except the Spectra task, we test our method for $N_{\rm obs}\in\{1,10,100,1000,10000\}$ points and evaluate the effectiveness of our methods.
-We note that in the absence of misspecification, RVNP collapses towards NPE and is well-specified and calibrated.
-RVNP and RVNP-T can recover robust posterior inference without adopting priors over misspecification or introducing ad hoc hyperparameters that must be tuned.
-We display examples of posterior inference in the $N_{\rm obs}=1000$ case for RVNP against NPE for each of the tasks for a single example in Figure 2.
-We argue that the success of RVNP, the interpretability of the error correction model, and the lack of hyperparameters to be tuned (other than the optimiser hyperparameters) point to a significant contribution to robust amortised SBI.
-In all tasks, our algorithm is overconfident for $N_{\rm obs}=1$.
-This is not surprising, as without any other information, our model collapses to the NPE solution.
-With stronger priors over the misspecification, we could make this step more robust, but this defeats the purpose of allowing the data to drive the error model.
+We found that RVNP and RVNP-T can recover robust posterior inference in an amortised manner across a range of different tasks, including cases where a significant number of points have ID data.
+RVNP-T does not increase the robustness of the posterior across any task, and should be avoided unless the practitioner believes the error model has been inferred globally and wants to infer posteriors on streaming data in a fast manner.
+Furthermore, we found that sample-importance-resampling from the variational posterior increases the posterior performance slightly.
+Our methods performed well against the benchmark algorithms except for the $N_{\rm obs}=1$ case, where we find that our algorithm is always overconfident to a certain degree.
+This is not surprising, as RVNP is highly over-parametrised without enough data to learn the error model reliably.
+With stronger priors on misspecification and reducing the flexibility of the error model, we could make this step more robust.
+However, this defeats the purpose of allowing the data to drive the error model.
+Furthermore, we do not attempt to address robustness in the non-amortised inference problem, but include it for completeness.
+We show examples of posterior inference in the $N_{\rm obs}=1000$ case for RVNP against NPE for each of the tasks for a single example in Figure 2 We argue that the success of RVNP, the interpretability of the error correction model, and the lack of misspecification-dependent hyperparameters to set, amount a significant contribution to robust amortised SBI.
 4.1 Metrics for assessing misspecification
-We consider three main metrics to assess the robustness of the inference.
+We consider four metrics to assess the robustness of the inference.
 Let $\bm{\theta}^{*}$ be the true value of the parameter.
-Assuming that we have a labelled test set $T=\{\bm{\theta}^{*(i)},\bm{x_{\rm obs}}^{(i)}\}_{i=1}^{N_{\rm test}}$, we evaluate the log posterior probability (LPP) of the true parameter over the dataset to assess the performance of RVNP.
-LPP has been extensively used in SBI literature (Papamakarios & Murray 2016; Hermans et al.
-2020; Ward et al.
-2022;Kelly et al.
-2024; Wehenkel et al.
-2025).
-Given a credible level $\gamma$, let $\rm{HDR}_{p(\bm{\theta}|\,\bm{x}_{\rm obs})}(1-\gamma)$ represent the $1-\gamma$ highest posterior density region of the posterior $p(\bm{\theta}|\,\bm{x}_{\rm obs})$.
+Assuming that we have a labelled test set $T=\{\bm{\theta}^{*(i)},\bm{x_{\rm obs}}^{(i)}\}_{i=1}^{N_{\rm test}}$, we evaluate the log posterior probability (LPP) of the true parameter for all points in the dataset.
+LPP has been extensively used in SBI literature to assess performance, and is a useful metric for assessing the ability to predict the truth from observation (2016arXiv160506376P; hermans2020likelihoodfreemcmcamortisedapproximate; ward2022robust;kelly2024misspecificationrobust; wehenkel2025addressing).
+Given a credible level $\gamma$, let $\rm{HDR}_{q_{\bm{\phi}}(\bm{\theta}|\,\bm{x}_{\rm obs})}(1-\gamma)$ represent the $1-\gamma$ highest posterior density region of the posterior estimator $q_{\bm{\phi}}(\bm{\theta}|\,\bm{x}_{\rm obs})$.
 The expected posterior coverage (EPC) at a given confidence level over a test set is given by
-$\rm{EPC}(\gamma):=\mathbb{E}_{\bm{\theta}^{*},\bm{x}_{\rm obs}\sim T}[\mathbf{1}\{\theta^{*}\in\rm{HDR}_{p(\bm{\theta}|\,\bm{x}_{\rm obs})}(1-\gamma)\}],$ (16)
+$\rm{EPC}(\gamma):=\mathbb{E}_{\bm{\theta}^{*},\bm{x}_{\rm obs}\sim T}[\mathbf{1}\{\theta^{*}\in\rm{HDR}_{q_{\bm{\phi}}(\bm{\theta}|\,\bm{x}_{\rm obs})}(1-\gamma)\}],$ (16)
 where $\mathbf{1}$ is the indicator function.
-EPC is a commonly used metric to assess robustness and calibration of posterior distributions, particularly when looking at single observation situations.
-When comparing posterior calibration across a range of amortised datasets, we compute the average expected posterior coverage AEPC
+EPC is a commonly used metric to assess robustness and calibration of posterior distributions (kelly2025simulation), particularly when looking at single observation situations.
+When comparing posterior calibration across a range of amortised datasets, we compute the average expected posterior coverage (AEPC)
 $\alpha:=\int_{0}^{1}[\rm{EPC}(\gamma)-\gamma]d\gamma,$ (17)
-which represents the average calibration across the test set. Finally, we also compute the normalised (root) mean squared error (NMSE):
-$\text{NMSE}=\frac{1}{N_{\rm obs}}\sum_{j=1}^{N_{\rm obs}}\frac{\sqrt{\frac{1}{S}\sum_{s=1}^{N_{\rm samples}}\left(\theta^{*}_{j}-\theta^{(s)}_{j}\right)^{2}}}{(\max(\theta_{\rm{prior}})-\min(\theta_{\rm{prior}}))},$ (18)
-which evaluates the normalised accuracy of the posterior prediction to the truth relative to the prior width.
+which represents the average calibration across the test set. The AEPC metric is evaluated using the density approach of hyndman1996hdr for calculating highest posterior density intervals .
+We also compute the average expected marginal posterior coverage (AEMPC) for each parameter $\bm{\theta}_{i}$, which is common when we only obtain samples of the posterior distribution (wehenkel2025addressing).
+The AMEC can be implemented in a more stable fashion than the posterior density approach but can oversimplify the complex regions of the posterior.
+It is calculated by first estimating
+$\rm{EMPC}(\gamma)_{i}:=\mathbb{E}_{\bm{\theta}_{i}^{*},\bm{x}_{\rm obs}\sim T}[\mathbf{1}\{\theta_{i}^{*}\in\rm{HDR}_{q_{\bm{\phi}}(\bm{\theta}_{i}|\,\bm{x}_{\rm obs})}(1-\gamma)\}].$ (18)
+The AEMPC is defined as
+$\alpha\,\rm{(marginal)}:=\frac{1}{m}\sum_{i=1}^{m}\int_{0}^{1}[\rm{EMPC}(\gamma)_{i}-\gamma]d\gamma,$ (19)
+which is necessary for comparing the coverage of the posterior when using sample-importance-resampling, as we only retrieve the unnormalised posterior associated with sample-importance-resampling distribution.
+We did not wish to rely on multi-dimensional density estimation methods to evaluate the posterior coverage from samples from the joint posterior, as we found that it depended on the density estimation algorithm.
+Finally, we also evaluated the normalised root mean squared error (NRMSE):
+$\text{NRMSE}=\frac{1}{N_{\rm obs}}\sum_{j=1}^{N_{\rm obs}}\frac{\sqrt{\frac{1}{S}\sum_{s=1}^{N_{\rm samples}}\left(\theta^{*}_{j}-\theta^{(s)}_{j}\right)^{2}}}{\rm{Std}(\theta_{\rm{prior}})},$ (20)
+which evaluates the normalised accuracy of the posterior prediction to the truth relative to the standard deviation of the prior training samples in each dimension.
+4.1.1 Description of Plots
+We evaluated the metrics on each task and displayed the results in Figures 3, 4, 5, and 6.
+The purple dashed line corresponds to NNPE, and the orange dashed line corresponds to NPE.
+The blue points are from RVNP with the darker blue indicating sample-importance-resampling.
+The green points are from RVNP-G and the darker green indicates sample-importance-resampling.
+The red points corresponds to the tuned variant RVNP.
+In Appendix A.2, we show how RVNP performs on well-specified tasks.
+The top-left plot in each image shows the AEPC ($\alpha$) as a function of the number of observations, with the light green hue showing the desired target region.
+A value of 0 in this metric indicates well-calibrated.
+Any values less than this implies overconfidence, and positive results imply posterior under-confidence.
+Under-confidence is more desirable than overconfidence.
+The values of NPE and NNPE are independent of the number of observations because they are not tuned dynamically.
+The top-right plot in each image shows the LPP as a function of the number of observations, with the error bars indicating the standard deviation.
+The purple hue indicates the standard deviation of LLP from NNPE, and the orange hue indicates the standard deviation of LLP from NPE.
+A higher value of LPP is better.
+The bottom-left plot in each image shows the AMEPC ($\alpha$ marginal) as a function of the number of observations, with the light green hue showing the desired target region.
+Similarly to AEPC, a value of 0 in this metric indicates well-calibrated.
+In the bottom-right plot, we show the NRMSE as a function of the number of observations, where lower values are better.
 4.2 Task A: CS
-We reproduce the cancer and stromal cell development benchmark task from Ward et al.
-(2022).
+We reproduce the cancer and stromal cell development benchmark task from ward2022robust.
 The simulator models the development of cancer and stromal cells in 2D space based on the locality of a cell relative to unobserved parents.
@@
 The summary statistics are as follows: N Cancer and N Stromal, the number of cancer and stromal cells, respectively; and (Mean Min Dist) and (Max Min Dist), the mean and maximum distance from the stromal cells to their nearest cancer cell, respectively.
-The Numba just-in-time implementation of this task was taken directly from the data products of Ward et al.
-(2022).
+The Numba just-in-time implementation of this task was taken directly from the data products of ward2022robust.
 Misspecification. The misspecification in the observed data is introduced by removing cells in the core regions of tumours, which mimics necrosis.
 Results.
-We display the results for the CS task in Figure 3 and provide an example posterior in the $N_{\rm obs}=1000$ case in Figure 2.
-Finally, we display the simulated observations together with observed points in Figure 7.
-In this task, the misspecification is minimal, and all versions of our algorithm can recover robust posterior inference beginning at the $N_{\rm obs}=10$ point.
-We find that all variants of RVNP perform similarly to NNPE in this task.
-This is not surprising, as the misspecification can be described by inflating the covariance along two specific axes.
-The RVNP algorithms slightly outperform NNPE in NMSE.
+We show the results for the CS task in Figure 3 and provide an example posterior of RVNP with $N_{\rm obs}=1000$ in Figure 2.
+We display the simulated observations together with observed points in Figure 16.
+In this task, the misspecification of removing cells in the core regions that mimics necrosis primarily impacts the number of cancer cells summary statistics.
+In Figure 8, we show the evaluation of RVNP on a well-specified version of the CS task.
+In Figure 12, we show the final two dimensions of the RVNP ($N_{\rm obs}=1000$) covariance matrix eigenvalues evaluated at different points of the training data.
+All versions of our algorithm can recover robust posterior inference after the $N_{\rm obs}=10$ point in the CS task, performing better than NNPE and NPE.
+However, NNPE still performs relatively well, as the misspecification can be described by inflating the covariance along two specific axes.
+Both the global and the neural error covariance matrix represent a constant covariance matrix that points in a given direction that best describes the misspecification, which offers a form of model criticism for this problem.
+Our methods have better posterior coverage than NNPE and NPE, while at the same time achieving high log-probability and samples that are closer to the truth.
+Furthermore, they do not rely on any parameters that control the influence of the misspecification on the posterior.
 Figure 3: Results for the CS task.
@@
 For the log-probability, higher values are better.
-For NMSE, lower values are better.
+For NRMSE, lower values are better.
+(SIR) indicates sample-importance-resampling.
 4.3 Task B: SIR
-We include the misspecified susceptible-infected-recovered (SIR) task from Ward et al.
-(2022), which takes the stochastic model of epidemic spread modelled conditional on $\bm{\theta}=(\beta,\gamma)$, the time-varying infection rate and the recovery rate, respectively.
+We include the misspecified susceptible-infected-recovered (SIR) task from ward2022robust, which takes the stochastic model of epidemic spread modelled conditional on $\bm{\theta}=(\beta,\gamma)$, the time-varying infection rate and the recovery rate, respectively.
 The SIR model emulates ideal disease transmission dynamics from the susceptible ($s$), infected ($i$), and recovered ($r$) parameters as
 $\frac{ds}{dt}=-\beta si,\quad\frac{di}{dt}=\beta si-\gamma i,\quad\frac{dr}{dt}=\gamma i.$ (21)
-Ward et al.
-(2022) employs a stochastic extension by using time-dependent transmission dynamics through a variable infection rate $\tilde{\beta}_{t}$, accounting for external factors such as policy interventions or pathogen mutations.
-This stochastic process is characterized using the basic reproduction number $R_{0t}=\frac{\tilde{\beta}_{t}}{\gamma}$, which follows the mean-reverting stochastic differential equation:
+ward2022robust employs a stochastic extension by using time-dependent transmission dynamics through a variable infection rate $\tilde{\beta}_{t}$, accounting for external factors such as policy interventions or pathogen mutations.
+This stochastic process is characterised using the basic reproduction number $R_{0t}=\frac{\tilde{\beta}_{t}}{\gamma}$, which follows the mean-reverting stochastic differential equation:
 $dR_{0t}=\eta\left(\frac{\beta}{\gamma}-R_{0t}\right)dt+\sigma R_{0}dW_{t},$ (22)
@@
 $\eta=0.05$ and $\sigma=0.05$ are fixed and the goal is to infer the parameters $\beta$ and $\gamma$.
-The Julia code to sample from this process was taken directly from the data products of Ward et al.
-(2022).
+The Julia code to sample from this process was taken directly from the data products of ward2022robust.
 The summary statistics produced in this task are the mean, median and maximum number of infections, the day of the maximum number of infections, and the day at which half of the total number of infections was reached, and the mean autocorrelation of infections with lag 1.
@@
 For the log-probability, higher values are better.
-For NMSE, lower values are better.
+For NRMSE, lower values are better.
+(SIR) indicates sample-importance-resampling, not to be confused with the name of the task.
 Misspecification To introduce misspecification in the observations, a small reporting delay is adopted where weekend infection counts are reduced by $5\%$ and are added to the Monday count.
 Results.
-We display the results for the SIR task in Figure 4 and provide an example posterior in the $N_{\rm obs}=1000$ case in Figure 2.
-Finally, we display the simulated observations together with observed points in Figure 8.
-In this task, the misspecification is extreme but occurs only along the final axis, and all versions of our algorithm can recover robust posterior inference beginning at the $N_{\rm obs}=10$ point.
-As the misspecification occurs along a given axis, we expect NNPE to perform well in this task.
-We find that each of our algorithms witnesses slight performance gains in each of the metrics as the number of observations increases.
+This misspecification in this task only has a significant impact on the autocorrelation summary statistic.
+We show the results for the SIR task in Figure 4 and provide an example of the posterior distribution ($N_{\rm obs}=1000$) in Figure 2.
+We display the simulated observations together with the observed points in Figure 17.
+In Figure 9, we show the evaluation of RVNP on a well-specified version of the SIR task.
+In Figure 13, we plot the final two dimensions of the RVNP ($N_{\rm obs}=1000$) covariance matrix eigenvalues evaluated at different points of the training data.
+In this task, the misspecification appears extremely OOD along the axis of the autocorrelation statistic.
+All versions of our algorithm can recover robust posterior inference after the $N_{\rm obs}=10$ point, and they perform better than NNPE and NPE in this task.
+However, NNPE still performs well, as the misspecification can be described by inflating the covariance along the final axis.
+NPE performs very poorly in this task, due to the extreme misspecification along the final axis.
+Our methods have better posterior coverage than NNPE and NPE, while at the same time achieving high log-probability, and samples that are closer to the truth.
+Furthermore, the global and neural error covariance matrix learn that the covariance inflates along the final axis.
+However, the neural error model outperforms the global error model, as the degree of misspecification along that axis depends on $\bm{\theta}$.
+The effect of this is that both the log-probability and the NRMSE are better for RVNP than the global version, as the global version scales up the covariance to account for this misspecification everywhere.
 4.4 Task C: Pendulum Task
@@
 $\phi$ is a stochastic phase shift.
-This task was inspired by the task from Wehenkel et al.
-(2025).
+This task was inspired by the task from wehenkel2025addressing.
 However, it differs significantly and was adjusted to test our claim that increasing the number of posterior observations will better constrain the parameters.
 Misspecification We synthesise a time calibration error in the instrumentation that causes the instrument to take 200 measurements every 0.075 seconds instead of the simulated 0.05 seconds.
-There is a significant probability that the misspecified point will appear ID due to the mild misspecification.
+There is a significant probability that the misspecified point will appear ID due to the form of the misspecification.
 Neural Statistic Estimation In this example, each of the data points is a single observation.
 The neural statistic estimator is an embedding of the full pendulum time series into a lower-dimensional representation of the data.
-Following Chen et al.
-(2021), we use the Shannon-Jensen InfoMax objective (Hjelm et al., 2019) to target sufficient neural statistics, $\iota_{\bm{\omega}}$, to encode the high-dimensional data.
+Following chen2021neuralapproximatesufficientstatistics, we use the Shannon-Jensen InfoMax objective (hjelm2018learning) to target sufficient neural statistics, $\iota_{\bm{\omega}}$, to encode the high-dimensional data.
 This objective function maximises the mutual information between $\bm{x}_{\rm sim}$ and $\bm{\theta}$ using a discriminator network.
+We give an overview of the chosen architecture in the appendices.
 Figure 5: Results for the Pendulum task.
@@
 For the log-probability, higher values are better.
-For NMSE, lower values are better.
+For NRMSE, lower values are better.
+(SIR) indicates sample-importance-resampling
 Results.
-We display the results for the Pendulum task in Figure 5 and provide an example posterior in the $N_{\rm obs}=1000$ case in Figure 2.
-Finally, we display the simulated observations together with observed points in Figure 9.
-In this task, the misspecification is geometrically significantly more complex.
-Moreover, due to the nature of the misspecification, most of the points will appear ID.
-Due to the complexity of the misspecification, NNPE struggles significantly.
-However, RVNP and RVNP-Global recover robust posterior inference for this task.
-In particular, RVNP is very well calibrated after the $N_{\rm obs}=100$ point and recovers robust posterior inference across a broad range in parameter space.
+We show the results for the Pendulum task in Figure 5.
+In Figure 2, we display an example of the posterior distribution in the $N_{\rm obs}=1000$ case.
+We show the simulated observations together with the observed points in Figure 18.
+In Figure 10, we evaluate RVNP on a well-specified version of the Pendulum task.
+In Figure 14, we plot the final two dimensions of the RVNP ($N_{\rm obs}=1000$) covariance matrix eigenvalues evaluated at different points of the training data.
+In this task, the misspecification appears more complex geometrically.
+Moreover, due to the nature of the misspecification, most of the points will appear ID relative to the simulated training points.
+All versions of our algorithm can recover more robust posterior inference beginning after the $N_{\rm obs}=10$ point than the NNPE and NPE.
+NNPE struggles significantly, most likely due to the spike-and-slab error model being unable to account for such a complex misspecification.
+However, RVNP and RVNP-G recover robust posterior inference for this task.
+In particular, RVNP is well calibrated after the $N_{\rm obs}=100$ point and recovers robust posterior inference across a broad range in parameter space.
+Furthermore, RVNP is reasonably well calibrated.
 Due to the complexity of the error model, the tuned variant, RVNP-T, does not see the same degree of performance increase.
 However, it does perform better than both NPE and NNPE.
+NPE performs very poorly in the pendulum task.
+Both covariance models have their largest eigenvector pointing in the direction of changing fundamental frequency, as expected.
 4.5 Task D: Spectra Task - Real Gaia BP/RP Data
-The third data release of the European Space Agency’s Gaia telescope (Gaia Collaboration et al., 2016) contain over 220 million flux-calibrated, low-resolution, optical stellar spectra.
+The third data release of the European Space Agency’s Gaia telescope (GAIA_MISSION) contain over 220 million flux-calibrated, low-resolution, optical stellar spectra.
 These spectra are measured by two instruments, the “Blue Photometer” (BP, 330-680 nm coverage in wavelength) and the “Red Photometer” (RP, 640-1050 nm).
-The processed and calibrated (De Angeli, F.
-et al.
-2023; Montegriffo, P.
-et al.
-2023) BP/RP (XP) spectra from Gaia DR3 are low-resolution, contaminated spectra which have multiple difficult systematics to overcome (Huang et al., 2024).
-However, the XP spectra are expected to contain significant information about different stellar parameters (Witten et al., 2022) and robust posterior inference using stellar evolution simulators conditional on the Gaia XP spectra would provide an efficient method for understanding the Milky Way.
-Simulator We use the MIST Choi et al.
-(2016) stellar evolution models to generate stellar parameters compatible with high Galactic latitudes and map each of the effective temperature, log-surface gravity, and metallicity to a medium-resolution synthetic Castelli & Kurucz (2004) model spectral energy distribution.
-This defines the simulator relating $\bm{\theta}=(T_{\rm eff},\log g,\rm{[Fe/H]})$ to $\bm{x}_{\rm sim}$.
-We cut the spectra in their native resolution between 330-1050 nm to define a $301$ dimensional vector which overlaps with the Gaia XP spectra wavelength range but at a significantly different resolution.
-Embedding We pre-train an NSE estimator using the same method as described in the pendulum synthetic task.
+The processed and calibrated (bprpspec; external_bprp) BP/RP (XP) spectra from Gaia DR3 are low-resolution, contaminated spectra which have multiple difficult systematics to overcome (Huang_2024).
+However, the XP spectra are expected to contain significant information about different stellar parameters (2022MNRAS.516.3254W) and robust posterior inference using stellar evolution simulators conditional on the Gaia XP spectra would provide an efficient method for understanding the Milky Way.
+Simulator We use the MIST MIST stellar evolution models to generate stellar parameters compatible with high Galactic latitudes (a region where observable contaminants are low) and map each of the effective temperature, log-surface gravity, and metallicity to a medium-resolution synthetic 2003IAUS.210P.A20C model spectrum.
+This defines a stochastic simulator mapping $\bm{\theta}=(T_{\rm eff},\log g,\rm{[Fe/H]})$ to $\bm{x}_{\rm sim}$.
+We restrict the spectra in their native resolution between 330-1050 nm to define a $301$ dimensional vector which overlaps with the Gaia XP spectra wavelength, but at a significantly different resolution.
+Embedding We pre-train an NSE estimator using the same method as described in the pendulum task.
 Misspecification: We can view the problem of inferring stellar parameters using real Gaia XP spectra and synthetic stellar evolution models as a misspecification problem.
 The Gaia XP spectra are expected to have two significant differences from the synthetic simulation model.
-Firstly, the XP spectra are lower resolution than the synthetic spectra and, therefore, contain less information than high-resolution spectra from the simulator.
-Secondly, there are inherent systematic errors in the processing of the XP spectra that cause a simulation gap even if we know the spectroscopic parameters using traditional high-resolution methods.
-For the real dataset, we target high Galactic latitudes due to the minimal impact of photometric extinction in these regions and a relatively homogenous population of main-sequence stars (O’Callaghan et al., 2024).
-We select all stars with absolute Galactic latitude $|b|>80^{\circ}$ that have valid LAMOST (Wang, 2022) spectroscopically determined stellar parameters.
+Firstly, the XP spectra are lower resolution than the synthetic spectra and, therefore, contain less information.
+Secondly, there are inherent systematic errors in the processing of the XP spectra and in the stellar evolution modelling that causes a simulation gap even if we knew the spectroscopic parameters exactly.
+For the real dataset, we target high Galactic latitudes due to the minimal impact of dust extinction and other contaminants (high_gal_me).
+We select all stars with absolute Galactic latitude $|b|>80^{\circ}$ that have valid LAMOST (LAMOSTDR8) spectroscopically determined stellar parameters.
 These spectroscopically determined stellar parameters will act as ground truth for our experiment, but we should note that they have their own errors arising from the spectroscopic determination.
@@
 This leaves us with a dataset of size $N_{\rm obs}=1053$.
-Misspecification Summary.
-For those not familiar with stellar astronomy, the observed dataset is generated from real data with ground truth values that were selected so that measurement error and external factors have a minimal impact on the data, helping us better isolate the model misspecification.
 Results.
-We display the results for the Spectra task in Figure 5 and provide an example posterior in the $N_{\rm obs}=1000$ case in Figure 2.
-Finally, we display the simulated observations together with observed points in Figure 10.
-In this task, the misspecification is geometrically very complex due to both the image of the neural statistic and the complexity of stellar evolution.
+We show the results for the Spectra task in Figure 5 and provide an example posterior in the $N_{\rm obs}=1000$ case in Figure 2.
+We display the simulated observations together with the observed points in Figure 19.
+In Figure 11, we evaluate RVNP on a well-specified version of the Spectra task.
+In Figure 15, we plot the final two dimensions of the RVNP ($N_{\rm obs}=1000$) covariance matrix eigenvalues evaluated at different points of the training data.
+In the spectra task, the misspecification is geometrically very complex.
+This arises from using a neural statistic embedding and from the complexity of stellar evolution models.
 Many of the points appear ID relative to the simulated points in this task.
-We naively applied the neural statistic, so that the neural statistic knows nothing about the structure of the Gaia XP spectra.
-We find that the neural statistic struggles to identify the metallicity parameter, most likely due to the neural statistic fitting for high-resolution features in the synthetic spectra.
-Due to the complexity of the misspecification, NNPE struggles significantly.
-However, RVNP and RVNP-Global recover robust posterior inference for this task.
-In particular, RVNP-Global is very well calibrated after the $N_{\rm obs}=10$ point and recovers robust posterior inference across a broad range in parameter space.
-In this task, we displayed the tuned variant of RVNP-T and found that it struggled significantly to recover robust posterior inference.
-It is worth noting that we chose a poor prior for this problem, and we hypothesise that this is why RVNP-Global performs better than RVNP-Global, as the neural network covariance matrix accesses many unseen parts of parameter space.
+We naively applied the neural statistic, so that the neural statistic knows nothing about the structure of the Gaia XP spectra, including instrument response and measurement error.
+The RVNP and RVNP global algorithms can recover robust posterior inference beginning after the $N_{\rm obs}=10$ point.
+We find that both of these algorithms perform better than NNPE and NPE.
+However, RVNP-T collapses to the NPE posterior after tuning.
+NNPE struggles significantly, most likely due to the spike-and-slab error model being unable to account for such a complex misspecification.
+RVNP-G is very well calibrated after the $N_{\rm obs}=100$ point and recovers robust posterior inference across a broad range of parameter space.
+Furthermore, RVNP is reasonably well calibrated, particularly when considering the marginal calibration metric.
+NPE performs very poorly in the pendulum task in terms of calibration and log-probability, but its sampled points are close to the truth, reflecting an extreme overconfidence.
+Investigating the covariance matrix of the error models helps us to understand why the global variant performs better.
+In certain points of parameter space where the covariance matrix has not observed data, the posterior collapses to the NPE estimate, and near these points the inferene is less reliable.
+The covariance matrix provides useful model criticism by considering the error as a function of $\bm{\theta}$.
+We find that we cannot recover the metallicity estimate reliably.
+This is not surprising because the resolution of the synthetic spectra is higher than that of the observed data, and the metallicity will appear as strong local features in the spectra.
+These features are not present in the Gaia BP/RP spectra.
+Domain adaptation approaches that can account for these difference may be desirable in this context.
 Figure 6: Results for the Spectra task.
@@
 For the log-probability, higher values are better.
-For NMSE, lower values are better.
+For NRMSE, lower values are better.
+(SIR) indicates sample-importance-resampling.
 5 Related work
-Misspecification in SBI Model misspecification is understood in likelihood-based methods (Davison 2003, pages 147–148); however, a systematic theory of likelihood-free methods is lacking.
-Misspecification in SBI has been studied in the context of ABC (Frazier et al.
-2020; Bharti et al.
-2022; Fujisawa et al.
-2021), BSL (Frazier et al., 2024), generalized Bayesian inference (Dellaporta et al., 2022), and neural conditional density estimation (Ward et al.
-2022; Kelly et al.
-2024; Huang et al.
-2023; Elsemüller et al.
-2025; Mishra et al.
-2025; Schmitt et al.
-2024; Wehenkel et al.
-2025).
-In this paper, we focused on neural-based methods, where empirically it has been shown that SBI struggles under model misspecification (Cannon et al.
-2022; Schmitt et al.
-2024).
+Misspecification in SBI Model misspecification is understood in likelihood-based methods (davison2003statistical, pages 147–148); however, a systematic theory of likelihood-free methods is lacking.
+Misspecification in SBI has been studied in the context of ABC (10.1111/rssb.12356; bharti2022approximate; fujisawa2021gamma), BSL (Frazier26092024), generalized Bayesian inference (dellaporta2022robust), and neural conditional density estimation (ward2022robust; kelly2024misspecificationrobust; huang2023learning; elsemüller2025doesunsuperviseddomainadaptation; mishra2025robust; schmitt2024detecting; wehenkel2025addressing).
+In this paper, we focused on neural-based methods, where empirically it has been shown that SBI struggles under model misspecification (cannon2022investigating; schmitt2024detecting).
 Robust posterior recovery under model misspecification is essential for the success of SBI, and different methods have emerged to mitigate against it.
-Kelly et al.
-(2025) identifies three main strategies currently used to account for model misspecification in SBI: robust summary statistics, generalised Bayesian inference, and error modelling and adjustment parameters.
-Most of the early solutions in robust SBI were intended for a single data set (Ward et al.
-2022; Kelly et al.
-2024; Huang et al.
-2023).
-Robust amortised SBI Recently, robust amortised neural SBI has been addressed using optimal transport for domain shifts when a calibration set exists (Wehenkel et al., 2025), using unsupervised domain adaptation (Elsemüller et al., 2025), and consistency losses regularisation (Mishra et al., 2025).
-Moreover, Glöckler et al.
-(2023) proposes regularisation techniques to increase the robustness of the learnt posterior against adversarial attacks.
+kelly2025simulation identifies three main strategies currently used to account for model misspecification in SBI: robust summary statistics, generalised Bayesian inference, and error modelling and adjustment parameters.
+Most of the early solutions in robust SBI were intended for a single data set (ward2022robust; kelly2024misspecificationrobust; huang2023learning).
+Robust amortised SBI Recently, robust amortised neural SBI has been addressed using optimal transport for domain shifts when a calibration set exists (wehenkel2025addressing), using unsupervised domain adaptation (elsemüller2025doesunsuperviseddomainadaptation), and consistency losses regularisation (mishra2025robust).
+Moreover, 2023arXiv230514984G proposes regularisation techniques to increase the robustness of the learnt posterior against adversarial attacks.
 Of these methods, only the consistency loss method targets both the likelihood and the posterior.
-The noisy neural posterior estimation from Ward et al.
-(2022) can also be viewed as an amortised SBI method, where a pre-defined error model is used during training to corrupt the simulations.
+The noisy neural posterior estimation from ward2022robust can also be viewed as an amortised SBI method, where a pre-defined error model is used during training to corrupt the simulations.
 Variational methods in SBI Variational methods have been used in multiple capacities to date.
-Wiqvist et al.
-(2021) introduced Sequential Neural Posterior and Likelihood Approximation, which proposes using variational inference (VI) to speed up the inference of likelihood-based methods, similar to the likelihood-based Bayesian approach to VI.
-Glöckler et al.
-(2022) introduces a framework that uses VI for simulation-based inference by using a pre-trained likelihood (or likelihood-ratio) and learn the posterior using VI, then refining the posterior using sampling importance resampling (Rubin, 1987).
-Nautiyal et al.
-(2024) introduces a generative modelling approach based on the variational inference framework and learns an encoder-decoder model in terms of latent variables.
+wiqvist2021sequential introduced Sequential Neural Posterior and Likelihood Approximation, which proposes using variational inference (VI) to speed up the inference of likelihood-based methods, similar to the likelihood-based Bayesian approach to VI.
+glockler2022variational introduces a framework that uses VI for simulation-based inference by using a pre-trained likelihood (or likelihood-ratio) and learn the posterior using VI, then refining the posterior using sampling importance resampling (Rubin1987SIR).
+nautiyal2024variational introduces a generative modelling approach based on the variational inference framework and learns an encoder-decoder model in terms of latent variables.
 They introduce a latent variable that can account for complex structures and dependencies in the simulator model.
-Simons et al.
-(2022) propose a simulation-based inference algorithm that iteratively updates particles to more match the posterior in a variational likelihood-free gradient descent manner.
-6 Discussion
+simons2022variational propose a simulation-based inference algorithm that iteratively updates particles to more match the posterior in a variational likelihood-free gradient descent manner.
+Moreover, dax2023neural use a trained NPE for importance-sampling in a likelihood-based manner.
+6 Discussion and limitations
 In this paper, we introduced RVNP(-T), a robust amortised Bayesian inference method for simulation-based inference that jointly infers the simulation-to-reality gap and the amortised posterior using an importance-weighted autoencoder framework.
-This is the first case of using variational inference for robust simulation-based inference.
-Moreover, it is the first approach that does not rely on tuning hyperparameters of the loss function to recover robust posterior inference.
-We argue that this is an important step toward reliable posterior inference in amortised SBI.
-Previous work’s reliance on tuning parameters or misspecification priors to control the robustness of the posterior implies that there is an unknown parameter that controls the posterior inference in a way that may be nontrivial for real inference tasks.
-Furthermore, the error model is interpretable, overcoming the issues with unsupervised domain adaptation.
-6.1 Neural statistic estimators
-In this paper, we adopt a neural statistic estimator (NSE) before training the neural likelihood surrogate.
-There is an architectural limitation from the necessity of learning the likelihood proxy in RVNP because learning the NSE simultaneously with the likelihood is not a well-posed problem (Brehmer & Cranmer, 2020).
-However, we argue that in amortised SBI, this is not a significant problem.
-Usually, amortised SBI requires a fixed simulation budget up front, and in neural posterior estimation (NPE), the neural conditional density estimator is trained on those simulations.
-When training an NSE for an NPE task, usually the NSE is trained simultaneously with the posterior proxy.
-However, this objective function has no theoretical guarantees of the sufficiency of the statistic (Chen et al., 2021).
-In this paper, we chose to pre-train the NSE using the InfoMax objective (Hjelm et al., 2019) because of its theoretical guarantees of sufficiency.
-We also considered hand-crafted statistics in this paper, as the robustness guarantees occur during the final variational inference step.
-This allows RVNP to include problems where an understanding of the error model is a significant part of the inference task.
-Such as in situations where the chosen summary statistics are in units known to the scientist, and they wish to fit for discrepancies between the model and the data.
-6.2 Connection to domain adaptation
-Unsupervised domain adaptation couples posterior learning to adapting the neural statistics.
-In this paper, we saw that for some tasks, learning a simple covariance matrix parameterised by a neural network on a subset of data does not necessarily generalise well to other areas of parameter space, even with strong inductive bias.
-Complex domain adaptation in a nonlinear fashion will give rise to infinite solutions that can account for the simulation-to-reality gap, and even in simple inductive bias error models throughout this paper, it highlights that the domain adaptation may cause serious issues if many of the points lie ID.
-A separate body of work should discuss error modelling using calibration to compare with the results of Wehenkel et al.
-(2025).
+RVNP is the first example of using variational inference for robust simulation-based inference, and was inspired by robust SBI approaches for synthetic likelihood (frazier2021robust), neural posterior estimation (ward2022robust), and sequential neural likelihood estimation (kelly2024misspecificationrobust).
+We argue that RVNP’s novelty and its ability to recover robust amortised posterior inference without setting misspecification-related tuning parameters constitutes an important step toward reliable posterior inference in amortised SBI.
+Previous work’s reliance on setting a hyperparameter or misspecification prior to control the influence of the misspecification on the posterior inference is undesirable, and as more data are observed we will have a better understanding of the simulation-to-reality gap, which should inform the inference without strong assumptions on the degree of misspecification.
+Moreover, RVNP is reliable in the well-specified case.
+Our work is not without limitations, and in what follows we discuss the scope of RVNP, how it relates to previous error modelling approaches, the use of neural statistic embeddings, and the implications of using variational inference to infer the posterior.
+6.1 Error Modelling
+Including an error model is an assumption about how to model the true data-generative process (DGP) under the assumption that observed data appears corrupted relative to the model simulations.
+In most real-world problems, the exact corruption is unknown and there are infinitely many distributions and assumptions that can bridge the simulation-to-reality gap.
+In this paper, we imposed a strong inductive bias on the distribution that bridges it, so that the error model accounts for the corruption by inflating the error on the synthetic DGP.
+This choice is a neural adaptation of the covariance inflation from frazier2021robust.
+The simulation-to-reality gap should inform our error model as we observed more data from the true DGP, particularly if the misspecification is a global phenomenon (Figure 1).
+By jointly inferring the posterior and the error model, we can allow the error model parameters to be inferred directly from the data without having tunable parameters that dictate the strength of the misspecification (for example, a shrinkage prior over the adjustment parameters (kelly2024misspecificationrobust) or the misspecification regularisation in unsupervised domain adaptation (huang2023learning)).
+Similarly to the spike-and-slab error model introduced by ward2022robust, the error models introduced in this paper allows us to have a form of model criticism.
+We argue that without prior knowledge about the misspecification, inflating the error of the output of the simulator is the most sensible approach, as it is interpretable and allows for a level of trust in the inference procedure.
+The spike-and-slab model invokes a similar assumption: that the true corruption model is unknown and the misspecification encourages the posterior to forget certain summary statistics if they lie OOD.
+This relies on the misspecification appearing along individual axes in summary-statistic space, which is unrealistic in many situations because misspecification has no guarantee of appearing along given axes.
+Furthermore, the success of the spike-and-slab error model is subject to hyperparameter choices that dictate the strength of misspecification.
+We found that in the pendulum and the spectra task NNPE performed poorly for these reasons.
+The parameter adjustment method of kelly2024misspecificationrobust can be thought of as an error model $p_{\bm{\xi}}(\bm{x}_{\rm obs}|\,\bm{x}_{\rm sim})=\delta(\bm{x}_{\rm obs}-\bm{x}_{\rm sim}-\bm{\xi})$ after integrating with respect to $\bm{x}_{\rm sim}$, where $\bm{\xi}$ are the linear adjustment parameters.
+This approach will move the observed point in summary-statistic space to a region of high probability with respect to the simulated samples.
+On a point-by-point basis, there are infinitely many solutions to this problem, and there is no guarantee of where an OOD point should be mapped back to in the original sample without extra information on the misspecification, such as a calibration set.
+Moreover, kelly2024misspecificationrobust state that the ability to recover robust posterior inference in their benchmark tasks depends on the hyperparameters that control the strength of the misspecification, which is difficult to set in practice.
+The flexibility of the RVNP error model also gives rise to certain limitations.
+In RVNP, we are required to choose an error model architecture together with network hyperparameters.
+Although the neural-network parametrised error model performs more favourably than the global covariance in most of the tasks, there are two issues which can arise from its implementation.
+Firstly, we fixed the architecture capacity, implying that the neural-network covariance matrix is massively over-parametrised in the $N_{\rm obs}=1$ and $10$ cases.
+Secondly, as highlighted in the spectra task, the highly nonlinear bridging of the simulation-to-reality through covariance inflation gap may not be suitable if many of the points lie within distribution in a manner that cannot be distinguished by $\bm{\theta}$.
+In the spectra task, significant subsets of $\Theta$ are mapped to observed points in a manner that the covariance matrix collapsed to a delta on the subset.
+We believe that a reliable calibration set that covers such subsets can address these collapses.
+The impact of a complex bridging of the simulation-to-reality gap has also been witnessed in unsupervised domain adaptation (elsemüller2025doesunsuperviseddomainadaptation).
+Introducing an error model inherently adds noise into the forward model.
+We found that the lowest the covariance would decrease to in the well-specified tasks amounted to a standard deviation of $0.01$ on the Gaussian uncertainty in each direction.
+While small, if the simulator is very sensitive to changes in the underlying parameters this could be consequential.
+2021arXiv211006581H observe that SBI methods tend to produce overconfident posterior approximations in general, and perhaps introducing an error model in this situation is less problematic.
+6.2 Choice of summary statistics
+In this paper, we used summary statistics that are hand-crafted and those that are inferred using a neural statistic estimator (NSE).
+Interpretable, hand-crafted summary statistics, or preparing the data into a manageable form with physically motivated units is often highly desirable (kelly2025simulation).
+In the context of misspecification, hand-crafted summary statistics allow RVNP to be used when inferring the misspecification for the purpose of model criticism, providing an understanding of the discrepancies between the simulator model and the data to the practitioner.
+Often, scientific discoveries are made in physical units by fitting for the discrepancy between the model and the observed data.
+For example, in astronomy dust extinction is inferred by comparing the global misspecification between the simulator and the observed spectral data.
+In RVNP, we return both the posterior and the forward model, allowing us to inspect and criticise the model, but also recover fast and robust posterior inference that can readily be applied to large datasets.
+Adopting a neural statistic estimator (NSE) to embed the data into a low-dimensional representation can be highly valuable in SBI if the data is complex and high-dimensional (deistler2025simulationbasedinferencepracticalguide).
+As RVNP is an extension of neural likelihood estimation (NLE) we cannot jointly learn the NSE and the likelihood proxy in an end-to-end fashion (brehmer2020flows).
+A downside of this is that we have to specify separate hyperparameters for the optimiser when inferring the NSE.
+However, staged training allows us to choose objectives that guarantee global sufficiency (chen2021neuralapproximatesufficientstatistics), which may be highly desirable for amortised Bayesian inference.
+As RVNP pre-trains the NSE, we can use the InfoMax objective (hjelm2018learning) and benefit from its theoretical guarantees of global sufficiency.
+A staged training allows us to decouple the inference from learning the NSE, and can help the Bayesian workflow if the main computational difficulty is in dealing with the high-dimensional data once simulated.
+This is true particularly for amortised learning where global sufficient statistics are needed and an up-front simulation budget is defined.
+Non-amortised methods train in rounds, often making end-to-end methods more appealing.
+When learning the NSE and the posterior in an end-to-end fashion, a form of domain adaptation is necessary to account for misspecification.
+While this will help against the observed data having a very poor representation under the NSE, there is an inherit trade-off between learning the posterior and accounting for the misspecification in a manner that depends on a user-defined hyperparameter.
+Although RVNP requires more tunable hyperparameters as a result of the variational inference (and the InfoMax objective, if adopted), it requires no parameters that dictate the strength of the misspecification.
+Similarly to error modelling, complex domain adaptation in a nonlinear fashion may have infinitely many solutions that can account for the simulation-to-reality gap in a manner that is highly uninterpretable, and offers very little model criticism.
+A separate body of work should discuss error modelling using calibration set, and compare with the results of wehenkel2025addressing.
 6.3 Prior misspecification
-A type of misspecification not explicitly addressed in this paper is if the prior $p(\bm{\theta})$ is itself misspecified.
-RVNP should have the same limitations as likelihood-based prior misspecification, except with two crucial differences.
-Any prior chosen must be defined within the support of the prior used to train the likelihood.
-Furthermore, changing the prior distribution may make the likelihood proxy learn different parts of the parameter space suboptimally.
-In the Spectra task, we defined an incorrect prior relative to the observed data, which occurred unintentionally due to the selection effects of the real data.
-It is worth considering the pendulum task for this problem.
-The OOD points appear OOD under this misspecification because the effect of the misspecification makes the points appear they have a higher fundamental frequency.
-However, the information about the misspecification is only available to us because we assume that the prior is well-specified.
-We would lose the robustness of the inference if we assumed a wider prior on the fundamental frequency, causing the synthetic DGP to cover the observed data points.
-6.4 Higher dimensional problems
-Our method inherits the limitations of both neural likelihood estimation and amortised variational inference.
-It is highly desirable to use NSE in higher-dimensional problems to reduce the dimension of the data to help the computational capabilities of the normalising flow on the likelihood estimation.
-In general, RVNP should scale well to higher-dimensional problems, but may struggle to compete with domain adaptation methods for more complex, multimodal data.
-Future work will look at the limitations of using pre-training the NSE and learning the likelihood proxy on the embedded summary statistics.
-Broader Impact Statement
+Prior misspecification was not explicitly addressed within the RVNP tasks.
+Assuming that the neural likelihood proxy is learnt exactly, RVNP should have the same limitations as likelihood-based prior misspecification if the support of the adjusted prior is a subset of the support of the original prior used to train the likelihood.
+In reality, the ability to train the likelihood is influenced by the prior over $\bm{\theta}$, and it may struggle to learn the distribution in regions where there are a low density of points.
+In the experiments, we chose relatively uninformed priors.
+However, we stress that the pendulum task benefits from the uniform bounds in $\bm{\theta}$ defined to generate the simulator samples.
+In the pendulum task, the points appear OOD under the misspecification because its effect makes the points appear as if they have a higher fundamental frequency relative to the summary statistics.
+However, the information about the misspecification is only available to us because we assume that the prior is well-specified and that there is an upper bound to the fundamental frequency.
+With expert knowledge this is not an unreasonable assumption.
+However, we would lose the robustness of the inference if we assumed a wider prior on the fundamental frequency, causing the synthetic DGP to cover the observed data points.
+Therefore, there is a trade-off between extending the prior to higher fundamental frequencies and robustness.
+We note that in the Spectra task we defined a wide, poorly-adopted prior relative to the observed data, which occurred unintentionally due to the selection effects of the real data.
+However, this did not cause any issues.
+There are, however, hard boundaries along a certain direction due to the sharp cut off expected from observed metallicities.
+6.4 Variational Inference and sample-importance-resampling
+RVNP builds on variational methods for solving the inverse problem under a learnt likelihood in SBI (glockler2022variational), and relies on a conditional neural density estimator to approximate the likelihood from a fixed budget of simulations from the simulator model.
+In this sense, RVNP extends NLE by accounting for misspecification using an error model.
+We invoke variational inference to overcome the computational expense of using Hamiltonian Monte Carlo 2011hmcm.book..113N or other Markov Chain methods to jointly infer the parameters of interest and the parameters of the error model.
+The sharing of the parameters over a large model makes the inference highly dependent on the gradients, and struggles to converge even for a diagonal Gaussian covariance error model.
+Furthermore, variational methods allow us to use more expressive models for the error model distribution.
+NLE methods require more hyperparamter tuning and model choice by adopting the variational inference posterior, which also comes at an increased computational expense (deistler2025simulationbasedinferencepracticalguide).
+However, if the amortised likelihood has been trained well, variational methods with sample-importance-resampling can compete with the accuracy of NPE (glockler2022variational).
+Importantly for the misspecification problem, while we increase the number of tunable parameters and architecture choices for the inference step, we do not adopt hyperparameters that control the strength of the misspecification.
+In this situation, the difficult arises from the inference step as opposed to accounting for the misspecification.
+We argue that this is sensible for misspecification.
+In general, we find that sample-importance-resampling is important for making the inference more robust.
+In particular, it is useful for "cleaning samples" that the posterior has generated which are not compatible with the prior distribution or the likelihood.
+We present an example of this in Figure LABEL:spectra_posterior_comparison_sir.
 Author Contributions
+The programming and the writing of the paper was carried out by MOC.
 Acknowledgments
+MOC is supported by the Gianna Angelopoulos Programme for Science, Technology and Innovation (GAPSTI) through the Science and Technology Facilities Council (STFC) studentship for astronomy.
+GG acknowledges support from The Leverhulme Trust, through Emeritus Fellowship 2025-007.
+KSM is supported by the European Union’s Horizon 2020 research and innovation programme under European Research Council Grant Agreement No 101002652 (BayeSN) and Marie Skłodowska-Curie Grant Agreement No 873089 (ASTROSTAT-II).
 Appendix A Appendix
-A.1 Misspecified Observations Overlaying Training Simulations
+In this section, we discuss some results that are not central to the main claims of the paper.
+A.1 Sample importance resampling for Spectra task well-specified
+Sample-importance-resampling is a useful technique in variational autoencoder frameworks, as it adjusts the variational posterior samples based on the learnt forward model.
+In Figure 7, we show an example of the effect of sample-importance-resampling in the spectra task.
+From this plot, it is clear that the samples from the variational posterior (green) are sampled in regions of parameter space with extremely low prior probability.
+The sample-importance-resampling has the effect of tightening the posterior by removing points which are incompatible with the prior and the simulator model.
+Figure 7: (SIR) indicates sample-importance-resampling.
+We show the effect of sample-importance-resampling on the well-specified spectra task.
+The effect largely reduces the number of points which have very low prior probability and are incompatible with the simulator likelihood proxy.
+A.2 Well-specified case
+In this subsection, we present plots of the metrics when we use RVNP to infer the posterior distribution when the model is well-specified.
+We find that RVNP is not significantly detrimental to posterior recovery in the well-specified case, as it will become under-confident and the sample-importance-resampling distribution becomes very close to the NPE estimate.
+RVNP is quite under-confident in the spectra and pendulum task.
+However, this is highly preferable to overconfidence.
+Moreover, the NPE task is highly over-fit in both of these tasks, and even a small perturbation in output space will give highly wrong posteriors.
+Such a tight posterior would rarely be used in practice.
+Figure 8: Results for the well-specified CS task.
+We conclude that RVNP and its variants can recover robust posterior inference in amortised simulation-based inference when the simulator is well specified.
+However, the error introduced by the error model will inflate the posterior.
+Sample-importance-resampling will sample closest to the NPE estimate.
+The hue in the middle plots indicates the error bar on the NPE and NNPE algorithms.
+For $\alpha$ nearest to 0 is best, with positive values representing underconfidence and negative values representing overconfidence.
+For the log-probability, higher values are better.
+For NRMSE, lower values are better.
+(SIR) indicates sample-importance-resampling.
+Results.
+Figure 9: Results for the well-specified SIR task.
+We conclude that RVNP and its variants can recover robust posterior inference in amortised simulation-based inference when the model is well-specified.
+However, the error introduced by the error model will inflate the posterior.
+Sample-importance-resampling will sample closest to the NPE estimate, and hybrid sample importance-resampling matches the NRMSE of the NPE.
+The hue in the middle plots indicates the error bar on the NPE and NNPE algorithms.
+For $\alpha$ nearest to 0 is best, with positive values representing underconfidence and negative values representing overconfidence.
+For the log-probability, higher values are better.
+For NRMSE, lower values are better.
+(SIR) indicates sample-importance-resampling and should not be confused with the SIR Task.
+Results.
+Figure 10: Results for the well-specified Pendulum task.
+We conclude that RVNP and its variants can recover robust posterior inference in amortised simulation-based inference when the model is well-specified.
+The hue in the middle plots indicates the error bar on the NPE and NNPE algorithms.
+For $\alpha$ nearest to 0 is best, with positive values representing underconfidence and negative values representing overconfidence.
+For the log-probability, higher values are better.
+For NRMSE, lower values are better.
+(SIR) indicates sample-importance-resampling
+Results.
+Figure 11: Results for the well-specified spectra task.
+We conclude that RVNP and its variants can recover robust posterior inference in amortised simulation-based inference.
+The hue in the middle plots indicates the error bar on the NPE and NNPE algorithms.
+For $\alpha$ nearest to 0 is best, with positive values representing underconfidence and negative values representing overconfidence.
+For the log-probability, higher values are better.
+For NRMSE, lower values are better.
+(SIR) indicates sample-importance-resampling
+A.3 Learnt Error Model for $N_{\rm obs}=1000$ in final two summary statistics
+We display the final two dimensions of the output of the simulator in each of the tasks.
+Figure 12: The final two dimensions of the output of the simulator in the CS task, with the bars indicating the direction of the eigenvectors of the covariance matrix in this subspace.
+After clipping the lengths of the vectors above a magnitude of 1 to ignore outliers, the lengths have been scaled to a maximum length of $0.3$ relative to each other.
+Figure 13: The final two dimensions of the output of the simulator in the SIR task, with the bars indicating the direction of the eigenvectors of the covariance matrix in this subspace.
+After clipping the lengths of the vectors above a magnitude of 5 to ignore outliers, the lengths have been scaled to a maximum length of $0.3$ relative to each other.
+Figure 14: The final two dimensions of the output of the simulator in the SIR task, with the bars indicating the direction of the eigenvectors of the covariance matrix in this subspace.
+After clipping the lengths of the vectors above a magnitude of 1 to ignore outliers, the lengths have been scaled to a maximum length of $0.3$ relative to each other.
+Figure 15: The final two dimensions of the output of the simulator in the SIR task, with the bars indicating the direction of the eigenvectors of the covariance matrix in this subspace.
+After clipping the lengths of the vectors above a magnitude of 1 to ignore outliers, the lengths have been scaled to a maximum length of $0.3$ relative to each other.
+A.4 Misspecified Observations Overlaying Training Simulations
 Here, we display the simulated samples and the misspecified points together for each task.
@@
 Figure 19: The simulated samples and the misspecified points for the Spectra Task.
-A.2 RVNP Training Pipeline
+A.5 RVNP Training Pipeline
 We provide an overview of the RVNP training pipeline in the following algorithms.
@@
 3:if train_flowtrain\_flow then
-4:  TrainNormalizingFlow()
+4:  TrainNormalisingFlow()
 5:end if
@@
 Algorithm 3 Train Normalising Flow Likelihood
-1:Input: Simulator dataset $D=\{(\mathbf{x}_{\rm sim},\bm{\theta})\}$, normalizing flow $p_{\Psi}$
+1:Input: Simulator dataset $D=\{(\mathbf{x}_{\rm sim},\bm{\theta})\}$, normalising flow $p_{\Psi}$
 2:Input: Number of epochs $E_{\rm flow}$, optimizer $\text{Optimizer}_{\Psi}$
@@
 Algorithm 4 Train Variational Posterior and Error Model (RVNP)
-1:Input: Observed dataset $\mathcal{O}$, normalizing flow $p_{\Psi}$, posterior $p_{\phi}$, error model network $\bm{\xi}_{\alpha}$
+1:Input: Observed dataset $\mathcal{O}$, normalising flow $p_{\Psi}$, posterior $q_{\bm{\phi}}$, error model network $\bm{\xi}_{\alpha}$
 2:Input: Number of epochs $E_{\rm var}$, MC samples $K$, optimizers $\text{Optimizer}_{\phi},\text{Optimizer}_{\alpha}$
@@
 4:  for each $\mathbf{x}_{\rm obs}^{(i)}\in\mathcal{O}$ do
-5:   Sample $\bm{\theta}^{(l)}\sim p_{\phi}(\bm{\theta}\mid\mathbf{x}_{\rm obs}^{(i)}),\;l=1..K$
+5:   Sample $\bm{\theta}^{(l)}\sim q_{\bm{\phi}}(\bm{\theta}\mid\mathbf{x}_{\rm obs}^{(i)}),\;l=1..K$
 6:   For each $\bm{\theta}^{(l)}$, sample $\mathbf{x}_{\rm sim}^{(l,m)}\sim p_{\Psi}(\mathbf{x}_{\rm sim}\mid\bm{\theta}^{(l)})$
 7:   Compute IWAE variational loss:
-$\mathcal{L}_{V}^{(i)}=-\log\frac{1}{K}\sum_{l=1}^{K}\frac{\mathbb{E}_{\mathbf{x}_{\rm sim}^{(l,m)}}\big{[}p_{\bm{\xi}_{\alpha}}(\mathbf{x}_{\rm obs}^{(i)}\mid\mathbf{x}_{\rm sim}^{(l,m)})\big{]}\,p(\bm{\theta}^{(l)})\,p(\bm{\xi}_{\alpha}(\bm{\theta}^{(l)}))}{p_{\phi}(\bm{\theta}^{(l)}\mid\mathbf{x}_{\rm obs}^{(i)})}$
+$\mathcal{L}_{V}^{(i)}=-\log\frac{1}{K}\sum_{l=1}^{K}\frac{\mathbb{E}_{\mathbf{x}_{\rm sim}^{(l,m)}}\big[p_{\bm{\xi}_{\alpha}}(\mathbf{x}_{\rm obs}^{(i)}\mid\mathbf{x}_{\rm sim}^{(l,m)})\big]\,p(\bm{\theta}^{(l)})\,p(\bm{\xi}_{\alpha}(\bm{\theta}^{(l)}))}{q_{\bm{\phi}}(\bm{\theta}^{(l)}\mid\mathbf{x}_{\rm obs}^{(i)})}$
 8:  end for
@@
 Algorithm 5 Posterior Tuning (RVNP-T)
-1:Input: Simulator dataset $D$, posterior $p_{\phi}$, error model $\bm{\xi}_{\alpha}$ (fixed)
+1:Input: Simulator dataset $D$, posterior $q_{\bm{\phi}}$, error model $\bm{\xi}_{\alpha}$ (fixed)
 2:Input: Number of epochs $E_{\rm tune}$, optimizer $\text{Optimizer}_{\phi}$
@@
 5:  Compute NPE-style loss with fixed error model:
-$\mathcal{L}_{\rm NPE}=-\mathbb{E}_{(\mathbf{x}_{\rm sim},\bm{\theta})\in B_{\rm sim}}\mathbb{E}_{p_{\bm{\xi}_{\alpha}}(\mathbf{x}_{\rm obs}\mid\mathbf{x}_{\rm sim})}[\log p_{\phi}(\bm{\theta}\mid\mathbf{x}_{\rm sim})]$
+$\mathcal{L}_{\rm NPE}=-\mathbb{E}_{(\mathbf{x}_{\rm sim},\bm{\theta})\in B_{\rm sim}}\mathbb{E}_{p_{\bm{\xi}_{\alpha}}(\mathbf{x}_{\rm obs}\mid\mathbf{x}_{\rm sim})}[\log q_{\bm{\phi}}(\bm{\theta}\mid\mathbf{x}_{\rm sim})]$
 6:  Update $\phi\leftarrow\text{Optimizer}_{\phi}(\nabla_{\phi}\mathcal{L}_{\rm NPE})$
 7:end for
-A.3 Training Procedure
-In every task, RVNP is defined using a variational posterior model based on a rational quadratic spline (RQS) flow with $B=10$ and $15$ knots.
+A.6 Training Procedure
+In every task, RVNP is defined using a variational posterior model based on a rational quadratic spline (RQS) flow with $B=10$ and $15$ knots using the masked autoregressive flow architecture template of papamakarios2018maskedautoregressiveflowdensity and implemented using the flowjax python package.
 The depth of the flow is set to $5$ layers, while the neural network conditioner has a hidden block dimension of $52$.
@@
 The simulator flow has the same architecture as the input and output dimensions swapped.
-The importance weighted autoencoder objective was trained with a batch size of $1024$ over $500$ iterations, using the Adam optimizer with learning rate $10^{-3}$, momentum term $\beta_{1}=0.9$, $\epsilon=10^{-8}$, weight decay $10^{-5}$, gradient clipping at $10.0$, and a warmup schedule of $1000$ steps.
+The importance weighted autoencoder objective was trained with a batch size of $1024$ over $500$ iterations, using the Adam optimiser with a learning rate of $10^{-3}$, momentum term $\beta_{1}=0.9$, $\epsilon=10^{-8}$, weight decay $10^{-5}$, gradient clipping at $10.0$, and a warmup schedule of $1000$ steps.
 Early stopping is applied with a patience of $100$ iterations, and $10\%$ of the data is reserved for validation.
@@
 In each task, the simulator was trained using the same optimiser parameters but using the maximum likelihood loss.
+When sampling with sampling-importance-resampling, we use $5000$ particles per observation and sample $100$ times per particle.
 Neural Statistic Estimator To train the neural statistic estimator on the InfoMax objective, we adopt a neural statistic and a discriminator model.
@@
 The hidden dimension for both the embeddings and the discriminator is 100.
-We describe the algorithm in 5.
+We describe the algorithm in 6.
 Algorithm 6 Spectra Encoder Forward Pass
@@
 5:return logitlogit
-InfoMax Loss Function
-To train the encoders, we maximise the mutual information (MI) between the latent embeddings $z$ and the conditioning variables $\theta$. Algorithm 8 summarises the loss function.
-Algorithm 8 InfoMax (Shannon) Loss Computation
+InfoMax Loss Function chen2021neuralapproximatesufficientstatistics show that finding an embedding $\iota_{\mathbf{\omega}}$ that maximizes the mutual information between $\mathbf{\theta}$ and $\iota_{\mathbf{\omega}}(\mathbf{x})$ tends towards sufficient statistics of $\mathbf{\theta}$ provided the dimension of the embedding space is not too small.
+Furthermore, hjelm2018learning introduce the InfoMax objective for learning deep representations of high-dimensional data based on maximizing the mutual information.
+After adopting a discriminator network $D_{\mathbf{\gamma}}$, we can generate a lower bound on the mutual information using the Shannon-Jensen divergence
+$\mathcal{I}_{\mathrm{JSD}}({\omega,\gamma}):=\mathbb{E}_{P}\big[-\operatorname{sp}\big(-D_{\mathbf{\gamma}}(\mathbf{\theta},\iota_{\mathbf{\omega}}(\mathbf{x}))\big)\big]-\mathbb{E}_{P\times\tilde{P}}\big[\operatorname{sp}\big(D_{\mathbf{\gamma}}(\mathbf{\theta},\iota_{\mathbf{\omega}}(\mathbf{x}))\big)\big],$ (23)
+where $\operatorname{sp}(u)=\log(1+e^{u})$ is the softplus function, $P$ is the joint distribution of $(\mathbf{\theta},\iota_{\mathbf{\omega}}(\mathbf{x}))$, and $\tilde{P}$ denotes the product of marginals.
+Jointly optimizing for the embedding and the discriminator maximises a lower bound on the mutual information.
+To train the encoders, we maximise the mutual information (MI) between the latent embeddings $z$ and the conditioning variables $\theta$.
+Algorithm 9 summarises the loss function.
+Algorithm 9 InfoMax (Shannon-Jensen) Loss Computation
 0: Input batch $x\in\mathbb{R}^{B\times L}$, real batch $x_{\text{real}}$, condition vectors $\theta\in\mathbb{R}^{B\times d_{\theta}}$, encoder $E(\cdot)$, discriminator $D(\cdot)$, number of shuffles $S$

### v3

@@
 A range of SBI methods have emerged to solve the likelihood-free inference problem, beginning with traditional approaches such as approximate Bayesian computation (ABC; 10.1214/aos/1176346785; abc_genetics) and Bayesian synthetic likelihood (BSL, Price02012018).
-Recent work has introduced methods based on neural density estimation, such as neural posterior estimation (NPE, 2016arXiv160506376P; NIPS2017_addfa9b7; greenberg2019automaticposteriortransformationlikelihoodfree), neural likelihood estimation (NLE, pmlr-v96-lueckmann19a; papamakarios2019sequentialneurallikelihoodfast), neural ratio estimation (NRE, pmlr-v33-izbicki14; cranmer2016approximatinglikelihoodratioscalibrated; hermans2020likelihoodfreemcmcamortisedapproximate; pmlr-v119-durkan20a), and modelling the joint distribution of data and simulation parameters (DBLP:conf/icml/GlocklerDWWM24).
+Recent work has introduced methods based on neural density estimation, such as neural posterior estimation (NPE, 2016arXiv160506376P; NIPS2017_addfa9b7; greenberg2019automaticposteriortransformationlikelihoodfree), neural likelihood estimation (NLE, pmlr-v96-lueckmann19a; papamakarios2019sequentialneurallikelihoodfast), neural ratio estimation (NRE, pmlr-v33-izbicki14; cranmer2016approximatinglikelihoodratioscalibrated; hermans2020likelihoodfreemcmcamortizedapproximate; pmlr-v119-durkan20a), and modelling the joint distribution of data and simulation parameters (DBLP:conf/icml/GlocklerDWWM24).
 SBI methods can be categorised into amortised and non-amortised inference methods.
-In the context of neural SBI, non-amortised methods such as sequential neural posterior estimation (SNPE; 2016arXiv160506376P; NIPS2017_addfa9b7; greenberg2019automaticposteriortransformationlikelihoodfree), sequential neural likelihood estimation (SNLE; papamakarios2019sequentialneurallikelihoodfast), and sequential neural ratio estimation (SNRE; hermans2020likelihoodfreemcmcamortisedapproximate) target a single posterior conditioned on fixed data, adapting their inference procedure with each simulation round.
+In the context of neural SBI, non-amortised methods such as sequential neural posterior estimation (SNPE; 2016arXiv160506376P; NIPS2017_addfa9b7; greenberg2019automaticposteriortransformationlikelihoodfree), sequential neural likelihood estimation (SNLE; papamakarios2019sequentialneurallikelihoodfast), and sequential neural ratio estimation (SNRE; hermans2020likelihoodfreemcmcamortizedapproximate) target a single posterior conditioned on fixed data, adapting their inference procedure with each simulation round.
 After an up-front simulation budget, amortised methods aim to learn a global inference model over a given prior, making them well-suited for scenarios where repeated or scalable inference is required.
@@
 Amortised SBI: In this paper, we refer to amortised SBI as amortising the Bayesian inference over all data points in $\mathcal{O}$.
-This definition is in line with amortised variational inference (margossian2023amortised).
-Our solution: We propose robust variational neural posterior estimation (RVNP) and its tuned variant (RVNP-T), methods that address the misspecification problem in amortised SBI.
-RVNP pre-trains a simulator likelihood $p_{\bm{\Psi}}(\bm{x}_{\rm sim}|\,\bm{\theta})$, adopts a flexible error model $p_{\bm{\xi}(\bm{\theta})}(\bm{x}_{\rm obs}|\,\bm{x}_{\rm sim})$, and uses an importance weighted autoencoder (burda2015importance) inference scheme to maximize the evidence of the true data under the variational posterior $q_{\bm{\phi}}(\bm{\theta}|\,\bm{x}_{\rm obs})$ for the parameters $\bm{\xi}$ and $\bm{\phi}$.
+This definition is in line with amortised variational inference (margossian2023amortized).
+Our solution: We propose robust variational neural posterior estimation (RVNP) and its tuned variant (RVNP-T), methods that address the misspecification problem in amortized SBI.
+RVNP pre-trains a simulator likelihood $p_{\bm{\Psi}}(\bm{x}_{\rm sim}|\,\bm{\theta})$, adopts a flexible error model $p_{\bm{\alpha}}(\bm{x}_{\rm obs}|\,\bm{x}_{\rm sim},\bm{\theta})$, and uses an importance weighted autoencoder (burda2015importance) inference scheme to maximize the evidence of the true data under the variational posterior $q_{\bm{\phi}}(\bm{\theta}|\,\bm{x}_{\rm obs})$ for the parameters $\bm{\alpha}$ and $\bm{\phi}$.
 The main claim of our paper is that RVNP and its tuned variant can recover robust amortised posterior inference under misspecification by bridging the simulation-to-reality gap using error modelling.
@@
 We let $p^{*}(\bm{x}_{\rm obs})$ denote the true, unknown, DGP.
-We define the error model as $p_{\bm{\xi}(\bm{\theta})}(\bm{x}_{\rm obs}|\,\bm{x}_{\rm sim})$, a family of distributions parametrized by $\bm{\xi}(\bm{\theta})\in\mathbb{R}^{k}$ and $\bm{x}_{\rm sim}$, which relates the simulations to observations.
-The density explicitly depends on $\bm{x}_{\rm sim}$ and implicitly on $\bm{\theta}$ due to the error model parameters $\bm{\xi}$ being a function of $\bm{\theta}$.
+We define the error model as $p_{\bm{\alpha}}(\bm{x}_{\rm obs}|\,\bm{x}_{\rm sim},\bm{\theta})$, a family of distributions parametrized by $\bm{\alpha}\in\mathbb{R}^{k}$, $\bm{x}_{\rm sim}$, and $\bm{\theta}$, which relates the simulations to observations.
 We assume that we have $D=\{\bm{\theta}^{(i)},\bm{x}_{\rm sim}^{(i)}\}_{i=1}^{N_{\rm sim}}$, a fixed number of points generated from the synthetic DGP $p(\bm{x}_{\rm sim}|\bm{\theta})p(\bm{\theta})$, and a set of observations, $O=\{\bm{x}_{\rm obs}^{(i)}\}_{i=1}^{N_{\rm obs}}$ each associated to a different, unknown true $\bm{\theta}^{*}$ value.
@@
 After choosing a conditional density estimation architecture $q_{\bm{\phi}}(\bm{\theta}|\,\bm{x}_{\rm obs})$ parametrized by $\bm{\phi}$, and an architecture for the neural statistic embedding, $\iota_{\omega}$, NPE (2016arXiv160506376P; NIPS2017_addfa9b7; greenberg2019automaticposteriortransformationlikelihoodfree) fits for the parameters $\bm{\omega}$ and $\bm{\phi}$ by minimizing the the expected forward Kullback-Leibler (KL) divergence between analytic and approximate posterior
-$\mathcal{L}_{\rm NPE}({\bm{\phi},\bm{\omega}})$ $=\mathbb{E}_{p^{*}(x_{\rm obs})}\big[\mathbb{KL}[p(\bm{\theta}|\,\iota_{\bm{\omega}}(\bm{x}_{\rm obs}))||q_{\bm{\phi}}(\bm{\theta}|\,\iota_{\bm{\omega}}(\bm{x}_{\rm obs}))]\big]$ (1)
-$=\mathbb{E}_{p^{*}(x_{\rm obs})}\big[\mathbb{E}_{p(\bm{\theta}|\,\iota_{\bm{\omega}}(\bm{x}_{\rm obs}))}[-\log q_{\bm{\phi}}(\bm{\theta}|\,\iota_{\bm{\omega}}(\bm{x}_{\rm obs}))]\big],$
-where the expectation is over the unknown true data-generating distribution $p^{*}(\bm{x}_{\rm obs})$.
-The second line follows from the first because the true unknown posterior and its entropy do not depend on the trainable parameters.
+$\mathbb{E}_{p^{*}(x_{\rm obs})}\big[\mathbb{KL}[p(\bm{\theta}|\,\bm{x}_{\rm obs})||q_{\bm{\phi}}(\bm{\theta}|\,\iota_{\bm{\omega}}(\bm{x}_{\rm obs}))]\big],$ (1)
+where the expectation is over the unknown true data-generating distribution $p^{*}(\bm{x}_{\rm obs})$. Minimizing this is equivalent to minimizing the NPE loss function
+$\mathcal{L}_{\rm NPE}({\bm{\phi},\bm{\omega}})$ $=\mathbb{E}_{p^{*}(x_{\rm obs})}\big[\mathbb{E}_{p(\bm{\theta}|\,\bm{x}_{\rm obs})}[-\log q_{\bm{\phi}}(\bm{\theta}|\,\bm{x}_{\rm obs})]\big],$ (2)
+since the true unknown posterior and its entropy do not depend on the trainable neural network parameters.
 This amortised posterior objective function is not realistic in most situations (schmitt2024detecting), as we rarely have enough real data to approximate the expectation with respect to $p^{*}(\bm{x}_{\rm obs})$, and the true posterior is unknown.
@@
 Neural likelihood estimation (NLE) learns a distribution maximising the conditional log-probability of the simulated data
-$\mathcal{L}(\bm{\Psi})_{\rm NLE}:=\mathbb{E}_{p(\bm{x}_{\rm sim}|\bm{\theta})p(\bm{\theta})}[\log p_{\bm{\Psi}}(\bm{x}_{\rm sim}|\bm{\theta})]$ (3)
-with respect to the flow parameters $\bm{\Psi}$ (where we dropped reference to $\iota_{\bm{\omega}}$), which is equivalent to minimising the KL divergence between the flow and the target distribution.
-Using neural embeddings for the summary statistics must be implemented in a two-stage manner because it cannot be trained in an end-to-end fashion along with the likelihood proxy (brehmer2020flows).
+$\mathcal{L}(\bm{\Psi})_{\rm NLE}:=\mathbb{E}_{p(\bm{x}_{\rm sim}|\bm{\theta})p(\bm{\theta})}[\log p_{\bm{\Psi}}(\iota_{\bm{\omega}}(\bm{x}_{\rm sim})|\bm{\theta})]$ (4)
+with respect to the flow parameters $\bm{\Psi}$, which is equivalent to minimising the KL divergence between the flow and the target distribution.
+The parameters for the neural summary statistics must be implemented in a two-stage manner because it cannot be trained in an end-to-end fashion along with the likelihood proxy (brehmer2020flows).
 Sampling from the posterior amounts to adopting either a Markov Chain Monte Carlo (MCMC) sampling method or a variational inference algorithm, which will increase the number of architecture choices and hyperparameter tuning, but can recover accurate posterior inference (glockler2022variational).
@@
 The simulator, $p(\bm{x}_{\rm sim}|\,\bm{\theta})$, is said to be misspecified if the true data generative process does not fall within the family of distributions defined by the simulator on the support of the prior of $\bm{\theta}$.
-That is, $q^{*}\notin\{p(x_{\rm sim}|\,\bm{\theta})\,;\bm{\theta}\in\bm{\Theta})\}$ (cannon2022investigating).
-For single observations, the definition of misspecification can be extended to using summary statistics (kelly2024misspecificationrobust) by defining $\bm{b}(\bm{\theta})=\mathbb{E}_{p(\bm{x}_{\rm sim}\mid\bm{\theta})}[\iota_{\omega}(\bm{x}_{\rm sim})]$ and $\bm{b}_{j}=\mathbb{E}_{q^{*}(\bm{x}_{\rm obs})}[\iota_{\omega}(\bm{x}_{\rm obs}^{(j)})]$, for each $\bm{x}_{\rm obs}^{(j)}\in O$ where $O=\{\bm{x}_{\rm obs}^{(i)}\}_{i=1}^{N_{\rm obs}}$ is the set of observations each associated to a different, unknown true $\bm{\theta^{(j)}}$ value.
-The simulator is said to be misspecified if there is no $\{\bm{\theta}\}_{i=1}^{N_{\rm obs}}$ in the support of the prior for which $\bm{b}(\bm{\theta}^{(j)})=\bm{b}_{j}$ for each $j$.
-In this paper, we adopt the alternate, but similar, definition of misspecification provided in wehenkel2025addressing, which is more aligned with amortised SBI.
+That is, $p^{*}\notin\{p(x_{\rm sim}|\,\bm{\theta})\,;\bm{\theta}\in\bm{\Theta})\}$ (cannon2022investigating).
+In this paper, we adopt the adjusted definition of misspecification provided in wehenkel2025addressing, which is more aligned with amortised SBI.
 They define a simulator to be misspecified if $\exists\mathcal{S}\subseteq\Theta\times\mathcal{X}_{\rm obs}:\forall(\bm{\theta},\bm{x}_{\rm obs})\in\mathcal{S}$
@@
 2.4 Importance weighted autoencoders
-Importance-weighted autoencoders (IWAE) were inspired by the vanilla variational autoencoder (VAE, kingma2013auto), which introduces a latent space ${\Theta}\subseteq\mathbb{R}^{m}$ and two distributions, $p_{\bm{\xi}}(\bm{x}|\,\bm{\theta})$ and $q_{\bm{\phi}}(\bm{\theta}|\,\bm{x})$, known as the decoder and the encoder respectively.
+Importance-weighted autoencoders (IWAE) were inspired by the vanilla variational autoencoder (VAE, kingma2013auto), which introduces a latent space ${\Theta}\subseteq\mathbb{R}^{m}$ and two distributions, $p_{\bm{\alpha}}(\bm{x}|\,\bm{\theta})$ and $q_{\bm{\phi}}(\bm{\theta}|\,\bm{x})$, known as the decoder and the encoder respectively.
 The vanilla VAE aims to maximise the evidence of the data using the evidence lower bound
-$\log p(\mathbf{x})\;\geq\;\mathbb{E}_{q_{\bm{\phi}}(\mathbf{\theta}|\mathbf{x})}\left[\log p_{\bm{\xi}}(\mathbf{x}|\mathbf{\theta})\right]-\mathrm{KL}\!\left(q_{\bm{\phi}}(\mathbf{\theta}|\mathbf{x})\,\|\,p(\mathbf{\theta})\right)$ (5)
-with respect to $\bm{\xi}$ and $\bm{\phi}$, where $p(\bm{\theta})$ is a prior over the latent parameter.
+$\log p(\mathbf{x})\;\geq\;\mathbb{E}_{q_{\bm{\phi}}(\mathbf{\theta}|\mathbf{x})}\left[\log p_{\bm{\alpha}}(\mathbf{x}|\mathbf{\theta})\right]-\mathrm{KL}\!\left(q_{\bm{\phi}}(\mathbf{\theta}|\mathbf{x})\,\|\,p(\mathbf{\theta})\right)$ (6)
+with respect to $\bm{\alpha}$ and $\bm{\phi}$, where $p(\bm{\theta})$ is a prior over the latent parameter.
 However, using this bound has been shown to induce mode-seeking behaviour, leading to overly simplified representations and poor inference.
@@
 The log-evidence for a single point is given as
-$\log p_{\bm{\theta}}(\bm{x})$ $\approx\log\left[\frac{1}{k}\sum_{l=1}^{k}\frac{p_{\bm{\xi}}\left(\textbf{x},\bm{\theta}^{(l)}\right)}{q_{\bm{\phi}}\left(\bm{\theta}^{(l)}|\textbf{x}\right)}\right]={\mathcal{L}}^{\text{IWAE}}_{k}\left(\bm{\xi},\bm{\phi};\textbf{x}\right)$ (6)
+$\log p_{\bm{\theta}}(\bm{x})$ $\approx\log\left[\frac{1}{k}\sum_{l=1}^{k}\frac{p_{\bm{\alpha}}\left(\textbf{x},\bm{\theta}^{(l)}\right)}{q_{\bm{\phi}}\left(\bm{\theta}^{(l)}|\textbf{x}\right)}\right]={\mathcal{L}}^{\text{IWAE}}_{k}\left(\bm{\alpha},\bm{\phi};\textbf{x}\right)$ (7)
 $\text{with}\quad\bm{\theta}^{(l)}\sim q_{\bm{\phi}}\left(\bm{\theta}|\textbf{x}\right),$ (8)
@@
   1. 1.
-    $p(\bm{\xi})$ a (pseudo-)prior over the error model parameters.
+    $p(\bm{\alpha})$ a prior over the error model weights.
   2. 2.
@@
   4. 4.
-    $\bm{x}_{\rm obs}\sim p_{\bm{\xi}(\bm{\theta})}(\bm{x}_{\rm obs}|\,\bm{x}_{\rm sim})$, where $p_{\bm{\xi}(\bm{\theta})}(\bm{x}_{\rm obs}|\,\bm{x}_{\rm sim})=\mathcal{N}(\bm{x}_{\rm obs};\,\bm{x}_{\rm sim},\bm{\xi}(\bm{\theta}))$ is an adopted error model conditional on $\bm{\xi}$, where the covariance matrix is the output of a neural network $\rm{NN}(\bm{\theta})$ parametrized by $\bm{\alpha}$.
+    $\bm{x}_{\rm obs}\sim p_{\bm{\alpha}}(\bm{x}_{\rm obs}|\,\bm{x}_{\rm sim},\bm{\theta})$, where $p_{\bm{\alpha}}(\bm{x}_{\rm obs}|\,\bm{x}_{\rm sim},\bm{\theta})=\mathcal{N}(\bm{x}_{\rm obs};\,\bm{x}_{\rm sim},\bm{\xi}(\bm{\theta}\,;\,\bm{\alpha}))$ is an adopted error model where the covariance matrix is the output of a neural network $\rm{NN}(\bm{\theta})$ parametrized by $\bm{\alpha}$.
+We assume that the observed data $\bm{x}_{\rm obs}^{(i)}$ are conditionally independent given their latent parameters $\bm{\theta}^{(i)}$ and the error model parameters $\bm{\alpha}$.
 Under this generative model, the posterior distribution is proportional to
-$p(\{\bm{\theta}^{(i)}\}_{i=1}^{N_{\rm obs}},\bm{\xi}\mid O)\propto\prod_{i=1}^{N_{\rm obs}}\int p_{\bm{\xi}(\bm{\theta}^{(i)})}(\bm{x}^{(i)}_{\rm obs}\mid\bm{x}^{(i)}_{\rm sim})\,p_{\bm{\Psi}}(\bm{x}^{(i)}_{\rm sim}\mid\bm{\theta}^{(i)})\,p(\bm{\theta}^{(i)})\,p(\bm{\xi}(\bm{\theta}^{(i)}))\,d\bm{x}^{(i)}_{\rm sim}$.
+$p(\{\bm{\theta}^{(i)}\}_{i=1}^{N_{\rm obs}},\bm{\alpha}\mid O)\propto p(\bm{\alpha})\prod_{i=1}^{N_{\rm obs}}\int p_{\bm{\alpha}}(\bm{x}_{\rm obs}|\,\bm{x}_{\rm sim},\bm{\theta})\,p_{\bm{\Psi}}(\bm{x}^{(i)}_{\rm sim}\mid\bm{\theta}^{(i)})\,p(\bm{\theta}^{(i)})\,d\bm{x}^{(i)}_{\rm sim}$.
 (11)
@@
 From the posterior distribution of our forward model (Equation 11), we can express the log-evidence of the data as
-$\log p(O)=\sum_{i=1}^{N_{\rm obs}}\log\mathbb{E}_{p_{\bm{\Psi}}(\bm{x}^{(i)}_{\rm sim}|\,\bm{\theta}^{(i)})}[p_{\bm{\xi}(\bm{\theta}^{(i)})}(\bm{x}^{(i)}_{\rm obs}|\,\bm{x}^{(i)}_{\rm sim})p(\bm{\theta}^{(i)})p(\bm{\xi}(\bm{\theta}^{(i)}))].$ (11)
+$\log p(O)=\sum_{i=1}^{N_{\rm obs}}\log\mathbb{E}_{p(\bm{\theta}^{(i)})}\mathbb{E}_{p_{\bm{\Psi}}(\bm{x}^{(i)}_{\rm sim}|\,\bm{\theta}^{(i)})}[p_{\bm{\alpha}}(\bm{x}^{(i)}_{\rm obs}|\,\bm{x}^{(i)}_{\rm sim},\bm{\theta}^{(i)})]+\log p(\bm{\alpha}).$ (12)
 We can use the IWAE lower bound on the log-evidence (Equation 7) to derive the variational loss function for RVNP as
-$\mathcal{L}(\bm{\phi},\bm{\alpha})_{V}\approx-\sum_{i=1}^{N_{\rm obs}}\log\left[\frac{1}{K}\sum_{l=1}^{K}\frac{\mathbb{E}_{p_{\bm{\Psi}}(\mathbf{x}_{\rm sim}\mid\bm{\theta}^{(l)})}\!\left[\,p_{\bm{\xi}(\bm{\theta}^{(i)}\,;\,\bm{\alpha})}\!\left(\mathbf{x}_{\rm obs}^{(i)}\mid\mathbf{x}_{\rm sim}\right)\right]\,p(\bm{\theta}^{(l)})p(\bm{\xi}(\bm{\theta}^{(i)}\,;\,\bm{\alpha}))}{q_{\bm{\phi}}(\bm{\theta}^{(l)}\mid\mathbf{x}_{\rm obs}^{(i)})}\right]$ (12)
-where $\bm{\theta}^{(l)}\sim q_{\bm{\phi}}(\bm{\theta}\mid\mathbf{x}_{\rm obs}^{(i)})$, $\bm{\alpha}$ are the weights of the neural network $\bm{\xi}(\bm{\theta}^{(i)}\,;\,\bm{\alpha})$, and for each $\bm{\theta}^{(l)}$ we approximate $\mathbb{E}_{p_{\bm{\Psi}}(\mathbf{x}_{\rm sim}\mid\bm{\theta}^{(l)})}\!\left[\,p_{\bm{\xi}(\bm{\theta}^{(i)}\,;\,\bm{\alpha})}\!\left(\mathbf{x}_{\rm obs}^{(i)}\mid\mathbf{x}_{\rm sim}\right)\right]$ using a Monte Carlo (MC) estimate.
+$\mathcal{L}(\bm{\phi},\bm{\alpha})_{V}\approx-\sum_{i=1}^{N_{\rm obs}}\log\left[\frac{1}{K}\sum_{l=1}^{K}\frac{\mathbb{E}_{p_{\bm{\Psi}}(\mathbf{x}_{\rm sim}\mid\bm{\theta}^{(l)})}\!\left[\,p_{\bm{\alpha}}(\bm{x}^{(i)}_{\rm obs}|\,\bm{x}^{(i)}_{\rm sim},\bm{\theta}^{(l)})\right]\,p(\bm{\theta}^{(l)})}{q_{\bm{\phi}}(\bm{\theta}^{(l)}\mid\mathbf{x}_{\rm obs}^{(i)})}\right]-\log p(\bm{\alpha})$ (13)
+where $\bm{\theta}^{(l)}\sim q_{\bm{\phi}}(\bm{\theta}\mid\mathbf{x}_{\rm obs}^{(i)})$, and for each $\bm{\theta}^{(l)}$ we approximate $\mathbb{E}_{p_{\bm{\Psi}}(\mathbf{x}_{\rm sim}\mid\bm{\theta}^{(l)})}\!\left[\,p_{\bm{\alpha}}\!\left(\mathbf{x}_{\rm obs}^{(i)}\mid\mathbf{x}_{\rm sim},\bm{\theta}^{(l)}\right)\right]$ using a Monte Carlo (MC) estimate.
 We use the logsumexp function to ensure the MC estimates are stable.
@@
 RVNP-T, the tuned variant of RVNP, includes an extra tuning step that fixes the neural network parameters of the error model $\bm{\alpha}$, and uses the original simulated dataset $D$ to optimise the adjusted NPE objective
-$\mathcal{L}(\bm{\phi})_{\rm{NPE}(\alpha)}:=-\mathbb{E}_{p(\bm{x}_{\rm sim},\bm{\theta})}\mathbb{E}_{p_{\bm{\xi}(\bm{\theta})}(\bm{x}_{\rm obs}|\,\bm{x}_{\rm sim})}[\log q_{\bm{\phi}}(\bm{\theta}|\,\bm{x}_{\rm obs})].$ (13)
+$\mathcal{L}(\bm{\phi})_{\rm{NPE}(\alpha)}:=-\mathbb{E}_{p(\bm{x}_{\rm sim},\bm{\theta})}\mathbb{E}_{p_{\bm{\alpha}}(\bm{x}_{\rm obs}|\,\bm{x}_{\rm sim},\bm{\theta})}[\log q_{\bm{\phi}}(\bm{\theta}|\,\bm{x}_{\rm obs})].$ (14)
 This final objective can be identified with the noisy neural posterior estimation (NNPE, ward2022robust) objective and the error augmentation method suggested by Cranmer_2020.
@@
 The default error model that we adopt in RVNP is given by the Gaussian covariance matrix
-$\bm{\xi}(\bm{\theta})=\rm{Diag}(\rm{NN}(\bm{\theta}\,;\bm{\alpha}))+\bm{\Lambda},$ (14)
+$\bm{\xi}(\bm{\theta}\,;\,\bm{\alpha}))=\rm{Diag}(\rm{NN}(\bm{\theta}\,;\bm{\alpha}))+\bm{\Lambda},$ (15)
 a neural network that outputs the diagonal components of the covariance matrix and the non-diagonal components $\bm{\Lambda}$ are globally inferred.
 The second model we include is a global Gaussian model
-$\bm{\xi}(\bm{\theta})=\bm{\Sigma},$ (15)
+$\bm{\xi}(\bm{\theta}\,;\,\bm{\alpha})=\bm{\alpha},$ (16)
 defined through a full rank Gaussian covariance matrix that is parametrised in terms of a Cholesky decomposition.
@@
 3.5 Prior over the correction model parameters
-In RVNP, a (pseudo-)prior over the error model parameters $p(\bm{\xi}(\bm{\theta}))$ can be introduced to regularise the influence of the error model and make the inference more robust in the single-posterior case.
-In our experiments, we only use a uniform prior distribution over the error model parameters, as we focus on the robustness imparted by multiple observations.
-Algorithm 1 Robust Variational Neural Posterior Estimation (RVNP)
-[…]
-2:repeat
-3:  Sample minibatch $B\subset D$
-4:  $\mathcal{L}_{\rm NLE}(\bm{\Psi})\leftarrow-\mathbb{E}_{B}[\log p_{\bm{\Psi}}(\bm{x}_{\rm sim}\mid\bm{\theta})]$
-5:  $\bm{\Psi}\leftarrow\bm{\Psi}-\eta\nabla_{\bm{\Psi}}\mathcal{L}_{\rm NLE}$
-6:until convergence
-7:return $p_{\bm{\Psi}^{*}}(\bm{x}_{\rm sim}\mid\bm{\theta})$
-8:
-9:LearnPosteriorAndErrorModel($p_{\bm{\Psi}^{*}}$, $\mathcal{O}=\{\bm{x}_{\rm obs}\}$)
-10:repeat
-11:  for each $\bm{x}_{\rm obs}\in$ minibatch $\mathcal{O}$ do
-12:   $\{\bm{\theta}^{(l)}\}_{l=1}^{K}\sim q_{\bm{\phi}}(\bm{\theta}\mid\bm{x}_{\rm obs})$
-13:   for $l=1$ to $K$ do
-14:    $\{\bm{x}_{\rm sim}^{(l,m)}\}_{m=1}^{M}\sim p_{\bm{\Psi}^{*}}(\bm{x}_{\rm sim}\mid\bm{\theta}^{(l)})$
-15:    $w^{(l)}\leftarrow\frac{1}{M}\sum_{m=1}^{M}\frac{p_{\bm{\xi}_{\alpha}(\bm{\theta}^{(l)})}(\bm{x}_{\rm obs}\mid\bm{x}_{\rm sim}^{(l,m)})p(\bm{\theta}^{(l)})p(\bm{\xi}_{\alpha})}{q_{\bm{\phi}}(\bm{\theta}^{(l)}\mid\bm{x}_{\rm obs})}$
-16:   end for
-17:   $\mathcal{L}_{V}\leftarrow-\log\left(\frac{1}{K}\sum_{l=1}^{K}w^{(l)}\right)$
-18:   $(\bm{\phi},\bm{\alpha})\leftarrow(\bm{\phi},\bm{\alpha})-\eta\nabla_{(\bm{\phi},\bm{\alpha})}\mathcal{L}_{V}$
-19:  end for
-20:until convergence
-21:return $q_{\bm{\phi}^{*}}(\bm{\theta}\mid\bm{x}_{\rm obs})$, $p_{\bm{\xi}_{\alpha}^{*}(\bm{\theta})}(\bm{x}_{\rm obs}\mid\bm{x}_{\rm sim})$
-22:
-23:TunePosterior($q_{\bm{\phi}^{*}}$, $p_{\bm{\xi}_{\alpha}^{*}}$, $D$) // Optional: RVNP-T
-24:repeat
-25:  Sample $(\bm{x}_{\rm sim},\bm{\theta})\sim D$, $\bm{x}_{\rm obs}\sim p_{\bm{\xi}_{\alpha}^{*}(\bm{\theta})}(\bm{x}_{\rm obs}\mid\bm{x}_{\rm sim})$
-26:  $\mathcal{L}_{\rm tune}(\bm{\phi})\leftarrow-\log q_{\bm{\phi}}(\bm{\theta}\mid\bm{x}_{\rm obs})$
-27:  $\bm{\phi}\leftarrow\bm{\phi}-\eta\nabla_{\bm{\phi}}\mathcal{L}_{\rm tune}$
-28:until convergence
-29:return $q_{\bm{\phi}^{**}}(\bm{\theta}\mid\bm{x}_{\rm obs})$
-30:
-31:Infer($\bm{x}_{\rm obs}$, $N_{\rm samples}$) // With optional SIR
-32:for $i=1$ to $N_{\rm samples}$ do
-33:  $\bm{\theta}^{(i)}\sim q_{\bm{\phi}^{*}}(\bm{\theta}\mid\bm{x}_{\rm obs})$
-34:end for
-35: Optionally apply sample-importance-resampling to $\{\bm{\theta}^{(i)}\}$
-36:return $\{\bm{\theta}^{(i)}\}_{i=1}^{N_{\rm samples}}$
+In RVNP, a prior over the error model parameters $p(\bm{\alpha})$ can be introduced to regularise the influence of the error model parameters.
+In our experiments, we assume $\log p(\bm{\alpha})=0$ over the error model parameters.
 4 Experiments and results
@@
 Assuming that we have a labelled test set $T=\{\bm{\theta}^{*(i)},\bm{x_{\rm obs}}^{(i)}\}_{i=1}^{N_{\rm test}}$, we evaluate the log posterior probability (LPP) of the true parameter for all points in the dataset.
-LPP has been extensively used in SBI literature to assess performance, and is a useful metric for assessing the ability to predict the truth from observation (2016arXiv160506376P; hermans2020likelihoodfreemcmcamortisedapproximate; ward2022robust;kelly2024misspecificationrobust; wehenkel2025addressing).
+LPP has been extensively used in SBI literature to assess performance, and is a useful metric for assessing the ability to predict the truth from observation (2016arXiv160506376P; hermans2020likelihoodfreemcmcamortizedapproximate; ward2022robust;kelly2024misspecificationrobust; wehenkel2025addressing).
 Given a credible level $\gamma$, let $\rm{HDR}_{q_{\bm{\phi}}(\bm{\theta}|\,\bm{x}_{\rm obs})}(1-\gamma)$ represent the $1-\gamma$ highest posterior density region of the posterior estimator $q_{\bm{\phi}}(\bm{\theta}|\,\bm{x}_{\rm obs})$.
@@
 RVNP is the first example of using variational inference for robust simulation-based inference, and was inspired by robust SBI approaches for synthetic likelihood (frazier2021robust), neural posterior estimation (ward2022robust), and sequential neural likelihood estimation (kelly2024misspecificationrobust).
-We argue that RVNP’s novelty and its ability to recover robust amortised posterior inference without setting misspecification-related tuning parameters constitutes an important step toward reliable posterior inference in amortised SBI.
+We argue that RVNP’s novelty and its ability to recover robust amortized posterior inference without setting misspecification-related tuning parameters constitutes an important step toward reliable posterior inference in amortised SBI.
 Previous work’s reliance on setting a hyperparameter or misspecification prior to control the influence of the misspecification on the posterior inference is undesirable, and as more data are observed we will have a better understanding of the simulation-to-reality gap, which should inform the inference without strong assumptions on the degree of misspecification.
@@
 We found that in the pendulum and the spectra task NNPE performed poorly for these reasons.
-The parameter adjustment method of kelly2024misspecificationrobust can be thought of as an error model $p_{\bm{\xi}}(\bm{x}_{\rm obs}|\,\bm{x}_{\rm sim})=\delta(\bm{x}_{\rm obs}-\bm{x}_{\rm sim}-\bm{\xi})$ after integrating with respect to $\bm{x}_{\rm sim}$, where $\bm{\xi}$ are the linear adjustment parameters.
+The parameter adjustment method of kelly2024misspecificationrobust can be thought of as an error model $p_{\bm{\alpha}}(\bm{x}_{\rm obs}|\,\bm{x}_{\rm sim})=\delta(\bm{x}_{\rm obs}-\bm{x}_{\rm sim}-\bm{\alpha})$ after integrating with respect to $\bm{x}_{\rm sim}$, where $\bm{\alpha}$ are the linear adjustment parameters.
 This approach will move the observed point in summary-statistic space to a region of high probability with respect to the simulated samples.
@@
 In particular, it is useful for "cleaning samples" that the posterior has generated which are not compatible with the prior distribution or the likelihood.
-We present an example of this in Figure LABEL:spectra_posterior_comparison_sir.
+We present an example of this in Figure 7.
 Author Contributions
@@
 We provide an overview of the RVNP training pipeline in the following algorithms.
-Algorithm 2 RVNP Full Pipeline
-1:Input: Flags train_flow,train_variational,tune_posteriortrain\_flow,train\_variational,tune\_posterior
-2:Input: All datasets and models
-3:if train_flowtrain\_flow then
-4:  TrainNormalisingFlow()
-5:end if
-6:if train_variationaltrain\_variational then
-7:  TrainVariationalPosteriorAndErrorModel()
-8:end if
-9:if tune_posteriortune\_posterior then
-10:  PosteriorTuning()
-11:end if
-Algorithm 3 Train Normalising Flow Likelihood
-1:Input: Simulator dataset $D=\{(\mathbf{x}_{\rm sim},\bm{\theta})\}$, normalising flow $p_{\Psi}$
-2:Input: Number of epochs $E_{\rm flow}$, optimizer $\text{Optimizer}_{\Psi}$
-3:for epoch =1=1 to $E_{\rm flow}$ do
-4:  Sample minibatch $B_{\rm sim}\subset D$
-5:  Compute NLE loss:
-$\mathcal{L}_{\Psi}=-\mathbb{E}_{B_{\rm sim}}[\log p_{\Psi}(\mathbf{x}_{\rm sim}\mid\bm{\theta})]$
-6:  Update $\Psi\leftarrow\text{Optimizer}_{\Psi}(\nabla_{\Psi}\mathcal{L}_{\Psi})$
-7:end for
-Algorithm 4 Train Variational Posterior and Error Model (RVNP)
-1:Input: Observed dataset $\mathcal{O}$, normalising flow $p_{\Psi}$, posterior $q_{\bm{\phi}}$, error model network $\bm{\xi}_{\alpha}$
-2:Input: Number of epochs $E_{\rm var}$, MC samples $K$, optimizers $\text{Optimizer}_{\phi},\text{Optimizer}_{\alpha}$
-3:for epoch =1=1 to $E_{\rm var}$ do
-4:  for each $\mathbf{x}_{\rm obs}^{(i)}\in\mathcal{O}$ do
-5:   Sample $\bm{\theta}^{(l)}\sim q_{\bm{\phi}}(\bm{\theta}\mid\mathbf{x}_{\rm obs}^{(i)}),\;l=1..K$
-6:   For each $\bm{\theta}^{(l)}$, sample $\mathbf{x}_{\rm sim}^{(l,m)}\sim p_{\Psi}(\mathbf{x}_{\rm sim}\mid\bm{\theta}^{(l)})$
-7:   Compute IWAE variational loss:
-$\mathcal{L}_{V}^{(i)}=-\log\frac{1}{K}\sum_{l=1}^{K}\frac{\mathbb{E}_{\mathbf{x}_{\rm sim}^{(l,m)}}\big[p_{\bm{\xi}_{\alpha}}(\mathbf{x}_{\rm obs}^{(i)}\mid\mathbf{x}_{\rm sim}^{(l,m)})\big]\,p(\bm{\theta}^{(l)})\,p(\bm{\xi}_{\alpha}(\bm{\theta}^{(l)}))}{q_{\bm{\phi}}(\bm{\theta}^{(l)}\mid\mathbf{x}_{\rm obs}^{(i)})}$
-8:  end for
-9:  Update $\phi,\alpha$ via $\text{Optimizer}_{\phi},\text{Optimizer}_{\alpha}$
-10:end for
-Algorithm 5 Posterior Tuning (RVNP-T)
-1:Input: Simulator dataset $D$, posterior $q_{\bm{\phi}}$, error model $\bm{\xi}_{\alpha}$ (fixed)
-2:Input: Number of epochs $E_{\rm tune}$, optimizer $\text{Optimizer}_{\phi}$
-3:for epoch =1=1 to $E_{\rm tune}$ do
-4:  Sample minibatch $B_{\rm sim}\subset D$
-5:  Compute NPE-style loss with fixed error model:
-$\mathcal{L}_{\rm NPE}=-\mathbb{E}_{(\mathbf{x}_{\rm sim},\bm{\theta})\in B_{\rm sim}}\mathbb{E}_{p_{\bm{\xi}_{\alpha}}(\mathbf{x}_{\rm obs}\mid\mathbf{x}_{\rm sim})}[\log q_{\bm{\phi}}(\bm{\theta}\mid\mathbf{x}_{\rm sim})]$
-6:  Update $\phi\leftarrow\text{Optimizer}_{\phi}(\nabla_{\phi}\mathcal{L}_{\rm NPE})$
-7:end for
+Algorithm 1 Robust Variational Neural Posterior Estimation (RVNP)
+[…]
+2:repeat
+3:  Sample minibatch $B\subset D$
+4:  $\mathcal{L}_{\rm NLE}(\bm{\Psi})\leftarrow-\mathbb{E}_{B}[\log p_{\bm{\Psi}}(\bm{x}_{\rm sim}\mid\bm{\theta})]$
+5:  $\bm{\Psi}\leftarrow\bm{\Psi}-\eta\nabla_{\bm{\Psi}}\mathcal{L}_{\rm NLE}$
+6:until convergence
+7:return $p_{\bm{\Psi}^{*}}(\bm{x}_{\rm sim}\mid\bm{\theta})$
+8:
+9:LearnPosteriorAndErrorModel($p_{\bm{\Psi}^{*}}$, $\mathcal{O}=\{\bm{x}_{\rm obs}\}$)
+10:repeat
+11:  for each $\bm{x}_{\rm obs}\in$ minibatch $\mathcal{O}$ do
+12:   $\{\bm{\theta}^{(l)}\}_{l=1}^{K}\sim q_{\bm{\phi}}(\bm{\theta}\mid\bm{x}_{\rm obs})$
+13:   for $l=1$ to $K$ do
+14:    $\{\bm{x}_{\rm sim}^{(l,m)}\}_{m=1}^{M}\sim p_{\bm{\Psi}^{*}}(\bm{x}_{\rm sim}\mid\bm{\theta}^{(l)})$
+15:    $w^{(l)}\leftarrow\frac{1}{M}\sum_{m=1}^{M}\frac{p_{\bm{\alpha}}(\bm{x}_{\rm obs}\mid\bm{x}_{\rm sim}^{(l,m)},\bm{\theta}^{(l)})p(\bm{\theta}^{(l)})}{q_{\bm{\phi}}(\bm{\theta}^{(l)}\mid\bm{x}_{\rm obs})}$
+16:   end for
+17:   $\mathcal{L}_{V}\leftarrow-\log\left(\frac{1}{K}\sum_{l=1}^{K}w^{(l)}\right)-\log p(\bm{\alpha)}$
+18:   $(\bm{\phi},\bm{\alpha})\leftarrow(\bm{\phi},\bm{\alpha})-\eta\nabla_{(\bm{\phi},\bm{\alpha})}\mathcal{L}_{V}$
+19:  end for
+20:until convergence
+21:return $q_{\bm{\phi}^{*}}(\bm{\theta}\mid\bm{x}_{\rm obs})$, $p_{\bm{\alpha}^{*}}(\bm{x}_{\rm obs}\mid\bm{x}_{\rm sim},\bm{\theta})$
+22:
+23:TunePosterior($q_{\bm{\phi}^{*}}$, $p_{\alpha^{*}}$, $D$) // Optional: RVNP-T
+24:repeat
+25:  Sample $(\bm{x}_{\rm sim},\bm{\theta})\sim D$, $\bm{x}_{\rm obs}\sim p_{\bm{\alpha}^{*}}(\bm{x}_{\rm obs}\mid\bm{x}_{\rm sim},\bm{\theta})$
+26:  $\mathcal{L}_{\rm tune}(\bm{\phi})\leftarrow-\log q_{\bm{\phi}}(\bm{\theta}\mid\bm{x}_{\rm obs})$
+27:  $\bm{\phi}\leftarrow\bm{\phi}-\eta\nabla_{\bm{\phi}}\mathcal{L}_{\rm tune}$
+28:until convergence
+29:return $q_{\bm{\phi}^{**}}(\bm{\theta}\mid\bm{x}_{\rm obs})$
+30:
+31:Infer($\bm{x}_{\rm obs}$, $N_{\rm samples}$) // With optional SIR
+32:for $i=1$ to $N_{\rm samples}$ do
+33:  $\bm{\theta}^{(i)}\sim q_{\bm{\phi}^{*}}(\bm{\theta}\mid\bm{x}_{\rm obs})$
+34:end for
+35: Optionally apply sample-importance-resampling to $\{\bm{\theta}^{(i)}\}$
+36:return $\{\bm{\theta}^{(i)}\}_{i=1}^{N_{\rm samples}}$
 A.6 Training Procedure
@@
 The hidden dimension for both the embeddings and the discriminator is 100.
-We describe the algorithm in 6.
+We describe the algorithm in 2.
 Algorithm 2 Spectra Encoder Forward Pass